### v1

# Quasinormal spectra of a wormhole family: overtone features and a parameter-controlled redshift

Abstract
Though investigated extensively in the past, we take a detailed relook at the study of quasinormal modes (QNM) in a known family of wormholes (ultrastatic ($g_{00}=-1$), as well as spacetimes with different, non-constant $g_{00}$), which includes the familiar Bronnikov-Ellis spacetime as a special case. Our focus here is to go beyond the fundamental mode and obtain some of the QNM overtones using a suitable numerical scheme. Scalar and axial gravitational QNMs including two or three overtones are obtained for the family of ultrastatic geometries and their parameter dependencies are shown explicitly. Further, we comment on how (a) the overtones may influence the time-domain profile in a perturbation and (b) in what sense, the use of overtones may help in distinguishing between different geometries within the family.
Finally, we show how an effect somewhat similar to the so-called ‘environment induced redshift’ of QNMs, introduced recently (arXiv:2412.18651), may be obtained for spacetimes with non-constant $g_{00}$, via an appropriate tuning of available metric parameters which systematically modify the shapes of the effective potentials arising in the perturbation equations.
I Introduction and summary
The notion of a Lorentzian wormhole spacetime appeared first in the well-known work of Einstein and Rosen Einstein and Rosen (1935), popularly referred to as the Einstein-Rosen bridge spacetime. However, it was Wheeler’s ‘geons’ Wheeler (1955) in the 1950s and 1960s which highlighted wormhole topology of the electric field lines as a model of ‘charge without charge’ and also introduced the term ‘wormhole’ Misner and Wheeler (1957) in the literature. Since then, there have been important landmarks in ‘wormholery’. The early 1970s saw the appearance of the first ‘exact wormhole solution’ in General Relativity coupled to matter, albeit with a somewhat odd (negative energy density) matter field. This rather simple geometry is now referred to as Bronnikov-Ellis geometry Bronnikov (1973); Ellis (1973). In the late 1980s, Euclidean wormholes were very popular among Euclidean quantum gravity proponents at that time. Around the same time, Morris and Thorne Morris and Thorne (1988) and Morris-Thorne-Yurtsever Morris et al. (1988) came up with a couple of seminal papers which re-invented the wormhole idea in a Lorentzian context by bringing in aspects of energy condition violations, traversability and the exotic wormhole time-machine model. The Morris-Thorne paradigm was hotly pursued by many–most notably by Visser Visser (1995) who came up with numerous constructions (including the cut-and-paste method, surgery ) of Lorentzian wormholes.
The entire scenario was largely theoretical, with very little effort being given to even suggest possible observational consequences. A large part of the efforts focused on how to avoid energy condition violations Visser et al. (2003); Maeda and Nozawa (2008), and modified/alternate theories of gravity Bhawal and Kar (1992); Kanti et al. (2012); Hochberg (1990); Agnese and La Camera (1995); Shaikh and Kar (2016); Kumar et al. (2024); Radhakrishnan et al. (2024); Dehghani and Dayyani (2009); Sadatian and Hosseini (2024) seemed to be a sort of way out. It was noticed that in a modified theory of gravity, the purely geometric violation of the timelike or null convergence condition does not necessarily imply a violation of the weak, null or dominant energy conditions. Hence, energy conditions, postulated as physical requirements on matter, may hold even though the convergence conditions may be violated, as required for a wormhole to exist. Thus, wormholes seemed more viable in various modified theories of gravity.
Given the above fact that wormholes could be around if any modified theory of gravity is acceptable, even in a restricted domain of the theory parameters, it is obvious that one must find observational signatures. Gravitational lensing of wormholes Cramer et al. (1995); Dey and Sen (2008); Shaikh et al. (2019); Godani and Samanta (2021); Nandi et al. (2006); Eiroa et al. (2001); Izmailov et al. (2019) was one of the first among the observationally relevant works pursued by researchers. It still continues as an important possible signature.
Over the last decade or so, astronomy and astrophysics has reached a different level of observational relevance through the emergence of a large amount of data and results in gravitational wave astronomy (LIGO-VIRGO-KAGRA collaboration Abbott et al. (2023)) and infrared imaging of astrophysical objects (the Event Horizon Telescope (EHT) Akiyama et al. (2019, 2022)). The results obtained tally well with General Relativity predictions, leaving little room for modified gravity. However, there does remain a small but definite possibility of accommodating a modified gravity solution or a sort of exoticity within GR (eg. the Bronnikov-Ellis geometry). It is this possibility which has driven some researchers (including us) to construct models, do calculations with the aim of looking for signatures of wormholes in available data. With more observed events and data in the future, one can always hope for further confirmation of GR or, perhaps, some acceptance of a modification of GR, of course, within a specific domain of parameter values.
One of the central observationally important arenas in GW astrophysics, which has been flooded, in the last decade, by a huge amount of theoretical work by various groups, is the study of quasinormal modes Vishveshwara (1970) in perturbation theory (see Berti et al. (2025) for a very recent review). Scalar and gravitational perturbations have been worked out for diverse black hole Berti ; Berti et al. (2009); Cardoso and Lemos (2001); Nollert (1999); Konoplya (2002, 2003); Rincón and Santos (2020); Blázquez-Salcedo et al. (2018a) and wormhole Bueno et al. (2018); Konoplya and Molina (2005); Kim (2008); Alfaro et al. (2024); Jirón et al. (2025) spacetimes in GR as well as in modified gravity theories. Any such study on a spacetime not among the conventional solutions of GR (eg. Schwarzschild or Kerr) is referred to as a study on a ‘mimicker’. More specifically, wormholes constitute a prominent family among such ‘black hole mimickers’ Abdikamalov et al. (2019); Konoplya et al. (2019); Urbano and Veermäe (2019); Narzilloev et al. (2020); Olivares et al. (2020); Shaikh et al. (2022); Shao et al. (2021); Siemonsen (2024); Pani et al. (2010); Cardoso and Pani (2019); Konoplya and Zhidenko (2016); Glampedakis and Pappas (2018); De Simone et al. (2025). Collectively, such mimickers are also referred to as ‘exotic compact objects’. The main idea here is that these spacetimes could be among possible stable end-states in a post-merger scenario Sanchis-Gual et al. (2019); Cardoso et al. (2016); Krishnendu et al. (2017). Whether this is a feasible option at all, will be evident as more observational results emerge.
Thus, with the above-mentioned outlook and motivation, here we choose to study quasinormal modes (scalar and axial gravitational) of a family of wormholes which are generalisations of the well-known Bronnikov-Ellis spacetimes. It must be stated that much has been done with regard to quasinormal modes for these spacetimes in the recent past. So, it’s worth asking– what is it that we intend to focus on, which could be new and interesting?
Our two-fold aim, in this work, may be stated precisely as follows.
$\bullet$ Most of the work on QNMs of the family of spacetimes studied here is focused on the fundamental QNMs, which surely have the largest contribution (amplitude-wise) in a time domain profile. Despite this, with possible detection Capano et al. (2023) (though disputed Cotesta et al. (2022)) of overtones and the possible detectability of overtones in LISA Oshita and Tsuna (2023), it is always important to track down the contributions of the so-called ‘overtones’. This, indeed, has been a line of thought of late, with varied motivations. Researchers were keen on finding overtones from the onset of QNM analysis. Chandrasekhar and Detweiler Chandrasekhar and Detweiler (1975) computed the fundamental mode along with a couple of overtones for the Schwarzschild metric. Later, the QNM spectra with overtones for RN Ferrari and Mashhoon (1984), Kerr Detweiler (1980); Leaver (1985), Kerr-Newman Berti and Kokkotas (2005), regular blackhole Konoplya et al. (2022); Konoplya and Zhidenko (2024); Zhang et al. (2024), parametrized black hole Hirano et al. (2024) metrics were computed. The authors of Kim (2008); Batic and Dutykh (2024); Alfaro et al. (2024); Ávalos and Contreras (2022); Jirón et al. (2025); González et al. (2022) computed QNM overtones for various wormhole geometries.
Our aim here is to first obtain these overtones using a numerical scheme which allows us to calculate as many overtones as possible. Once we have the overtones for the classes of geometries we consider here, we use them to figure out how they may affect the approach to stability for a perturbed spacetime.
$\bullet$ It has been found (theoretically) that black hole quasinormal modes, when obtained for a scenario where the black hole is in an environment (e.g. a halo), exhibit a ‘redshift’ w.r.t. the increase of the environment parameter Barausse et al. (2014); Pezzella et al. (2025); Liu et al. (2025).
We show here, for the QNMs and overtones found by us for the wormhole family, that there is indeed a ‘redshift’ when we modify the so-called effective potential by tuning a metric parameter appropriately. In this context, we also provide additional examples of a similar redshift in the simplest of examples–eg. the square barrier or the Poschl-Teller potential.
Our article, which is largely a sequel to Dutta Roy et al. (2020), is organised as follows. In the next section, we provide a summary of the wormhole family we study in this article. Section III quickly outlines the methods adopted in finding the QNMs–more specifically, which method is better suited for finding overtones is emphasised and discussed in more detail here. In Section IV, we summarise the numerical results obtained with explicit plots. Section V discusses the inferences which may be drawn from our studies on the QNMs and overtones. The above-mentioned ‘redshift’ effect is discussed extensively in Section VI. Finally, in Section VII, we conclude with remarks on possible future work. The appendix contains the detailed numerical values of the QNMs, thereby supporting the plots shown in Section IV.
II A brief review on the family of wormhole geometries
We begin with a brief review of Morris-Thorne wormholes Morris and Thorne (1988)(1988) and the theoretical framework for constructing such wormholes within GR. The line element corresponding to a generic, spherically symmetric, static spacetime is given as
$ds^{2}=-e^{2\Psi(r)}dt^{2}+\frac{dr^{2}}{1-\frac{b(r)}{r}}+r^{2}d\theta^{2}+r^{2}sin^{2}\theta d\varphi^{2}$ (1)
where, $\Psi(r)$, in wormhole parlance is named the redshift function and $b(r)$, the shape function. Morris–Thorne defined a Lorentzian traversable wormhole via a set of conditions on $\Psi(r)$ and $b(r)$. These were: (a) $e^{2\Psi(r)}$ never zero (no horizon), (b) $\frac{b(r)}{r}\leq 1$ with $b(r=b_{0})=b_{0}$ (shape and throat) and (c) $\frac{b(r)}{r}\rightarrow 0$ as $r\rightarrow\infty$ (asymptotic flatness).
Much before the Morris-Thorne paper, in 1973, Bronnikov Bronnikov (1973) and Ellis Ellis (1973) independently constructed a static, spherically symmetric and horizon-less wormhole solution to the Einstein equations with a minimally coupled phantom scalar field (scalar field with negative kinetic energy). This solution, commonly referred to as Bronnikov–Ellis wormhole, is the simplest Morris-Thorne wormhole with $\Psi(r)=0$ and $b(r)=\frac{b_{0}^{2}}{r}$ and is represented by the line element
$ds^{2}=-dt^{2}+\frac{dr^{2}}{1-\frac{b_{0}^{2}}{r^{2}}}+r^{2}d\theta^{2}+r^{2}sin^{2}\theta d\varphi^{2}$ (2)
One may also rewrite the line element using the proper radial distance, i.e $r^{2}=b_{0}^{2}+l^{2}$ (where $-\infty\leq l\leq\infty$). In this representation, we have
$ds^{2}=-dt^{2}+dl^{2}+\left(l^{2}+b_{0}^{2}\right)d\Omega_{2}^{2}$ (3)
We may visualise the above line element as representing two asymptotically flat regions ($l\rightarrow\pm\infty$) connected by a throat/bridge of radius $b_{0}$. Note that the radial coordinate $r$ runs from $b_{0}$ to $\infty$ whereas $l$ runs from $\infty$- ∞ to $\infty$ with $l=0$ corresponding to $r=b_{0}$.
We will now present generalisations of the above-mentioned simple geometry. The first one will be an ultrastatic spacetime (i.e. with $g_{00}=-1$) with a different shape function, whereas the other two will have non-constant redshift functions.
II.1 The ultra-static wormhole family
The Bronnikov-Ellis solution may be viewed as a special case of a more general traversable wormhole solution. Such a generalised Bronnikov–Ellis wormhole was proposed in Kar et al. (1995). Here, $b(r)$ is a two-parameter function ($n$ and $b_{0}$). When $n=2$, we recover the original Bronnikov–Ellis geometry. The metric of such a generalised Bronnikov–Ellis wormhole Kar et al. (1995); Dutta Roy et al. (2020) is
$ds^{2}=-dt^{2}+\frac{dr^{2}}{1-\frac{b(r)}{r}}+r^{2}d\theta^{2}+r^{2}sin^{2}\theta d\varphi^{2}$ (4)
$b(r)=r-r^{3-2n}\left(r^{n}-b_{0}^{n}\right){}^{2-\frac{2}{n}}$
Alternatively, the line element can be expressed as
$ds^{2}=-dt^{2}+dl^{2}+(l^{n}+b_{0}^{n})^{\frac{2}{n}}(d\theta^{2}+sin^{2}\theta d\varphi^{2})$ (5)
Here, $l$ ([,])absent($\in$[$\infty$,$\infty$])( ∈ [ - ∞ , ∞ ] ) is the tortoise coordinate
$dl^{2}=\frac{dr^{2}}{1-\frac{b(r)}{r}}\implies r(l)=(l^{n}+b_{0}^{n})^{\frac{1}{n}}$
We will consider only the even values of the parameter $n$ so that $r(l)$ remains smooth and non-zero everywhere in the domain of $l$.
II.2 Bronnikov–Ellis wormhole with asymptotic mass
Till now, we have discussed the ultrastatic, symmetric and massless Bronnikov–Ellis geometry and its generalisation. However, the authors of Bronnikov (1973), Ellis (1973) (see also Blázquez-Salcedo et al. (2018b)) have proposed a metric for a massive, static, Bronnikov–Ellis wormhole for which the line element is of the following form:
$ds^{2}=-e^{f(l)}dt^{2}+e^{-f(l)}[dl^{2}+(l^{2}+b_{0}^{2})d\Omega^{2}]$ (6)
with
$f(l)=\frac{C}{b_{0}}\left(arctan\left(\frac{l}{b_{0}}\right)-\frac{\pi}{2}\right)$
In the asymptotic limit $C=2M$, where $M$ is the mass extracted in the asymptotically flat region. We may extend this geometry further by incorporating the $r(l)$ discussed in the previous subsection. This leads to the new line element given as,
$ds^{2}=-e^{f(l)}dt^{2}+e^{-f(l)}[dl^{2}+(l^{n}+b_{0}^{n})^{\frac{2}{n}}d\Omega^{2}]$ (7)
For $M=0$, we recover the symmetric generalised Bronnikov–Ellis wormhole, while any non-zero value of $M$ will introduce asymmetry in the wormhole shape. The QNMs for the $n=2$, $M\neq 0$ geometry will be discussed in Section VI.3.
The energy conditions in general relativity serve as criteria to ensure the physical viability of solutions to Einstein’s field equations. The creation and maintenance of traversable wormholes were first rigorously analysed by Morris and Thorne Morris and Thorne (1988), who explored the energy conditions that such structures must satisfy. Their study revealed that traversable wormholes in general relativity (GR) inevitably violate the Weak Energy Condition (WEC), at least at the throat of the wormhole. Such a violation implies the presence of exotic matter —i.e. matter with negative energy density. Subsequent studies have confirmed that all static wormhole solutions in GR inherently require exotic matter for their support. This reliance on exotic matter is considered a major drawback, as it challenges the existence of wormholes within the framework of classical GR.
To analyse this, consider the energy-momentum tensor $T_{\mu\nu}$, which describes the distribution of energy and matter in spacetime. In the orthonormal frame basis, its diagonal components are defined as: $T^{\mu}_{\nu}=diag(-\rho(l),\tau(l),p(l),p(l)).$ Assuming the Einstein field equations $G_{\mu\nu}=\kappa T_{\mu\nu}$ and $\kappa=1$ units, the energy-momentum tensor (written down from the metric and its Einstein tensor) of the generalised Bronnikov–Ellis wormhole with asymptotic mass is:
$\rho(l)=\rho_{\phi}+\rho_{e}=$ $\left[-\frac{c_{0}^{2}e^{f(l)}}{2\left(b_{0}^{\text{n}}+l^{\text{n}}\right){}^{\frac{4}{n}}}\right]+\Bigg{[}e^{f(l)}\Bigg{(}\frac{c_{0}^{2}}{2\left(b_{0}^{n}+l^{n}\right){}^{\frac{4}{n}}}+\frac{2l^{n-2}\left(lf^{\prime}(l)-n+1\right)}{b_{0}^{n}+l^{n}}$
$+\frac{1}{\left(b_{0}^{n}+l^{n}\right){}^{\frac{2}{n}}}+\frac{(2n-3)l^{2n-2}}{\left(b_{0}^{n}+l^{n}\right){}^{2}}+f^{\prime\prime}(l)-\frac{f^{\prime}(l)^{2}}{4}\Bigg{)}\Bigg{]}$ (8)
$\tau(l)=\tau_{\phi}+\tau_{e}=\left[-\frac{c_{0}^{2}e^{f(l)}}{2\left(b_{0}^{\text{n}}+l^{\text{n}}\right){}^{\frac{4}{n}}}\right]+\left[e^{f(l)}\left(\frac{c_{0}^{2}}{2\left(b_{0}^{n}+l^{n}\right){}^{\frac{4}{n}}}-\frac{1}{\left(b_{0}^{n}+l^{n}\right){}^{\frac{2}{n}}}+\frac{l^{2n-2}}{\left(b_{0}^{n}+l^{n}\right){}^{2}}-\frac{f^{\prime}(l)^{2}}{4}\right)\right]$ (9)
$p(l)=p_{\phi}+p_{e}=\left[\frac{c_{0}^{2}e^{f(l)}}{2\left(b_{0}^{\text{n}}+l^{\text{n}}\right){}^{\frac{4}{n}}}\right]+\left[e^{f(l)}\left(-\frac{c_{0}^{2}}{2\left(b_{0}^{n}+l^{n}\right){}^{\frac{4}{n}}}+\frac{(n-1)b_{0}^{n}l^{n-2}}{\left(b_{0}^{n}+l^{n}\right){}^{2}}+\frac{f^{\prime}(l)^{2}}{4}\right)\right]$ (10)
It is clear that the energy densities and pressures change with change in $n$ or $M$ ($C$). In each of the above expressions, the first term of the right-hand side is due to the phantom field, and the remaining terms denote the effects of extra matter. For the phantom field, the components of the energy-momentum tensor have the form
$\rho_{\phi}=\tau_{\phi}=-p_{\phi}=-\frac{e^{f(l)}}{2}\phi^{\prime 2}\hskip 28.45274pt\phi^{\prime}=\frac{d\phi}{dl}$ (11)
Here the scalar field $\phi(l)$ is found by solving the Klein-Gordon equation $\Box\phi=0$. Interestingly, the $f(l)$ in the line element does not appear in the KG equation. Solving for $\phi(l)$, we get,
$\phi(l)=c_{0}\frac{l\,\,_{2}F_{1}[1/n,2/n,1+1/n,-(l/b_{0})^{n}]}{b_{0}^{2}}$
where ${}_{2}F_{1}[...]$ denotes the hypergeometric function. For $n=2$, $\phi(l)=\frac{c_{0}}{b_{0}}\arctan(\frac{l}{b_{0}})$–same as $f(l)$, after adjusting and adding constants. If $2c_{0}^{2}=C^{2}+4b_{0}^{2}$, then for the $n=2$ case, the contribution due to the extra term vanishes and the matter stress-energy is entirely due to the phantom scalar field.
Hence, the total action for this family of wormhole geometries can be expressed as a sum of the Einstein-Hilbert gravity action ($S_{g}$), the phantom scalar action ($S_{\phi}$) and an extra action ($S_{extra}$), the exact form of which is not known.
$S=S_{g}+S_{\phi}+S_{extra}=\int\sqrt{-g}Rd^{4}x+\int\sqrt{-g}\partial^{\nu}\phi\partial_{\nu}\phi d^{4}x+S_{extra}$ (12)
Now, if we check the WEC conditions ($\rho\geq 0$, $\rho+\tau\geq 0$ and $\rho+p\geq 0$) in the $M=0$ case, we find that for $n=2$, all three conditions are not simultaneously satisfied in any region of $l$. For $n>2$, though the first and third conditions are satisfied in a specific range of $l$, the second one is always violated (except at $l=0$).
Figure 1: For $n=8$ geometry: [left] $\rho$; [middle] $\rho+\tau$; [right] $\rho+p$. Blue solid line denotes $M=0$ and orange dashed line represents $M=0.25b_{0}$ (here $b_{0}=1$). In all the images, the x-axis represents the proper coordinate $l$.
However, in the case of $M\neq 0$ and $n>2$, one can find that all three conditions are satisfied in certain regions if the $|M|$ exceeds a specific value. For different geometries (different values of $n$), the critical value of $M$ and the acceptable regions (corresponding to a fixed value of $M$) are different.
II.3 Bronnikov–Ellis geometry with a chosen redshift function:
It is also possible to choose a suitable redshift function ($-g_{00}$ never zero) in an ad-hoc way while retaining the Bronnikov–Ellis $n=2$ character unaltered in the spacelike slice. The ensuing consequences for required matter and energy conditions can thereafter be looked at. As an example, let us consider the line element studied in Alfaro et al. (2024) and given by,
$e^{2\Psi(r)}=1-\frac{2M}{r}\hskip 56.9055ptb(r)=\frac{b_{0}^{2}}{r}$
where $b_{0}>2M$ is required to make sure that there is no horizon and we have a wormhole. Thus, we have the modified Bronnikov–Ellis geometry with line element given as,
$ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\frac{dr^{2}}{1-\frac{b_{0}^{2}}{r^{2}}}+r^{2}d\theta^{2}+r^{2}sin^{2}\theta d\varphi^{2}$ (13)
For this modified geometry, the components of the energy-momentum tensors ($T^{\mu}_{\nu}=diag(-\rho(r),\tau(r),p(r),p(r))$) have the form Alfaro et al. (2024)
$\rho(r)=-\frac{b_{0}^{2}}{r^{4}},\hskip 42.67912pt\tau(r)=\frac{2Mr-b_{0}^{2}}{r^{3}(r-2M)},\hskip 42.67912ptp(r)=\frac{(M-r)\left(b_{0}^{2}(M-r)+Mr^{2}\right)}{r^{4}(r-2M)^{2}}$ (14)
Note that the energy density is not affected by $M$, though the pressures change. The Weak Energy Condition is violated for all $r>b_{0}$ since the energy density is negative. The Null Energy Condition is also violated with complementary behaviour for the $\rho+\tau$ and $\rho+p$ inequalities w.r.t. increase in $M$ value. As $M$ increases the region of violation (near $r=b_{0}$) of the $\rho+\tau$ inequality shrinks whereas the region of violation (away from $r=b_{0}$) of the $\rho+p$ inequality becomes larger. We have plotted $\rho+\tau$ and $\rho+p$ as a function of the radial coordinate $r$, for two values of $M$ (Fig. 2), in order to show this feature in the NEC violations.
Putting $M=0$, we get back the original Bronnikov–Ellis wormhole. We will discuss the effective potential and the $\omega_{QNM}$ for this spacetime in a later section (VI.4).
Figure 2: For Bronnikov–Ellis wormhole with redshift function: [left] $\rho+\tau$; [right] $\rho+p$. Blue line denotes $M=0.2b_{0}$ and red line represents $M=0.4b_{0}$. (here $b_{0}=1$)
II.4 Effective potentials for scalar and axial perturbation
When an open system is perturbed, it emits radiation and eventually stabilises. QNMs represent the damped oscillations of perturbations in spacetime and can arise due to scalar, electromagnetic or gravitational perturbations.
The original Bronnikov–Ellis wormhole ($n=2$) is a solution to the Einstein equations coupled to a phantom scalar field. The $n>2$ geometries also have contributions from the phantom field. The perturbations in this phantom scalar field ($\Phi$) can give rise to scalar quasinormal modes. For this, we may write the scalar field as $\phi+\Phi$.
$\Box\Phi=\frac{1}{\sqrt{-g}}\partial_{\mu}(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\Phi)=0$ (15)
Putting the ansatz $\Phi=\frac{e^{f(l)/2}}{(l^{n}+b_{0}^{n})^{1/n}}\psi(l)\hskip 2.84544pte^{-i\omega t}Y(\theta,\varphi)$ (here $Y(\theta,\varphi)$ are the spherical harmonics; stationary ansatz: $e^{-i\omega t}$ ) in Eqn. (15), one gets the Schrodinger-like radial equation
$\frac{d^{2}\psi}{dl_{*}^{2}}+(\omega^{2}-V_{sc}(l))\psi=0\hskip 56.9055ptwhere\hskip 28.45274pt\frac{dl_{*}}{dl}=e^{-f}$ (16)
Here, $V_{sc}$ is the effective potential (scalar perturbation) for the massive generalised Bronnikov–Ellis wormhole (Eq. (7)), given by
$V_{sc}(l)=\frac{1}{4}e^{2f(l)}\left(\frac{4m\left(m+1\right)}{\left(l^{n}+b_{0}^{n}\right){}^{2/n}}+\frac{4(\text{n}-1)b_{0}^{\text{n}}l^{\text{n}-2}}{\left(l^{n}+b_{0}^{n}\right){}^{2}}-2f^{\prime\prime}(l)-f^{\prime}(l)^{2}\right)$ (17)
where $m$ is the azimuthal number (from now on, will be referred to as the angular momentum parameter).
In linear perturbation theory, gravitational perturbation is classified into two categories based on parity: axial (odd parity) and polar (even parity). In order to get the master equation for axial perturbation, we follow Chandrasekhar’s method Chandrasekhar (1975). Finally, we get
$\frac{d^{2}\psi}{dl_{*}^{2}}+(\omega^{2}-V_{ax}(l))\psi=0$ (18)
where the effective potential has the form:
$V_{ax}=\frac{1}{4}e^{2f(l)}\left(-\frac{4l^{n-2}\left(2lf^{\prime}(l)+n-1\right)}{l^{n}+b_{0}^{n}}+\frac{4\left(m^{2}+m-2\right)}{\left(l^{n}+b_{0}^{n}\right){}^{2/n}}+\frac{4(n+1)l^{2n-2}}{\left(l^{n}+b_{0}^{n}\right){}^{2}}+2f^{\prime\prime}(l)+3f^{\prime}(l)^{2}\right)$ (19)
Plotting the potentials in the ultrastatic case ($M=0$) reveals that both $V_{sc}$ and $V_{ax}$ have a single-peaked nature (Fig. 4,4). One can notice that if $n>2$, $V_{sc}$ has double barriers for all values of the angular momentum parameter (m). The peaks move towards the origin and become less prominent with increasing m (Fig. 8). The effective potential due to axial perturbation exhibits a triple barrier up to a certain value of m (this critical value of m is different for different n values). If we further increase m, we will first observe potentials with steps and then single barriers with flattened peaks (See Fig. 16).
In the $M\neq 0$ cases, for $n=2$ geometry, both axial and scalar potentials have asymmetric single peaks that get shifted from the origin. The shift and asymmetry are also noticeable for $n>2$ cases, where the height of the peaks on the positive side of $l_{*}$ exceeds that of the negative side (if $M>0$).
III Methods used to find the QNMs
Computation of quasinormal modes requires solving linearised perturbation equations of spacetime under specific boundary conditions (purely outgoing at spatial infinities of the tortoise coordinate). These equations are, in general, non-trivial to solve analytically, and so we employ a variety of numerical, analytical and semi-analytical methods. Below we provide short summaries of the key methods we have used here to determine the QNMs of the generalised Bronnikov–Ellis wormholes.
III.1 Time domain and Prony fit
The time domain profile represents the temporal evolution of the field $\psi$ in a fixed background geometry. It was originally formulated in Gundlach et al. (1994). In the previous section, we derived the master wave equation, which may also be rewritten without the separability ansatz for the field. Thus, the equation, using the new coordinates $u$ and $v$, such that $du=dt-dx,\hskip 5.69046ptdv=dt+dx$, reads Konoplya and Zhidenko (2011):
$\frac{\partial^{2}\psi}{\partial u\partial v}+V(u,v)\psi=0.$ (20)
Rewriting the time evolution operator in the new coordinates and discretising the wave equation, we obtain,
$\psi(N)=\psi(W)+\psi(E)-\psi(S)-\frac{h^{2}}{8}V(S)[\psi(W)+\psi(E)]+O(h^{4}),$ (21)
where $S=(u,v)$, $W=(u+h,v)$, $E=(u,v+h)$, and $N=(u+h,v+h)$ and $h$ is the integration step size for both $u$ and $v$. Eqn. (21) enables us to compute $\psi$ iteratively on the $u-v$ grid based on the initial data specified on null surfaces $u=u_{0}$ and $v=v_{0}$. The values of $\psi(N)$, thus calculated, give us the time-domain profile for the given potential.
The time-domain profile of $\psi$ is then fitted using the Prony method $\psi(t)\approx\sum_{j=1}^{p}C_{j}e^{i\omega_{j}t}$. Here $\omega_{j}$ are the quasinormal frequencies and $C_{j}$ are the corresponding amplitudes. We assume that the ringdown starts at $t=0$ and ends at $t=N_{1}h$, and approximate the profile data as a sum of $p\hskip 2.84544pt(\leq\frac{N_{1}+1}{2})$ terms in the following way:
$x_{n}\approx\psi(nh)=\sum_{j=1}^{p}C_{j}z_{j}^{n},\hskip 56.9055ptz_{j}=e^{-i\omega_{j}nh}$ (22)
In the Prony fit method, it is assumed that the $\psi(nh)$ satisfies the recurrence relation
$\sum_{m=0}^{p}a_{m}\psi((n-m)h)=0\hskip 5.69046pt\implies\hskip 5.69046pt\sum_{m=0}^{p}a_{m}x_{n-m}=0$ (23)
where $a_{0}=1$. This leads to the matrix equation:
$\begin{bmatrix}x_{p-1}&x_{p-2}&\ldots&x_{0}\\ x_{p}&x_{p-1}&\ldots&x_{1}\\ \vdots&\vdots&\ddots&\vdots\\ x_{N-1}&x_{N-2}&\ldots&x_{N-p}\end{bmatrix}\begin{bmatrix}a_{1}\\ a_{2}\\ \vdots\\ a_{p}\end{bmatrix}=-\begin{bmatrix}x_{p}\\ x_{p+1}\\ \vdots\\ x_{N}\end{bmatrix}\hskip 28.45274pt\implies X\cdot a=-x,$ (24)
Eqn. (24) is solved to get $a=-(X^{\dagger}X)^{-1}X^{\dagger}x$, where $X^{\dagger}$ denotes the Hermitian transpose of $X$. This yields all the coefficients $a_{m}$, and thus we may form the master polynomial from Eqn. (23). This polynomial can be solved numerically to obtain the roots $z_{j}$, and subsequently, the QNMs are calculated from these roots
$\omega_{j}=i\frac{\ln(z_{j})}{h}.$ (25)
Having computed the quasi-normal frequencies, we can also obtain their corresponding amplitudes using the Prony fit. For this, we need to define a $(p\times N)$ matrix $A_{ij}=z_{i-1}^{j}$ and a vector $v=[x_{1},x_{2},\ldots,x_{N}]$. We get the amplitudes by calculating $C=[C_{1},C_{2},\ldots,C_{p}]=(A^{\dagger}A)^{-1}A^{\dagger}v$.
In the method discussed above, we get $p$-number of solutions, but most of them are numerical artefacts and not actual QNMs of the system. To mitigate this issue, we can repeat the above steps with different values of $h$ and $N$ and accept only the results common to both sets of solutions (with appropriate tolerance) as the actual QNMs.
The time-domain and Prony fit method works well for most effective potentials which do not have any discontinuities. But this method often fails to return the $\omega_{QNM}$ with low relative amplitude (we notice this feature for odd overtones in Sec. IV).
III.2 The spectral method
Using the corresponding equation for a perturbation (scalar or axial gravitational), another method known as the spectral method, may also be employed to find the QNMs of the system. In our analysis, we will use the Chebyshev grid and thus we have to map our coordinate to a finite interval, i.e. $[-1,1]$. The tortoise coordinate $l\in(-\infty,+\infty)$ may be mapped to $x\in[-1,1]$ using the transformation $x=\frac{2}{\pi}arctan(l)$ Batic and Dutykh (2024).
For the generalised Bronnikov–Ellis wormhole with $M=0$, we write the Schrodinger-like equation in terms of the tortoise coordinate ($l$),
$\frac{d^{2}\psi}{dl^{2}}+(\omega^{2}-V(l))\psi=0$ (26)
Now, we need to ensure that the ‘purely outgoing’ boundary conditions at both infinities of the tortoise coordinate. For that, we examine the behaviour of $\psi$ near the throat and at infinity and define a new variable $\phi(x)$ using a suitable ansatz: Batic and Dutykh (2024).
$\psi(x)=exp(i\omega\frac{2l}{\pi}arctan(l))\phi(x)$
Next, the function $\phi$ is discretized with the Chebyshev polynomial of first kind $C_{i}(x)$ as Jansen (2017)
$\phi(x)\approx\sum_{i=0}^{N}\phi(x_{i})C_{i}(x)$ (27)
with $x_{i}$ being the collocation points (Chebyshev roots) given by $x_{i}=\frac{(2i+1)\pi}{2(N+1)}$ (i = 0,1,2,…,N).
Thus, we can express the master equation in the form
$c_{2}(x,\omega)\phi^{\prime\prime}(x)+c_{1}(x,\omega)\phi^{\prime}(x)+c_{0}(x,\omega)\phi(x)=0$ (28)
and define $c_{p,0}$ as part of $c_{0}$ with coefficient of $\omega^{p}$, $c_{q,2}$ as part of $c_{2}$ with coefficient of $\omega^{q}$ and so on Jansen (2017). We compute these expressions and $\phi$ at the co-location points and get numerical matrices of the form
[…]
where, $D^{(n)}_{ij}$ represents n-th derivative of $C_{i}(x_{j})$. If the highest power of $\omega$ in Eqn. (28) is $p$, then we will need to compute $p$ such matrices. Finally, (28) is expressed as
$(M_{0}+\omega M_{1}+\omega^{2}M_{2}+...+\omega^{p}M_{p})\phi=0$ (29)
Solving the matrix equation (29) will return the quasi-normal frequencies of the system.
III.3 Other methods: merits and demerits
Apart from the methods discussed above, there are a plethora of analytic and numerical methods that can be employed to compute the quasinormal spectra. One can compare numerically obtained QNMs with exact results for the few effective potentials which are analytically solvableFerrari and Mashhoon (1984); Heidari and Hassanabadi (2023). In some cases, the exactly solvable potentials are used as approximations for the real potentials in black hole perturbations (for example, in the Schwarzschild, a matching may be done using the Poschl–Teller potential, as shown many years ago in Ferrari and Mashhoon (1984) ). The QNMs found following this route however do not match accurately when compared with results using other numerical methods. In particular, the $Re(\omega)$ for the overtones show a mismatch.
The direct integration method is another method that we have used to verify the fundamental modes for our geometries. This method involves integrating the perturbation equation outward from one boundary, inward from another boundary and matching at a midpoint to ensure continuity Chandrasekhar and Detweiler (1975). The matching generates a characteristic equation, the roots of which are the QNMs. This method is well suited for fundamental modes of any generic $V_{eff}$ but is not very useful for overtones Pani (2013) (lecture notes).
The WKB method, a well-known semi-analytic, approximation tool, can be extended to the study of QNMs. In both the asymptotically constant regions ($r_{*}\xrightarrow{}\pm\infty$), the solution is approximated as linear combinations of in-going and out-going waves Schutz and Will (1985) and matched with solutions obtained using the behaviour of the effective potential near its maximum. The first two corrections were provided in Iyer and Will (1987), and Konoplya (2003) provided the higher-order corrections. This approach is particularly well-suited for effective potentials which have a single peak and approach a constant value at both horizons.
The continued fraction method is a highly accurate and widely used technique for computing both fundamental modes and overtones Leaver (1985). Here, one transforms the Schrodinger-like equation using a proper ansatz. Next, the solution is expressed in a series form, and a finite-term recurrence relation is obtained, putting the series solution in the transformed equation. This relation is then solved to compute the $\omega_{QNM}$. However, this method is challenging to implement for complicated potentials.
Apart from the methods listed above, there are many more. These include the Asymptotic Iteration Method (AIM) Cho et al. (2012), Physics-Informed-Neural-Networks (PINNs) Cornell et al. (2022); Luna et al. (2024) etc. A recent summary of available methods and codes may be found in Berti et al. (2025).
In the results presented below, we have primarily used the Prony fit (time domain) and the spectral method (SM) while finding the QNMs.
IV Results: quasinormal spectra (fundamental and overtones for the ultra-static case)
The problem of computing quasi-normal modes for the Bronnikov–Ellis wormhole has been tackled in many previous works. The authors of Kim (2008); Blázquez-Salcedo et al. (2018b) and Dutta Roy et al. (2020); Dutta Roy (2022) have presented the fundamental modes for the original Bronnikov–Ellis ($n=2$ case) and the generalised Bronnikov–Ellis wormholes respectively.
However, as pointed out in Sago et al. (2021); Berti et al. (2007); Giesler et al. (2019); Nee et al. (2023), the fundamental modes may not represent the full picture, as the overtones play a crucial role in fitting the ringdown (RD) waveform. The inclusion of overtones (at least, the first few overtones) can help us achieve a better fit of the RD profile at earlier times when we expect a better signal-to-noise ratio (SNR). Overtones can also be instrumental in providing better estimates of the parameters of the underlying geometry. Though the overtones corresponding to the $n=2$ geometry were computed in Alfaro et al. (2024); Batic and Dutykh (2024), we compute the full QNM spectra for the generalised Bronnikov–Ellis spacetimes (even $n>2$ cases). We note many key differences between the $n=2$ and the $n>2$ cases.
The authors of Baibhav et al. (2023) have argued that the overtones having lower real parts compared to the fundamental mode (true for $n=2$), fit away the non-linearities of the RD signal. The same argument will not be true for $n>2$ geometries (where $Re(\omega)$ for overtones are higher in value than the fundamental mode). Any possible future detection of such RD signals may further cement the role of overtones.
As stated above, while computing the $\omega_{QNM}$, we have used a frequency-domain method (spectral method (SM)) in conjunction with a time-domain method (Prony fit). While the frequency-domain method is indispensable for extracting all the modes (up to a certain overtone number), the time-domain provides us with some valuable insight into the relative amplitudes of the overtones, thus pointing to the fact that sometimes even higher-order overtones (in our case: even modes) can have greater importance than the lower-order (in our case: odd modes) ones.
The results from the two methods are in good agreement with each other. However, we could not compute the same number of overtones for every geometry. For lower values of $n$ and angular momentum parameter $m$, we could get very few overtones.
In the series of figures below, we demonstrate the effective potentials and QNMs for scalar and axial gravitational perturbations. Figs. 2-5 are for $n=2$, Figs. $6-9$ for $n=4$, Figs. 10-13 are for $n=6$ and Figs. 14-17 for $n=8$. In the following section, we discuss the results here and try to arrive at specific conclusions on the nature of the QNMs and their link with the different geometries.
QNMs for n = 2 geometry:
Figure 3: Effective potential (Scalar) for Bronnikov–Ellis wormhole with $n=2$.
Figure 4: Effective potential (Axial) for Bronnikov–Ellis wormhole with $n=2$.
Figure 5: QNMs for n = 2 (Scalar) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 8)
Figure 6: QNMs for n = 2 (Axial) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 8)
QNMs for n = 4 geometry:
Figure 7: Effective potential (Scalar) for generalised Bronnikov–Ellis wormhole with $n=4$.
Figure 8: Effective potential (Axial) for generalised Bronnikov–Ellis wormhole with $n=4$.
Figure 9: QNMs for n = 4 (Scalar) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 9)
Figure 10: QNMs for n = 4 (Axial) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 9)
QNMs for n = 6 geometry:
Figure 11: Effective potential (Scalar) for generalised Bronnikov–Ellis wormhole with $n=6$.
Figure 12: Effective potential (Axial) for generalised Bronnikov–Ellis wormhole with $n=6$.
Figure 13: QNMs for n = 6 (Scalar) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 10)
Figure 14: QNMs for n = 6 (Axial) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 10)
QNMs for n = 8 geometry:
Figure 15: Effective potential (Scalar) for generalised Bronnikov–Ellis wormhole with $n=8$.
Figure 16: Effective potential (Axial) for generalised Bronnikov–Ellis wormhole with $n=8$.
Figure 17: QNMs for n = 8 (Scalar) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 11)
Figure 18: QNMs for n = 8 (Axial) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 11)
V Inferences drawn from the QNM spectra: emphasis on overtones
In Section IV, we presented the fundamental modes and overtones of the generalised, ultrastatic, Bronnikov–Ellis wormhole. Now, one may ask whether these QNMs can help distinguish different geometries (i.e. different $n$ values) of this wormhole family. To answer this query, we focus on the plots ($Re(\omega)$ vs $-Im(\omega)$) given above in the previous section and try to identify some general trends in the QNM spectra. In our analysis, we denote the mode numbers with $p$, where the mode with the smallest value of $|Im(\omega)|$ (the longest-lived mode) is the fundamental mode (see e.g. Giesler et al. (2019)) with $p=0$ and modes with higher $|Im(\omega)|$ are the overtones.
For both scalar and axial perturbations, in the $n=2$ case, the real parts of the $\omega_{QNM}$ decrease with increase in mode number ($p$), while for the geometries with higher $n$ values, the higher overtones correspond to an increase in $Re(\omega)$.
From the tables in Appendix A, we notice that for a fixed large value of the angular momentum parameter ($m$), the fundamental modes of higher $n$ geometries are very close in value. However, these wormhole spacetimes can also be distinguished from each other if we take the overtones into account. As we move to larger $n(>2)$ values, we observe that the rate of increase of $Re(\omega)$ with the mode number gets expedited. This is true for both scalar and axial perturbations (see Figs. 6, 6, 10, 10…, 18).
The imaginary parts of the quasinormal spectra are also an important feature distinguishing the geometries. In the scalar perturbation case, for a fixed angular momentum parameter value, the greater the $n$ value, the smaller the magnitude of the imaginary parts of the QNMs. The separation between the $|Im(\omega)|$ also keeps decreasing (See Figs. 6, 10, 14, 18). This implies that for wormholes with larger values of $n$, the fundamental modes and overtones have a lower decay rate and, therefore, are less stable. As we keep on increasing $m$, for the axial perturbation case, the absolute value of the imaginary parts of the fundamental modes first increase and then decrease (See Figs. 6, 10, 14, 18).
Having identified the broader trends from the plots, we can try to explain them from the effective potential. The decrease of $Re(\omega)$ with overtone mode number is only observed in the $n=2$ case, which has a comparatively narrower and sharper single-peaked potential (See Figs. 4, 4). This trend is reversed for all even $n>2$ wormholes, which have a comparatively broader and flatter effective potential with two peaks for scalar perturbation (See Figs. 8, 12, 16) and one (for higher values of m)/three (for smaller values of m) peak(s) for axial perturbations (See Figs. 8, 12, 16). On careful observation, one can notice that as we go to higher values of $n$, the central region of the potential becomes flatter. This can be a possible explanation behind the enhanced rate of increase of the real parts for higher $n$ geometries.
Fixing the value of $n$, if we increase the angular momentum parameter ($m$), the height of the effective potential increases, which is related to increasing the real part of the QNMs. The double peak feature also becomes less prominent for the scalar perturbation case (See Figs. 8, 12, 16). This may explain the decrease of $|Im(\omega)|$ with $m$ in the quasi-normal spectra. For axial perturbations in $n>2$ geometries, the transition of the potential from triple barrier to step-like and then to a single barrier, with increasing m (See Figs. 8, 12, 16), may cause the initial increase and subsequent decrease of $|Im(\omega)|$ for the fundamental modess with m.
Variation Scalar Axial
The rate of increase The rate of increase
Increase n: of $Re(\omega)$ with the mode of $Re(\omega)$ with the mode
(for $n>2$) number gets expedited. number gets expedited.
$Re(\omega)$ increases and $Re(\omega)$ increases and
Increase m: $|Im(\omega)|$ decreases $|Im(\omega)|$ (fundamental mode) first
increases and then decreases.
The rate of increase of The rate of increase of
Increase p: $Re(\omega)$ does not change $Re(\omega)$ increases after a
(for $n>2$) with overtone number. certain overtone number.
$|Im(\omega)|$ increases. $|Im(\omega)|$ increases.
Table 1: Key features of the QNM spectra for massless generalised Bronnikov-Ellis wormhole.
V.1 Approximate Analytic Fit
To further advance our study of the QNM spectra and their dependence on different parameters such as the geometry ($n$), angular momentum parameter ($m$), overtone order ($p$) and the throat ($b_{0}$), we can fit our numerical results (for scalar perturbations) to analytical models.
Here, the throat radius is simply a scaling parameter, and the quasi-normal frequencies are inversely proportional to $b_{0}$. However, the other parameters do not follow such simple rules. To get the fits corresponding to these parameters, we have used the NonLinearModelFit function of Mathematica 14.2. For all the fits obtained in this section, we have used the results from the spectral method.
As mentioned above, the change in the QNMs with changing $n$ values may give us some valuable insights. To this end, we have analysed the dependence of $Re(\omega)$ on overtone mode number ($p$) and the angular momentum parameter ($m$), keeping the geometry parameter $n$ fixed.
Figure 19: QNMs for n = 2 (Scalar Perturbation) case.
Here, we observe that the real parts of the quasi-normal frequencies decrease with overtone mode $p$ but increase with angular momentum $m$.
Figure 20: [Left] QNMs for n = 4; [Right] QNMs for n = 8.
For geometries with higher $n$ $(>2)$, the real parts increase with $m$ as well as with overtone mode number $(p)$. We notice, with increasing $p$ $Re(\omega)$ increases faster for $n=8$ case as compared to say, the $n=4$ case. We have also fitted the imaginary part with the geometry parameter and the real part, for different overtone numbers ($p$), for scalar perturbations.
Figure 21: Fit of Im($\omega$) with n and Re($\omega$): [left] fundamental modes; [middle] 1st overtones; [right] 2nd overtones.
Thus far, we have (i) fitted $Re(\omega_{QNM})$ with angular momentum and overtone mode number and (ii) the $-Im(\omega_{QNM})$ with geometry parameter $n$ and $Re(\omega_{QNM})$. Hence, for given values of $n$, $m$ and $p$, we can numerically estimate the value of $\omega_{QNM}$ (scalar perturbation). The fitting functions discussed so far are approximate. They do not return exact results but provide information about overall trends, which are useful while interpreting our results.
V.2 Detectable Modes
Since we have now obtained the quasinormal modes including some overtones, it is important to ask whether or not these modes are detectable. If they are within observable limits, this must be for certain choices of the parameters involved. To figure this out, we will need to determine the numerical values for the quasinormal frequencies ($\nu$) and decay times ($\tau$), which can be found from
[…]
Advanced LIGO is sensitive to the frequency range of $10\hskip 5.69046ptHz$ to $10\hskip 5.69046ptkHz$. For the generalised Bronnikov-Ellis geometries, we have calculated the QNFs, and all of them can be within the aLIGO frequency range when the throat radius is roughly between $50M_{\odot}$ and $1000M_{\odot}$. For now, we choose $b_{0}=200M_{\odot}$ and present the frequencies and damping times for a few cases, as an illustration. In Table II below we show the values for $n=2$ and $m=5$ (scalar perturbation).
Mode Frequency (Hz) Damping Time (s) Amplitude
0 889.5 0.01235 0.2396
1 872.6 0.00408 2.01×$\times 10^{-5}$
2 845.8 0.00244 0.1219
Table 2: Frequency and Damping time for $n=2$ and $m=5$. These values are calculated from the Prony fit (details in Appendix A).
Figure 22: Log plot of time-domain profile for $\psi$ for $n=2$ and $m=5$.
Figure 23: Time-domain waveform for $n=2$ and $m=5$.
For $n=8$ and $m=5$ case (scalar):
Mode Frequency (Hz) Damping Time (s) Amplitude
0 904.2 0.06467 0.1550
1 967.6 0.01617 1.789×$\times 10^{-5}$
2 1067.7 0.00733 0.1368
3 1200.3 0.00444 0.0006
4 1359.8 0.00296 0.1100
Table 3: Frequency and Damping time for $n=8$ and $m=5$. These values are calculated from the prony fit (details mentioned in Appendix A).
Figure 24: Log plot time-domain profile for $\psi$ for $n=8$ and $m=5$.
Figure 25: Time-domain waveform for $n=8$ and $m=5$.
In the ringdown waveforms, we have represented the actual time-domain (obtained using the procedures mentioned in III.1) wave as the blue points, the fundamental mode as green lines and the resultant wave of the fundamental and overtones ($\omega$ and amplitudes computed using the Prony fit) as the magenta lines (Figs. 23,25). For both cases, the magenta lines better fit the time domain than only the fundamental mode. However, the difference between the green and magenta lines is only noticeable for the first few cycles, as the overtones have higher imaginary parts and decay faster. While the inclusion of overtones makes minimal improvement in the fitting for the $n=2$ case, the effect is more significant for the $n=8$ case. This is because for the $n=8$ case, we could find more overtones, and these overtones have amplitudes comparable to the fundamental mode, whereas, in the $n=2$ case, the amplitudes of the overtones are significantly lower than the fundamental mode. For $n=4,6$ cases, we computed $2-3$ overtones and their corresponding amplitudes using the Prony fit, and similar to the $n=8$ case, we observe that inclusion of the overtones significantly improves the fit at early times. For all the geometries (i.e. different $n$), the more overtones we could get (up to $p=4$), the better we were able to model the time-domain signal, indicating that obtaining more overtones may result in a better fit of the time-domain profile. However, we need to be cautious not to overfit our results by incorporating an excessive number of overtones.
One would expect the amplitudes of the modes to depend on our initial value of $\psi$ but not on which part of the time domain we are fitting. However, for the generalised Bronnikov–Ellis wormhole, only the amplitudes associated with the fundamental modes and 2nd overtones (to some extent) are time-independent. For almost all the other $\omega_{QNM}$, if we change the initial time of the fitting, the corresponding amplitudes remain stable up to some fitting domain, but beyond it, they begin to vary.
Figure 26: Amplitude calculated for $n=2$ and $m=5$ with different initial time.
Figure 27: Amplitude calculated for $n=8$ and $m=5$ with different initial time.
Thus, we again fit the ringdown profiles with the modes with a stable amplitude
Figure 28: Time-domain profile for all computed overtones and the modes with stable amplitude for $n=2$ and $m=5$.
Figure 29: Time-domain profile for all computed overtones and the modes with stable amplitude for $n=8$ and $m=5$.
In both plots (Figs. 29, 29), the black line represents the contribution of the fundamental and 2nd overtone (the modes with stable amplitude), and the magenta line has the contribution of all the modes from tables 2 and 3. For the $n=2$ case, the black and the magenta lines overlap (this is expected as we filtered out the 1st overtone with a very small amplitude). In the $n=8$ case, the two lines deviate at a very early time. The inclusion of all the modes gives a slightly better fit as the modes that were filtered out are also stable in the initial times, and their inclusion does not change the late-time behaviour, as these modes are damped quickly. Thus, we conclude that the inclusion of overtones can indeed help us achieve a better fit of the time-domain profile at earlier times.
VI Redshift
In the previous sections, we have computed the quasinormal modes for the massless generalised Bronnikov–Ellis wormhole. However, if the geometry under study is surrounded by a dark matter halo or any generic matter distribution, the QNMs calculated for the isolated compact object may differ from the observed values, as the environment may affect the generation of GW as well as their propagation. Such environments cause the spectrum to be red-shifted as documented in Barausse et al. (2014); Pezzella et al. (2025); Liu et al. (2025) for the black hole spacetime embedded in a halo.
In the case of the Bronnikov–Ellis wormhole, similar redshifts in the fundamental mode can be found, if one considers the wormhole with non-zero ADM mass, as noted by the authors of Blázquez-Salcedo et al. (2018b). We will first find the QNMs (fundamental and overtones) for the $M\neq 0$ case in the $n=2$ Bronnikov–Ellis geometry. Thereafter, in our second example, we will find the QNMs for a modified $n=2$ geometry with a new redshift function having an extra parameter $M$ Alfaro et al. (2024). In these two scenarios, using the QNMs obtained, we then verify if we have a parameter-controlled redshift, similar to the environment-induced redshift mentioned above. The modelling of such red-shifts is crucial as the red-shifted modes may appear to be QNMs of the original isolated geometry with a larger throat radius (as $\omega$ varies inversely with $b_{0}$).
Let us first fix our general idea using some well-studied potentials as toy examples. We will try to find out how the $\omega_{QNM}$ vary if we change the width or other parameters present in two well-known potentials which have been investigated and used extensively in the past.
VI.1 Rectangular Barrier:
The rectangular barrier has the form:
$V(x)=\begin{cases}V_{0}&\text{for }0\leq x\leq a,\\ 0&\text{otherwise}.\end{cases}$ (30)
The exact transmission amplitude is $a_{t}=4\omega ke^{-i\omega a}((\omega+k)^{2}e^{-ika}-(\omega-k)^{2}e^{ika})^{-1}$, where $k=\sqrt{\omega^{2}-V_{0}}$.
In order to obtain the quasi-normal spectra, we need purely outgoing waves, i.e. $a_{t}\rightarrow\infty$, Boonserm and Visser (2011) which implies $(\omega+k)^{2}e^{-ika}-(\omega-k)^{2}e^{ika}=0$, and we can solve this equation numerically to get the QNMs of the rectangular barrier.
Here, we are interested in how the QNM spectra change if we change the width (a) of the potential while keeping the height of the barrier fixed (say $V_{0}=1$).
Figure 30: Rectangular Barriers with the same height ($V_{0}=1$) and slightly different widths.
Figure 31: QNMs for Rect. Barrier of different widths. Results of the semi-analytical method and Prony fit are represented by a circle and a plus, respectively.
We have computed the QNM spectra (Fig. 31) for a rectangular barrier with three different barrier widths (Fig. 31). We observe that, as we increase the width of the barrier, the $Re(\omega)$ decreases. Thus, we encounter a redshift for wider barriers. The QNM separation is more pronounced for higher overtones (see Figure 31, rightwards).
VI.2 Poschl-Teller Potential:
The potential is defined as: $V=V_{0}sech^{2}(\alpha x)$ and the corresponding master equation is:
$\frac{d^{2}\psi}{dx^{2}}+\left(\omega^{2}-\frac{V_{0}}{\cosh^{2}(\alpha x)}\right)\psi=0.$
Using $\xi=\tanh(\alpha x)$ and rewriting the equation, we get
$\frac{d}{d\xi}\left((1-\xi^{2})\frac{d\psi}{d\xi}\right)+\left[s(s+1)+\frac{\omega^{2}}{\alpha^{2}(1-\xi^{2})}\right]\psi=0\hskip 28.45274ptwhere\hskip 8.5359pts(s+1)=-\frac{V_{0}}{\alpha^{2}}$ (31)
The asymptotic solution at $x\to-\infty$ is given by Landau and Lifshitz (2013):
$\psi\sim e^{-i\omega x}\frac{\Gamma(\frac{i\omega}{\alpha})\Gamma(1-\frac{i\omega}{\alpha})}{\Gamma(-s)\Gamma(1+s)}+e^{i\omega x}\frac{\Gamma(-\frac{i\omega}{\alpha})\Gamma(1-\frac{i\omega}{\alpha})}{\Gamma(-\frac{i\omega}{\alpha}-s)\Gamma(-\frac{i\omega}{\alpha}+s+1)}$
At the asymptotic limits, the solution must be purely outgoing. Thus, we have to find the values of $\omega$ such that the coefficient of $e^{i\omega x}$ becomes zero. We can solve this analytically, and the QNMs are given by Ferrari and Mashhoon (1984); Beyer (1999); Cardona and Molina (2017)
$\omega=\alpha\left[-i\left(n+\frac{1}{2}\right)\pm\sqrt{\frac{V_{0}}{\alpha^{2}}-\frac{1}{4}}\right]$ (32)
where, $n=0$ corresponds to the fundamental mode and $n=1,2,3,...$ represent overtones.
Figure 32: Poschl-Teller potential of different values of $\alpha$.
Figure 33: QNMs for Poschl-Teller of different values of $\alpha$.
From the plots above, one can note that contrary to the rectangular barrier case, for the Poschl-Teller potential, as we increase the value of $\alpha$ (decrease width), keeping the height fixed (Fig. 33), the real part of the $\omega_{QNM}$ decreases (Fig. 33). So, increasing the value of $\alpha$ in the Poschl-Teller potential causes a redshift.
In both the rectangular barrier and the Poschl-Teller potential case, reducing the height of the potential barrier (keeping the width fixed) causes a reduction in the $Re(\omega)$, thus causing a redshift.
VI.3 Bronnikov–Ellis geometry with asymptotic mass:
Let us now see if a similar redshift effect arises for the QNMs of perturbations of Bronnikov–Ellis geometries. The ultrastatic Bronnikov–Ellis geometry is not of use since it has only one parameter $b_{0}$ (throat radius) which essentially appears as an overall scale factor in the QNMs. Thus, we either modify the shape function with extra parameters, in which case it is no longer Bronnikov–Ellis or, alternatively, we may bring in the redshift function which will have an extra parameter. We choose to work with this latter modification.
The well-known, original line element for a Bronnikov–Ellis geometry with a redshift function is given as,
$ds^{2}=-e^{f(l)}dt^{2}+e^{-f(l)}[dl^{2}+(l^{2}+b_{0}^{2})d\Omega_{2}^{2}]$ (33)
where, $f(l)=\frac{C}{b_{0}}\left(arctan\left(\frac{l}{b_{0}}\right)-\frac{\pi}{2}\right)$. Note that due to the presence of $f(l)$, the energy density as well as the pressures change, as discussed in Section II.2.
We will work with this line element in order to demonstrate the redshift effect in the QNMs.
The ADM mass ($M$) of this spacetime is related to the metric parameter $C$ as $C=2M$. One can plot the effective potentials for both scalar and axial perturbations numerically and observe that the symmetry in the potential barriers is lost with the introduction of non-zero asymptotic mass ($M$).
The QNM spectra computed using the Prony fit (from $t=0$ to $t=40$ with $h=1/40$ using Mathematica 14.2 for $M\neq 0$, $b_{0}=1$), for this geometry, are presented in the tables 4,5 . However, in the figures we have presented some results from the spectral method.
The effective potential (scalar perturbation) for this metric can be expressed as Eq. (17). Here, we have plotted the effective potentials for the $n=2$ case ($m=5$ in Fig. 35 and $m=7$ in Fig. 35) and obtained single-peaked potentials with the maxima of the potential shifting to the right as we increase the value of $C$ ($=2M$). The QNMs corresponding to $m=5$ and $m=7$ are plotted in Figs. 37 and 37 respectively.
Figure 34: Effective potentials (Scalar perturbation) for different values of M. (n = 2 and m = 5).
Figure 35: Effective potentials (Scalar perturbation) for different values of M. (n = 2 and m = 7).
Figure 36: QNMs (Scalar) for different values of M. (n = 2 and m = 5). The results of SM and Prony fit are represented by circles and plus signs, respectively.
Figure 37: QNMs (Scalar) for different values of M. (n = 2 and m = 7). The results of SM and Prony fit are represented by circles and plus signs, respectively.
$m$ Mode $M=0b_{0}$ $M=0.025b_{0}$ $M=0.05b_{0}$ $M=0.075b_{0}$ $M=0.1b_{0}$
3 0 $3.53443-0.50686i$ $3.27166-0.46911i$ $3.03576-0.43526i$ $2.82383-0.40465i$ $2.63255-0.37783i$
5 0 $5.52282-0.50282i$ $5.11243-0.46532i$ $4.74378-0.43185i$ $4.41259-0.40161i$ $4.11427-0.37699i$
1 $5.41796-1.52258i$ $5.04381-1.38243i$ $4.65962-1.30864i$ $4.33811-1.21722i$ $4.08699-1.09940i$
2 $5.25123-2.54327i$ $4.85855-2.36301i$ $4.44292-2.19880i$ $4.14809-2.04224i$-
7 0 $7.51934-0.50106i$ $6.95985-0.46386i$ $6.45810-0.43037i$ $6.00711-0.40048i$ $5.60108-0.37349i$
1 $7.45576-1.50817i$ $6.91398-1.38873i$ $6.40556-1.30364i$ $5.95396-1.20411i$ $5.55057-1.11919i$
2 $7.31985-2.52412i$ $6.79114-2.33889i$ $6.28485-2.12915i$ $5.84509-2.02687i$ $5.46827-1.89265i$
9 0 $9.51561-0.50062i$ $8.81008-0.46296i$ $8.17473-0.42973i$ $7.60378-0.39983i$ $7.09023-0.37301i$
1 $9.43351-1.50092i$ $8.76270-1.41073i$ $8.13121-1.29187i$ $7.55941-1.20308i$ $7.04588-1.12225i$
2 $9.35764-2.51354i$ $8.64660-2.31498i$ $8.02958-2.14977i$ $7.50736-2.00358i$ $6.98478-1.87115i$
Table 4: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Scalar perturbation) with $M=0,0.025b_{0},0.05b_{0},0.075b_{0},0.1b_{0}$. (The results are obtained using Prony fit.) [The fundamental modes were computed in Blázquez-Salcedo et al. (2018b). We extended the analysis to include overtones.]
The effective potential for axial perturbation has the form Eq. (19) and are plotted below. With increasing M ($=\frac{C}{2}$), the height of the barrier reduces and the potentials become more asymmetric (See Figs. 39, 39). The figures 41 and 41 display the QNM spectra for axial perturbation in $m=5$ and $m=7$ cases.
Figure 38: Effective potentials (Axial Perturbation) for different values of M. (n = 2 and m = 5).
Figure 39: Effective potentials (Axial Perturbation) for different values of M. (n = 2 and m = 7).
Figure 40: QNMs (Axial Perturbation) for different values of M. (n = 2 and m = 5). The results of SM and Prony fit are represented by circles and plus signs, respectively.
Figure 41: QNMs (Axial Perturbation) for different values of M. (n = 2 and m = 7). The results of SM and Prony fit are represented by circles and plus signs, respectively.
$m$ Mode $M=0b_{0}$ $M=0.025b_{0}$ $M=0.05b_{0}$ $M=0.075b_{0}$ $M=0.1b_{0}$
5 0 $5.15176-0.44366i$ $4.77236-0.43237i$ $4.42882-0.40144i$ $4.12030-0.37367i$ $3.84389-0.34924i$
1 $5.06686-1.39798i$ $4.73155-1.33422i$ $4.36034-1.21916i$ $4.06004-1.12686i$ $3.76455-0.99294i$
2 $4.91958-2.39447i$ $4.55509-2.21569i$- -- --
7 0 $7.25073-0.48257i$ $6.71107-0.44682i$ $6.22750-0.41477i$ $5.79355-0.38607i$ $5.40297-0.36046i$
1 $7.19387-1.45787i$ $6.68408-1.34773i$ $6.17347-1.25075i$ $5.74208-1.16371i$ $5.35677-1.08268i$
2 $7.05993-2.43004i$ $6.52963-2.26610i$ $6.04865-2.09636i$ $5.63629-1.94025i$ $5.24520-1.82708i$
9 0 $9.30408-0.48927i$ $8.61413-0.45250i$ $7.99283-0.41976i$ $7.43555-0.39076i$ $6.97356-0.37212i$
1 $9.25255-1.47571i$ $8.56364-1.36779i$ $7.88056-1.26327i$ $7.40840-1.17056i$ $6.88158-1.09045i$
2 $9.15068-2.45648i$ $8.47519-2.27130i$ $7.84263-2.09637i$ $7.27180-1.96059i$ $6.88756-1.86712i$
Table 5: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Axial perturbation) with $M=0,0.025b_{0},0.05b_{0},0.075b_{0},0.1b_{0}$. (The results are obtained using Prony fit.) [The fundamental modes were computed in Blázquez-Salcedo et al. (2018b). We extended the analysis to include overtones.]
For this geometry, the increase of asymptotic mass (M) can be associated with redshift, i.e. for higher values of M, we have lower values of $Re(\omega)$. This phenomenon is similar to that observed by the authors of Pezzella et al. (2025). We have presented the quasi-normal modes for small values of M, as we could not get overtones for higher values of M. The fundamental modes are in agreement with the results of Blázquez-Salcedo et al. (2018b).
It is interesting to note that both the real and imaginary parts of the $\omega_{QNM}$ decrease with increasing M, with the fundamental mode and the overtones following the same pattern. We can estimate the QNMs for a certain M value from the spectra of the original Bronnikov–Ellis wormhole ($M=0$ case) using the purely empirical formula
$\omega_{M}=\omega_{0}exp\left(-\frac{\pi M}{1+0.8M}\right)$
where, $\omega_{0}$ is quasi-normal mode corresponding to $M=0$ case and $\omega_{M}$ denotes quasi-normal modes corresponding to $M\neq 0$ cases. This relation is true for both scalar and axial perturbations of the $n=2$ geometry and holds for all the values of the overtone number and the angular momentum parameter (i.e. the ones we have obtained numerically). We are not aware of the origin of this empirical relation, though, for small values of the exponent, an approximate linear relation seems to work, as evident from the plots.
In Figs. 43 (scalar) and 43 (axial), we have plotted the QNMs (obtained using the Prony Fit method) and the above empirical fitting function for the $n=2$ and $m=7$ case.
Figure 42: Fit of the scalar $\omega_{QNM}$ ($m=7$) using the empirical fitting function.
Figure 43: Fit of the axial $\omega_{QNM}$ ($m=7$) using the empirical fitting function.
For the $n=2$ case, one can observe that the increasing ADM mass ($M$) has a very similar effect on fundamental mode and overtones. Due to this feature of the QNM spectra, the ringdown signal of a Bronnikov-Ellis wormhole with non-zero $M$ can be mistaken as that of the massless Bronnikov-Ellis wormhole with a larger throat radius ($b_{0}$).
VI.4 Bronnikov–Ellis geometry with a chosen redshift function:
Apart from the geometry we studied in the previous subsection, we may also consider the Bronnikov–Ellis wormhole (the $n=2$ case) with a different non-zero redshift function, introduced and discussed earlier in Section II.3. The line element is given as,
$ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\frac{dr^{2}}{1-\frac{b_{0}^{2}}{r^{2}}}+r^{2}d\theta^{2}+r^{2}sin^{2}\theta d\varphi^{2}$
Here, as stated before, the energy density does not change on varying $M$, but the pressures change. We will now see if a redshift effect of the QNMs is present when we tune the value of $M$.
The Schrodinger-like equation for perturbations of this spacetime is
$\frac{d^{2}\psi(r)}{dr_{*}^{2}}+(\omega^{2}-V_{eff}(r))\psi(r)=0$ (34)
where the tortoise coordinate $r_{*}$ has the form
$r_{*}=$ $\frac{1}{\sqrt{r\left(r^{2}-b_{0}^{2}\right)(r-2M)}(4M^{2}-b_{0}^{2})}\bigg{(}b_{0}\bigg{(}b_{0}^{2}k_{1}(r-2M)^{2}\left(E\left(k_{0}|a_{1}\right)-F\left(k_{0}|a_{1}\right)\right)+2b_{0}M\big{(}k_{1}(r-2M)^{2}\big{(}E\left(k_{0}|a_{1}\right)$
$-\Pi\left(a_{0};k_{0}|a_{1}\right)\big{)}-2Mr\big{)}-4k_{1}M^{2}(r-2M)^{2}\left(F\left(k_{0}|a_{1}\right)-\Pi\left(a_{0};k_{0}|a_{1}\right)\right)+b_{0}^{3}r\bigg{)}+r^{3}(4M^{2}-b_{0}^{2})\bigg{)}$
with $a_{0}=\frac{2b_{0}}{b_{0}+2M}$, $a_{1}=\frac{4M}{b_{0}+2M}$, $k_{0}=sin^{-1}\sqrt{\frac{\left(b_{0}+2M\right)\left(r-b_{0}\right)}{2b_{0}(r-2M)}}$, $k_{1}=\sqrt{\frac{r\left(b_{0}-2M\right)\left(b_{0}^{2}-4M^{2}\right)\left(r^{2}-b_{0}^{2}\right)}{b_{0}^{3}(r-2M)^{3}}}$, $E(...)=EllipticE$, $F(...)=EllipticF$ and $\Pi(...)=EllipticPi$ function.
Now, to compute the QNMs with the spectral method, it is convenient to consider two separate ansatz Alfaro et al. (2024) (one for odd overtones and one for even).
$\psi(r)_{even}=e^{i\omega r}r^{iM\omega}\phi(r)_{even}\hskip 56.9055pt\psi(r)_{odd}=(r-b_{0})^{\frac{1}{2}}e^{i\omega r}r^{iM\omega-\frac{1}{2}}\phi(r)_{odd}$
Further, we have used the transformation $x=1-\frac{2b_{0}}{r}$ to map the radial coordinate $r\in(b_{0},\infty)$ to $x\in[-1,1]$ Batic and Dutykh (2024). This procedure is applicable for both scalar and axial perturbations. The tables below (6, 7) contain results obtained using the spectral method. However, we have also shown the results from the Prony fit method in the plots.
For scalar perturbation the expression for $V_{eff}$ is
$V_{sc}(r)=\frac{b_{0}^{2}(r-3M)+r^{2}\left(\left(-2m^{2}-2m+1\right)M+m(m+1)r\right)}{r^{5}}$
As shown in the figures below (Figs. 45, 45), the effective potential changes from being a single barrier to a double barrier for higher $M$. The potentials have been plotted numerically.
Figure 44: Effective potentials (Scalar Perturbation) for different values of $M$. (n = 2 and m = 5).
Figure 45: Effective potentials (Scalar Perturbation) for different values of $M$. (n = 2 and m = 7).
For this geometry, as we increase the value of $M$, both the real and imaginary parts of the QNM spectra decrease (See Figs. 47 ,47). Thus, an increase in $M$ causes a redshift. On careful observation, one notices that for lower values of $M$, the real parts of quasinormal modes decrease with the overtone number. However, this pattern reverses when we go to higher $M$ values. The height of the effective potential decreases as we keep on increasing $M$, and after a certain value, the effective potential becomes double-peaked. This may explain the above feature of the quasi-normal spectra.
For a fixed overtone number, the $\omega_{QNM}$ of scalar perturbation corresponding to different $M$ values can be roughly estimated by linear fits. The fitting parameter will vary with the overtone number and angular momentum parameter (m).
Figure 46: QNMs (Scalar) for different values of $M$. (n = 2 and m = 5). The results of SM and Prony fit are represented by circles and plus signs, respectively.
Figure 47: QNMs (Scalar) for different values of $M$. (n = 2 and m = 7). The results of SM and Prony fit are represented by circles and plus signs, respectively.
$m$ Mode $M=0b_{0}$ $M=0.1b_{0}$ $M=0.2b_{0}$ $M=0.3b_{0}$ $M=0.4b_{0}$
1 0 $1.57271-0.52970i$ $1.42443-0.44987i$ $1.26184-0.35838i$ $1.07980-0.25189i$ $0.86152-0.12862i$
1 $1.25957-1.70681i$ $1.18531-1.43536i$ $1.12961-1.14815i$ $1.08147-0.84604i$ $1.01104-0.51803i$
2 -- -- $0.97385-2.14007i$ $1.09839-1.63586i$ $1.18106-1.10171i$
3 0 $3.53429-0.50692i$ $3.17040-0.42628i$ $2.76272-0.32941i$ $2.29567-0.20688i$ $1.73832-0.06152i$
1 $3.39029-1.51877i$ $3.06406-1.29199i$ $2.71576-1.00573i$ $2.34754-0.67242i$ $1.93405-0.30467i$
2 $3.10063-2.62534i$ $2.85146-2.20091i$ $2.61615-1.73263i$ $2.38730-1.23570i$ $2.08989-0.70557i$
3 $2.71863-3.55029i$- $2.46305-2.54081i$ $2.39244-1.88891i$ $2.23465-1.18172i$
5 0 $5.52231-0.50296i$ $4.94551-0.42182i$ $4.29517-0.32260i$ $3.54142-0.19058i$ $2.63686-0.03204i$
1 $5.41372-1.50367i$ $4.87885-1.27072i$ $4.27080-0.97686i$ $3.60420-0.61373i$ $2.85899-0.20555i$
2 $5.25011-2.54570i$ $4.74524-2.13600i$ $4.21729-1.65629i$ $3.67073-1.11220i$ $3.02866-0.52636i$
3 $4.97014-3.61705i$ $4.54492-3.03031i$ $4.12908-2.37426i$ $3.71354-1.67104i$ $3.19144-0.91448i$
7 0 $7.51649-0.50162i$ $3.19144-0.91448i$ $5.83575-0.31996i$ $4.79524-0.18192i$ $3.53888-0.01635i$
1 $7.45225-1.50671i$ $6.67910-1.26365i$ $5.82019-0.96550i$ $4.85928-0.58145i$ $3.78057-0.14253i$
2 $7.31542-2.52531i$ $6.58191-2.11572i$ $5.78590-1.62677i$ $4.93587-1.04502i$ $3.95845-0.40847i$
3 $7.11330-3.56071i$ $6.43513-2.98262i$ $5.72857-2.31139i$ $4.99835-1.55865i$ $4.13449-0.74607i$
Table 6: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Scalar perturbation) with $M=0,0.1b_{0}$, $0.2b_{0}$, $0.3b_{0}$ and $0.4b_{0}$. [The QNM spectra for this geometry (single-peaked scalar potentials) were published in Alfaro et al. (2024). We extended this to double-peaked potentials (Higher $M$)]. [Results obtained using the spectral method ($b_{0}=1$).]
In the axial perturbation case, the effective potential in Eq. (34) has the form
$V_{ax}(r)=\frac{b_{0}^{2}(7M-3r)+r^{2}(m(m+1)r-(2m(m+1)+1)M)}{r^{5}}$
Similar to the scalar perturbation, the effective potential due to axial perturbation changes from being a single barrier to a double barrier for higher $M$ (See Figs. 49 ,49).
Figure 48: Effective potentials (Axial Perturbation) for different values of $M$. (n = 2 and m = 5).
Figure 49: Effective potentials (Axial Perturbation) for different values of $M$. (n = 2 and m = 7).
The quasi-normal spectra for axial perturbation have a similar redshift effect as the QNMs for scalar perturbation (Figs. 51 ,51). However, unlike the scalar perturbation case, here, a linear fit does not estimate the $\omega_{QNM}$ of the same overtone order.
For this modified Bronnikov-Ellis wormhole, the spectra change with increasing $M$. For the $M=0$ case, the real part decreases with increasing overtone number, but this trend gradually reverses with increasing $M$. Though both $Re(\omega)$ and $|Im(\omega)|$ decrease with $M$ (similar effect as increasing $b_{0}$ of the original Bronnikov-Ellis spacetime), the detection of the first few overtones can distinguish this geometry (with non-zero redshift function) from the original Bronnikov-Ellis wormhole.
Figure 50: QNMs (Axial) for different values of $M$. (n = 2 and m = 5). The results of SM and Prony fit are represented by circles and plus signs, respectively.
Figure 51: QNMs (Axial) for different values of $M$. (n = 2 and m = 7). The results of SM and Prony fit are represented by circles and plus signs, respectively.
$m$ Mode $M=0b_{0}$ $M=0.1b_{0}$ $M=0.2b_{0}$ $M=0.3b_{0}$ $M=0.4b_{0}$
3 0 $2.95241-0.40998i$ $2.65202-0.34116i$ $2.31704-0.25890i$ $1.93583-0.15765i$ $1.47895-0.04526i$
1 $2.86643-1.25767i$ $2.59943-1.05329i$ $2.32019-0.81701i$ $2.02320-0.54789i$ $1.68143-0.25349i$
2 $2.66056-2.19085i$ $2.47232-1.84041i$ $2.28366-1.46458i$ $2.09262-1.06263i$ $1.83772-0.62125i$
5 0 $5.15482-0.46725i$ $4.61743-0.39009i$ $4.01226-0.29556i$ $3.31317-0.17059i$ $2.47482-0.02713i$
1 $5.07286-1.40857i$ $4.56144-1.17693i$ $4.00038-0.89951i$ $3.38786-0.56119i$ $2.69934-0.18747i$
2 $4.90746-2.37118i$ $4.44773-1.98388i$ $3.96646-1.53674i$ $3.46529-1.03582i$ $2.86794-0.49395i$
3 $4.65661-3.37224i$ $4.27300-2.82560i$ $3.89988-2.21918i$ $3.51728-1.57299i$ $3.03059-0.87000i$
7 0 $7.24818-0.48308i$ $6.48787-0.40372i$ $5.62868-0.30567i$ $4.62788-0.17089i$ $3.42055-0.01444i$
1 $7.18438-1.45261i$ $6.44308-1.21430i$ $5.61809-0.92386i$ $4.69804-0.55209i$ $3.66405-0.13378i$
2 $7.05634-2.43256i$ $6.35314-2.03436i$ $5.59222-1.56056i$ $4.78041-1.00184i$ $3.84139-0.39161i$
3 $6.86588-3.43084i$ $6.21733-2.87112i$ $5.54494-2.22360i$ $4.84772-1.50282i$ $4.01693-0.72270i$
9 0 $9.30171-0.48964i$ $8.32354-0.40939i$ $7.21638-0.30982i$ $5.92045-0.16947i$ $4.34739-0.00712i$
1 $9.25035-1.47100i$ $8.28712-1.23004i$ $7.20768-0.93382i$ $5.98667-0.54158i$ $4.60839-0.09362i$
2 $9.14741-2.45870i$ $8.21414-2.05635i$ $7.18761-1.56991i$ $6.07049-0.97279i$ $4.79271-0.31009i$
3 $8.99239-3.45726i$ $8.10424-2.89222i$ $7.15227-2.22372i$ $6.14625-1.44932i$ $4.97551-0.60743i$
Table 7: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Axial perturbation) with $M=0,0.1b_{0}$, $0.2b_{0}$, $0.3b_{0}$ and $0.4b_{0}$. [The QNM spectra for this geometry (single-peaked scalar potentials) were published in Alfaro et al. (2024). We extended this to aial perturbation and double-peaked potentials (Higher $M$)]. [Results obtained using the spectral method ($b_{0}=1$).]
VII Remarks and future directions
In our work here, we have revisited the problem of computing the QNMs of the generalised Bronnikov-Ellis wormholes with emphasis on the overtones. Further, we have also studied the ‘redshift’ of QNMs induced by varying the metric parameters in analogy with the environment-induced proposed recently. Below, we briefly summarise our main results.
$\bullet$ We extend the ultrastatic, generalised (two parameter–$b_{0}$, $n>2$) Bronnikov-Ellis wormhole formulated in Kar et al. (1995) by constructing a modified metric with a non-zero asymptotic mass $M$ (furthering the analysis in Blázquez-Salcedo et al. (2018b) for $n=2$). The geometry, energy-momentum tensor and the energy conditions for the new three-parameter ($b_{0}$, $n$, $M$) space-times are discussed briefly, followed by the computation of the corresponding effective potentials for scalar and axial gravitational perturbations.
$\bullet$ We have used the spectral method (frequency domain) and the Prony Fit (time domain) to compute the QNM spectra (scalar and axial gravitational) for the ultrastatic case. Beyond the fundamental modes, we have computed 2-3 overtones (4 in some cases), which, as we show, may help in distinguishing different geometries (i.e. with different $n$). We have also pointed out key features of the QNM spectra and how the shape of the effective potentials may explain the trends in QNM behaviour.
$\bullet$ The range of the throat radius ($b_{0}$) for detectable QNMs (in aLIGO) in the ultrastatic case is found. We also establish, for a few cases, using the time-domain waveform, the crucial role that overtones may play in fitting the signal at early times (when SNR is high).
$\bullet$ Finally, we focused on the $n=2$ wormhole family with asymptotic mass $M$ and observed how the increase in $M$ causes a systematic redshift (lowering of values of $Re(\omega)$). We further observed this redshift effect (w.r.t. changing $M$) for the original Bronnikov–Ellis wormhole but with a chosen redshift function ($e^{2\Psi(r)}=1-\frac{2M}{r}$) Alfaro et al. (2024). The distinguishability of these geometries from their massless counterparts (with larger $b_{0}$) has also been mentioned.
Various extensions of this work here are possible. We have not really succeeded in obtaining a large number of overtones which can give us a clearer picture of the full quasinormal mode spectrum. This aspect surely needs to be handled better. Further, though we have identified distinct features and have tried to demonstrate how overtones may affect a signal, a better and more complete understanding is desirable. In addition, learning more about polar perturbations and their overtones remains a difficult task which needs to be addressed. Finally, in the context of the redshift, though we see a correlation with the behaviour of the effective potentials, the results are case-specific and need to be understood in a more generic way.
We hope to return to the abovementioned issues in future investigations.
ACKNOWLEDGEMENTS
ABM thanks the Indian Institute of Technology Kharagpur, India, for supporting him through a fellowship and allowing access to available computational facilities.
Appendix A QNMs for ultrastatic, generalized Bronnikov–Ellis wormhole
We present here, in table form, the values of QNMs (accurate up to 3 decimal places) for the family of wormholes (with $b_{0}=1$). The time-domain profiles are computed with initial conditions: $\psi(0,v)=0.5$ and $\psi(u,0)=exp\left(-\frac{(u-10)^{2}}{10}\right)$ and the $\omega_{QNM}$ are extracted by fitting these time-domains from $t=0$ to $t=20$ with a stepsize of $1/60$. For the starred results, the fitting window is $t=0$ to $t=60$ (note $m=0,1,2$ cases). The amplitudes are also calculated using the Prony fit method. We have not kept the results obtained using the Prony fit method, which significantly vary for different fitting windows.
In this ultra-static spacetime, if $(\omega_{R}-i\omega_{I})$ is a solution, then $(-\omega_{R}-i\omega_{I})$ is also a solution, both having the same amplitude. In the tables of the following sections, we show only the solution with a positive real part and double its amplitude (accounting for the amplitude of both solutions).
Scalar Perturbation Axial Perturbation
$m$ Mode Spectral Method Prony Fit Amplitude Spectral Method Prony Fit Amplitude
0 0 $0.68137-0.61775i$ $0.68402-0.61594i^{*}$- -- -- --
1 $0.49683-2.17988i$- -- -- -- --
1 0 $1.57271-0.52970i$ $1.57274-0.52969i^{*}$- -- -- --
1 $1.25957-1.70681i$- -- -- -- --
2 0 $2.54666-0.51267i$ $2.54925-0.51285i$ $0.35745915$ $1.73769-0.30514i$ $1.73771-0.30513i^{*}$ $0.2689666$
1 $2.35300-1.57522i$- -- $1.71239-1.02010i$- --
2 $1.94641-2.75472i$- -- $1.52495-2.03050i$- --
3 0 $3.53429-0.50692i$ $3.53443-0.50686i$ $0.30122677$ $2.95241-0.40998i$ $2.95253-0.41006i$ $0.29811729$
1 $3.39038-1.53568i$- -- $2.86643-1.25767i$- --
2 $3.09988-2.62324i$ $3.09682-2.62481i$ $0.16837004$ $2.66056-2.19085i$ $2.67518-2.17517i$ $0.171151$
4 0 $4.52705-0.50433i$ $4.52734-0.50424i$ $0.26517902$ $4.07625-0.44911i$ $4.07646-0.44904i$ $0.2640608$
1 $4.41536-1.52340i$- -- $3.99432-1.36285i$- --
2 $4.19023-2.57208i$ $4.19075-2.57242i$ $0.1397532$ $3.80100-2.30237i$ $3.79816-2.30549i$ $0.14061934$
5 0 $5.52231-0.50296i$ $5.52282-0.50282i$ $0.23964086$ $5.15482-0.46725i$ $5.15524-0.46714i$ $0.239214$
1 $5.42996-1.51789i$ $5.41796-1.52258i$ $2.009\times 10^{-5}$ $5.07306-1.40721i$- --
2 $5.24705-2.54683i$ $5.25123-2.54327i$ $0.12190640$ $4.90770-2.37002i$ $4.90815-2.37157i$ $0.12362226$
6 0 $6.51897-0.50214i$ $6.51980-0.50195i$ $0.22033499$ $6.20882-0.47711i$ $6.20954-0.47695i$ $0.2201443$
1 $6.44246-1.50964i$ $6.44313-1.53221i$ $1.886\times 10^{-5}$ $6.13729-1.43537i$- --
2 $6.28646-2.53346i$ $6.28819-2.53269i$ $0.11151401$ $5.99151-2.40927i$ $5.99267-2.40806i$ $0.1115693$
7 0 $7.51649-0.50162i$ $7.51934-0.50106i$ $0.20454891$ $7.24818-0.48307i$ $7.24931-0.48285i$ $0.20497677$
1 $7.44933-1.50895i$ $7.45576-1.50817i$ $2.309\times 10^{-5}$ $7.24818-0.48307i$- --
2 $7.31541-2.52544i$ $7.31985-2.52412i$ $0.10156691$ $7.05635-2.43256i$ $7.05785-2.43166i$ $0.10294975$
Table 8: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=2$) under scalar perturbations for $m=0$ $7$ and axial perturbations for $m=2$ $7$. [The QNM spectra for the $n=2$ case were published in Kim (2008); Batic and Dutykh (2024); Blázquez-Salcedo et al. (2018b); Alfaro et al. (2024)]
Scalar Perturbation Axial Perturbation
$m$ Mode Spectral Method Prony Fit Amplitude Spectral Method Prony Fit Amplitude
0 0 $0.83639-0.39420i$ $0.83639-0.39418i^{*}$- -- -- --
1 $1.78264-1.42784i$- -- -- -- --
1 0 $1.68973-0.28405i$ $1.68974-0.28404i^{*}$- -- -- --
1 $2.28101-1.18068i$- -- -- -- --
2 0 $2.64867-0.24369i$ $2.64856-0.24379i$ $0.30535863$ $1.86212-0.27064i$ $1.86214-0.27064i^{*}$ $0.24995736$
1 $3.09550-0.97920i$- -- $1.97161-0.71811i$- --
2 $3.73267-2.07179i$ $3.72699-2.07264i$ $0.22283351$ $2.55089-1.37091i$- --
3 0 $3.60155-3.36888i$ $3.62865-0.22143i$ $0.24259909$ $3.14398-0.31728i$ $3.14405-0.31724i$ $0.24589997$
1 $4.01130-0.86237i$- -- $3.18751-0.89811i$- --
2 $4.52123-1.81580i$ $4.52041-1.81625i$ $0.18309362$ $3.48149-1.49460i$ $3.48082-1.49494i$ $0.1574885$
4 0 $4.6164-0.20649i$ $4.61663-0.20641i$ $0.20468468$ $4.28394-0.28792i$ $4.28414-0.28787i$ $0.21497073$
1 $4.96092-0.78763i$ $4.95844-0.78954i$ $0.00019476$ $4.40055-0.92088i$- --
2 $5.40832-1.63672i$ $5.41382-1.63803i$ $0.15290777$ $4.59293-1.60512i$ $4.59135-1.60323i$ $0.1200131$
3 $5.94108-2.65940i$- -- $4.98385-2.34634i$- --
4 -- -- -- $5.60020-3.24404i$ $5.57299-3.21739i$ $0.12091509$
5 0 $5.60815-0.19523i$ $5.60859-0.19513i$ $0.1788866$ $5.35662-0.25860i$ $5.35702-0.25851i$ $0.18764768$
1 $5.92623-0.73455i$ $5.92481-0.73471i$ $0.00011546$ $5.53163-0.86711i$ $5.54394-0.87820i$ $1.795\times 10^{-5}$
2 $6.33667-1.50966i$ $6.33719-1.51452i$ $0.13574164$ $5.74869-1.60115i$ $5.74805-1.60159i$ $0.11547397$
3 $6.81719-2.45107i$- -- $5.91553-2.26906i$- --
4 -- -- -- $6.46920-3.21880i$ $6.41666-3.21904i$ $0.0932583$
6 0 $6.60210-0.18623i$ $6.60282-0.18610i$ $0.15997014$ $6.39998-0.23613i$ $6.40069-0.23600i$ $0.1668965$
1 $6.90042-0.69409i$ $6.90291-0.69404i$ $1.259\times 10^{-5}$ $6.60250-0.80713i$ $6.60367-0.81068i$ $2.414\times 10^{-5}$
2 $7.28536-1.41469i$ $7.28540-1.41346i$ $0.11879929$ $6.85764-1.53123i$ $6.85660-1.52978i$ $0.10977835$
3 $7.72908-2.28317i$- -- $7.21059-2.33977i$- --
4 $8.24197-3.28172i$ $8.25637-3.28542i$ $0.09481262$ $7.48045-3.19207i$ $7.50483-3.1369i$ $0.07558219$
7 0 $7.59739-0.17875i$ $7.59853-0.17859i$ $0.1454134$ $7.42853-0.21893i$ $7.42962-0.21877i$ $0.15090914$
1 $7.88016-0.66170i$ $7.88124-0.66145i$ $4.624\times 10^{-5}$ $7.64284-0.75603i$ $7.64344-0.75595i$ $6.13\times 10^{-6}$
2 $8.24579-1.34029i$ $8.24539-1.33931i$ $0.10773933$ $7.91941-1.45107i$ $7.91918-1.45030i$ $0.10293865$
3 $8.65746-2.16049i$- -- $8.22111-2.20332i$- --
4 $9.13470-3.08560i$ $9.10375-3.06571i$ $0.0848906$ $8.54449-3.11196i$ $8.54457-3.06192i$ $0.06786849$
Table 9: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=4$) under scalar perturbations for $m=0$ $7$ and axial perturbations for $m=2$ $7$. [The fundamental modes for the $n>2$ cases were published in Dutta Roy et al. (2020) (scalar) and Dutta Roy (2022) (axial)]
Scalar Perturbation Axial Perturbation
$m$ Mode Spectral Method Prony Fit Amplitude Spectral Method Prony Fit Amplitude
0 0 $0.86815-0.34320i$ $0.86815-0.34319i^{*}$- -- -- --
1 $2.04851-1.11401i$- -- -- -- --
2 $3.38964-1.91968i$ $3.39861-1.92833i^{*}$- -- -- --
1 0 $1.71264-0.22384i$ $1.71265-0.22383i^{*}$- -- -- --
1 $2.49623-0.91263i$- -- -- -- --
2 $3.64769-1.77155i$ $3.64832-1.77127i^{*}$- -- -- --
2 0 $2.66518-0.17757i$ $2.66523-0.17779i$ $0.29315965$ $1.89443-0.24761i$ $1.89445-0.24761i^{*}$ $0.2266797$
1 $3.24903-0.73077i$- -- $2.04700-0.61385i$- --
2 $4.15868-1.54933i$ $4.15833-1.54843i$ $0.20476605$ $2.78063-1.07127i$ $2.77464-1.07131i^{*}$ $0.30966731$
3 $5.28869-2.41854i$- -- -- -- --
4 $6.52829-3.26724i$ $6.53412-3.27928i$ $0.40308849$- -- --
3 0 $3.64016-0.15237i$ $3.64024-0.15233i$ $0.22714669$ $3.18363-0.28391i$ $3.18372-0.28388i$ $0.23197406$
1 $4.12259-0.61681i$ $4.14739-0.63000i$ $0.00015229$ $3.25390-0.76629i$- --
2 $4.86332-1.34256i$ $4.86269-1.34213i$ $0.18258948$ $3.65647-1.17197i$ $3.65640-1.17160i$ $0.16536179$
3 $5.81271-2.19639i$- -- $4.52375-1.72969i$- --
4 $6.94662-3.06045i$ $6.95026-3.06342i$ $0.12904273$ $5.60501-2.47686i$ $5.60406-2.47709i$ $0.15103107$
4 0 $4.62419-0.13598i$ $4.62440-0.13592i$ $0.18790256$ $4.32457-0.23666i$ $4.32477-0.23661i$ $0.20448213$
1 $5.04395-0.54241i$ $5.04408-0.54139i$ $1.859\times 10^{-5}$ $4.47738-0.78946i$- --
2 $5.67924-1.18324i$ $5.67895-1.18265i$ $0.15644501$ $4.71713-1.30051i$ $4.71834-1.29988i$ $0.10481977$
3 $6.51135-1.97082i$- -- $5.30511-1.77828i$- --
4 $7.49665-2.83332i$ $7.49472-2.83864i$ $0.11727697$ $6.23010-2.39260i$ $6.22995-2.38931i$ $0.13343803$
5 0 $5.61290-0.12418i$ $5.61332-0.12410i$ $0.16150451$ $5.39035-0.19645i$ $5.39075-0.19637i$ $0.17505116$
1 $5.98915-0.48994i$ $5.98930-0.48917i$ $7.8\times 10^{-6}$ $5.62591-0.70862i$ $5.62610-0.71015i$ $4.315\times 10^{-5}$
2 $6.55411-1.06444i$ $6.55411-1.06349i$ $0.13585498$ $5.87683-1.32855i$ $5.87679-1.32565i$ $0.09890309$
3 $7.28065-1.79073i$- -- $6.26018-1.85524i$- --
4 $8.15775-2.61273i$ $8.15569-2.60335i$ $0.10450563$ $6.96511-2.38395i$ $6.9809-2.36109i$ $0.10370454$
6 0 $6.60441-0.11513i$ $6.60513-0.11501i$ $0.14239417$ $6.42671-0.16857i$ $6.42738-0.16847i$ $0.15284811$
1 $6.94825-0.45059i$ $6.94923-0.45040i$ $0.00018714$ $6.69168-0.62009i$ $6.69217-0.62012i$ $0.00011249$
2 $7.46242-0.97399i$ $7.46683-0.96882i$ $0.1139611$ $7.01628-1.24539i$ $7.01625-1.24492i$ $0.10544664$
3 $8.11603-1.64203i$- -- $7.35083-1.88535i$- --
4 $8.90307-2.41503i$ $8.79939-2.40900i$ $0.06811125$ $7.84927-2.43946i$ $7.84363-2.43251i$ $0.08016603$
7 0 $7.59771-0.10787i$ $7.59880-0.10775i$ $0.12789715$ $7.44938-0.14872i$ $7.45045-0.14859i$ $0.13601879$
1 $7.91626-0.41968i$ $7.91704-0.41916i$ $8.235\times 10^{-5}$ $7.71959-0.55112i$ $7.72040-0.55064i$ $3.859\times 10^{-5}$
2 $8.39152-0.90296i$ $8.39220-0.90173i$ $0.10758787$ $8.08261-1.12967i$ $8.08311-1.12881i$ $0.10269501$
3 $8.99190-1.52108i$ $8.93263-1.51227i$ $0.00374277$ $8.46623-1.80756i$- --
4 $9.70803-2.24495i$ $9.69532-2.23123i$ $0.09946345$ $8.87902-2.45825i$ $8.88319-2.49385i$ $0.06576324$
Table 10: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=6$) under scalar perturbations for $m=0$ $7$ and axial perturbations for $m=2$ $7$. [The fundamental modes for the $n>2$ cases were published in Dutta Roy et al. (2020) (scalar) and Dutta Roy (2022) (axial)]
Scalar Perturbation Axial Perturbation
$m$ Mode Spectral Method Prony Fit Amplitude Spectral Method Prony Fit Amplitude
0 0 $0.87997-0.32519i$ $0.87998-0.32518i^{*}$- -- -- --
1 $2.14466-0.97491i$- -- -- -- --
2 $3.57108-1.59477i$ $3.57054-1.59493i^{*}$- -- -- --
1 0 $1.72193-0.20105i$ $1.72194-0.20104i^{*}$- -- -- --
1 $2.57737-0.79070i$- -- -- -- --
2 $3.82365-1.47070i$ $3.82587-1.46941i^{*}$- -- -- --
2 0 $2.67187-0.15181i$ $2.67251-0.15238i$ $0.29083766$ $1.91003-0.23457i$ $1.91005-0.23456i^{*}$ $0.21776703$
1 $3.30697-0.61908i$- -- $2.08571-0.56896i$- --
2 $4.31495-1.28267i$ $4.31427-1.28190i$ $0.23210154$ $2.87446-0.93747i$ $2.87405-0.93741i^{*}$ $0.29184553$
3 $5.55107-1.96109i$- -- $4.04950-1.50405i$- --
4 $6.88381-2.61925i$ $6.88773-2.61918i$ $0.25385346$ $5.37019-2.11453i$ $5.36933-2.11477i^{*}$ $0.1670686$
3 0 $3.64456-0.12512i$ $3.64471-0.12510i$ $0.22202515$ $3.19594-0.26645i$ $3.19603-0.26648i$ $0.22253281$
1 $4.16276-0.50814i$ $4.17354-0.50844i$ $0.00037032$ $3.28608-0.69780i$- --
2 $4.98966-1.10188i$ $4.99007-1.10179i$ $0.18332356$ $3.73814-1.01798i$ $3.73866-1.01791i$ $0.17541431$
3 $6.06575-1.77885i$- -- $4.69073-1.44514i$- --
4 $7.29017-2.45688i$ $7.29374-2.46700i$ $0.13255104$ $5.86562-2.02004i$ $5.8667-2.02183i$ $0.15024292$
4 0 $4.62667-0.10807i$ $4.62688-0.10801i$ $0.18186996$ $4.33741-0.21472i$ $4.33761-0.21467i$ $0.19968836$
1 $5.07164-0.43530i$ $5.07146-0.43473i$ $5.0481\times 10^{-6}$ $4.49808-0.72204i$- --
2 $5.77715-0.95771i$ $5.77702-0.95720i$ $0.15714858$ $4.76792-1.13746i$ $4.76742-1.13632i$ $0.10174903$
3 $6.71074-1.59843i$- -- $5.44241-1.46518i$- --
4 $7.81676-2.27620i$ $7.81867-2.27681i$ $0.11845874$ $6.47562-1.94993i$ $6.47398-1.95101i$ $0.13691266$
5 0 $5.61380-0.09608i$ $5.61421-0.09601i$ $0.15498730$ $5.40063-0.17075i$ $5.40103-0.17067i$ $0.16992681$
1 $6.00767-0.38426i$ $6.00786-0.38394i$ $1.7893\times 10^{-5}$ $5.65481-0.64034i$- --
2 $6.62899-0.84805i$ $6.62886-0.84759i$ $0.13681384$ $5.90429-1.18229i$ $5.90559-1.17820i$ $0.08731672$
3 $7.45181-1.43958i$ $7.45249-1.39775i$ $0.00057971$ $6.34149-1.56146i$- --
4 $8.44520-2.09680i$ $8.44251-2.10081i$ $0.11003456$ $7.17592-1.93339i$ $7.18692-1.93325i$ $0.11707173$
6 0 $6.60399-0.08711i$ $6.60469-0.08702i$ $0.13560445$ $6.43410-0.14111i$ $6.43478-0.14102i$ $0.14713892$
1 $6.95973-0.34643i$ $6.96018-0.34614i$ $1.3489\times 10^{-5}$ $6.72107-0.54020i$ $6.72167-0.53992i$ $8.009\times 10^{-5}$
2 $7.51913-0.76397i$ $7.51916-0.76313i$ $0.12066340$ $7.05523-1.11725i$ $7.05465-1.11561i$ $0.09874688$
3 $8.25731-1.30704i$ $8.23586-1.33897i$ $0.00457058$ $7.39302-1.62858i$- --
4 $9.15431-1.93145i$ $9.15519-1.93202i$ $0.10141119$ $7.98533-1.99226i$ $7.98248-1.98655i$ $0.08462117$
7 0 $7.59621-0.08009i$ $7.59732-0.07999i$ $0.12093846$ $7.45434-0.12059i$ $7.45540-0.12047i$ $0.12992236$
1 $7.92220-0.31715i$ $7.92296-0.31677i$ $3.2848\times 10^{-5}$ $7.74342-0.46352i$ $7.74431-0.46321i$ $1.434\times 10^{-5}$
2 $8.43378-0.69798i$ $8.43393-0.69726i$ $0.10853332$ $8.13502-0.98974i$ $8.13560-0.98861i$ $0.10233494$
3 $9.10647-1.19796i$ $9.10056-1.19398i$ $0.00094986$ $8.51244-1.59506i$- --
4 $9.92416-1.78541i$ $9.92248-1.77980i$ $0.09058615$ $8.94988-2.06737i$ $8.95234-2.06370i$ $0.06334415$
Table 11: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=8$) under scalar perturbations for $m=0$ $7$ and axial perturbations for $m=2$ $7$. [The fundamental modes for the $n>2$ cases were published in Dutta Roy et al. (2020) (scalar) and Dutta Roy (2022) (axial)]

### v2

@@
 Further, we comment on how (a) the overtones may influence the time-domain profile in a perturbation and (b) in what sense, the use of overtones may help in distinguishing between different geometries within the family.
-Finally, we show how an effect somewhat similar to the so-called ‘environment induced redshift’ of QNMs, introduced recently (arXiv:2412.18651), may be obtained for spacetimes with non-constant $g_{00}$, via an appropriate tuning of available metric parameters which systematically modify the shapes of the effective potentials arising in the perturbation equations.
+Finally, we show how an effect somewhat similar to the so-called ‘environment induced redshift’ of QNMs, introduced recently (Phys.
+Rev.
+D 111, 064026 (2025)), may be obtained for spacetimes with non-constant $g_{00}$, via an appropriate tuning of available metric parameters which systematically modify the shapes of the effective potentials arising in the perturbation equations.
 I Introduction and summary
@@
 (2024); Radhakrishnan et al.
-(2024); Dehghani and Dayyani (2009); Sadatian and Hosseini (2024) seemed to be a sort of way out.
+(2024); Dehghani and Dayyani (2009); Sadatian and Hosseini (2024); Loewer et al.
+(2024); Battista et al.
+(2024) seemed to be a sort of way out.
 It was noticed that in a modified theory of gravity, the purely geometric violation of the timelike or null convergence condition does not necessarily imply a violation of the weak, null or dominant energy conditions.
@@
 (2001); Izmailov et al.
-(2019) was one of the first among the observationally relevant works pursued by researchers.
+(2019); Mitra and Ghosh (2024); Tsukamoto and Harada (2017) was one of the first among the observationally relevant works pursued by researchers.
 It still continues as an important possible signature.
-Over the last decade or so, astronomy and astrophysics has reached a different level of observational relevance through the emergence of a large amount of data and results in gravitational wave astronomy (LIGO-VIRGO-KAGRA collaboration Abbott et al.
-(2023)) and infrared imaging of astrophysical objects (the Event Horizon Telescope (EHT) Akiyama et al.
-(2019, 2022)).
+Over the last decade or so, astronomy and astrophysics has reached a different level of observational relevance through the emergence of a large amount of data and results in gravitational wave astronomy (LIGO-VIRGO-KAGRA collaboration Abbott and others (2023)) and infrared imaging of astrophysical objects (the Event Horizon Telescope (EHT) Akiyama and others (2019, 2022)).
 The results obtained tally well with General Relativity predictions, leaving little room for modified gravity.
@@
 (2009); Cardoso and Lemos (2001); Nollert (1999); Konoplya (2002, 2003); Rincón and Santos (2020); Blázquez-Salcedo et al.
-(2018a) and wormhole Bueno et al.
+(2018b) and wormhole Bueno et al.
 (2018); Konoplya and Molina (2005); Kim (2008); Alfaro et al.
@@
 (2025); González et al.
-(2022) computed QNM overtones for various wormhole geometries.
+(2022); Konoplya and Stashko (2025) computed QNM overtones for various wormhole geometries.
 Our aim here is to first obtain these overtones using a numerical scheme which allows us to calculate as many overtones as possible.
@@
 (2025); Liu et al.
-(2025).
+(2025); Chen and Kotlařík (2023).
 We show here, for the QNMs and overtones found by us for the wormhole family, that there is indeed a ‘redshift’ when we modify the so-called effective potential by tuning a metric parameter appropriately.
-In this context, we also provide additional examples of a similar redshift in the simplest of examples–eg.
-the square barrier or the Poschl-Teller potential.
 Our article, which is largely a sequel to Dutta Roy et al.
@@
 In the next section, we provide a summary of the wormhole family we study in this article.
-Section III quickly outlines the methods adopted in finding the QNMs–more specifically, which method is better suited for finding overtones is emphasised and discussed in more detail here.
-In Section IV, we summarise the numerical results obtained with explicit plots.
-Section V discusses the inferences which may be drawn from our studies on the QNMs and overtones.
-The above-mentioned ‘redshift’ effect is discussed extensively in Section VI.
-Finally, in Section VII, we conclude with remarks on possible future work.
-The appendix contains the detailed numerical values of the QNMs, thereby supporting the plots shown in Section IV.
+In Section III, we summarise the numerical results obtained with explicit plots.
+Section IV discusses the inferences which may be drawn from our studies on the QNMs and overtones.
+The above-mentioned ‘redshift’ effect is discussed extensively in Section V.
+Finally, in Section VI, we conclude with remarks on possible future work.
+Appendix A quickly outlines the methods adopted in finding the QNMs–more specifically, which method is better suited for finding overtones is emphasised and discussed in more detail here.
+The appendix B contains the detailed numerical values of the QNMs, thereby supporting the plots shown in Section III.
 II A brief review on the family of wormhole geometries
@@
 $ds^{2}=-dt^{2}+\frac{dr^{2}}{1-\frac{b_{0}^{2}}{r^{2}}}+r^{2}d\theta^{2}+r^{2}sin^{2}\theta d\varphi^{2}$ (2)
-One may also rewrite the line element using the proper radial distance, i.e $r^{2}=b_{0}^{2}+l^{2}$ (where $-\infty\leq l\leq\infty$). In this representation, we have
+One may also rewrite the line element using the proper radial distance, i.e $r^{2}=b_{0}^{2}+l^{2}$ (where $\infty\leq l\leq\infty$). In this representation, we have
 $ds^{2}=-dt^{2}+dl^{2}+\left(l^{2}+b_{0}^{2}\right)d\Omega_{2}^{2}$ (3)
 We may visualise the above line element as representing two asymptotically flat regions ($l\rightarrow\pm\infty$) connected by a throat/bridge of radius $b_{0}$.
-Note that the radial coordinate $r$ runs from $b_{0}$ to $\infty$ whereas $l$ runs from $\infty$- ∞ to $\infty$ with $l=0$ corresponding to $r=b_{0}$.
+Note that the radial coordinate $r$ runs from $b_{0}$ to $\infty$ whereas $l$ runs from -$\infty$ to $\infty$ with $l=0$ corresponding to $r=b_{0}$.
 We will now present generalisations of the above-mentioned simple geometry.
@@
 $ds^{2}=-dt^{2}+dl^{2}+(l^{n}+b_{0}^{n})^{\frac{2}{n}}(d\theta^{2}+sin^{2}\theta d\varphi^{2})$ (5)
-Here, $l$ ([,])absent($\in$[$\infty$,$\infty$])( ∈ [ - ∞ , ∞ ] ) is the tortoise coordinate
+Here, $l$ $(\in[-\infty,\infty])$ is the tortoise coordinate
 $dl^{2}=\frac{dr^{2}}{1-\frac{b(r)}{r}}\implies r(l)=(l^{n}+b_{0}^{n})^{\frac{1}{n}}$
@@
 However, the authors of Bronnikov (1973), Ellis (1973) (see also Blázquez-Salcedo et al.
-(2018b)) have proposed a metric for a massive, static, Bronnikov–Ellis wormhole for which the line element is of the following form:
+(2018a)) have proposed a metric for a massive, static, Bronnikov–Ellis wormhole for which the line element is of the following form:
 $ds^{2}=-e^{f(l)}dt^{2}+e^{-f(l)}[dl^{2}+(l^{2}+b_{0}^{2})d\Omega^{2}]$ (6)
 with
-$f(l)=\frac{C}{b_{0}}\left(arctan\left(\frac{l}{b_{0}}\right)-\frac{\pi}{2}\right)$
+$f(l)=\frac{C}{b_{0}}\left(\arctan\left(\frac{l}{b_{0}}\right)-\frac{\pi}{2}\right)$
 In the asymptotic limit $C=2M$, where $M$ is the mass extracted in the asymptotically flat region.
@@
 For $M=0$, we recover the symmetric generalised Bronnikov–Ellis wormhole, while any non-zero value of $M$ will introduce asymmetry in the wormhole shape.
-The QNMs for the $n=2$, $M\neq 0$ geometry will be discussed in Section VI.3.
+The QNMs for the $n=2$, $M\neq 0$ geometry will be discussed in Section V.1.
 The energy conditions in general relativity serve as criteria to ensure the physical viability of solutions to Einstein’s field equations.
@@
 In all the images, the x-axis represents the proper coordinate $l$.
-However, in the case of $M\neq 0$ and $n>2$, one can find that all three conditions are satisfied in certain regions if the $|M|$ exceeds a specific value.
+However, in the case of $M\neq 0$ and $n>2$, one can find that all three conditions are satisfied in certain regions if the $|M|$ exceeds a specific value (Fig.
+1).
 For different geometries (different values of $n$), the critical value of $M$ and the acceptable regions (corresponding to a fixed value of $M$) are different.
 II.3 Bronnikov–Ellis geometry with a chosen redshift function:
-It is also possible to choose a suitable redshift function ($-g_{00}$ never zero) in an ad-hoc way while retaining the Bronnikov–Ellis $n=2$ character unaltered in the spacelike slice.
+It is also possible to choose a suitable redshift function ($g_{00}$ never zero) in an ad-hoc way while retaining the Bronnikov–Ellis $n=2$ character unaltered in the spacelike slice.
 The ensuing consequences for required matter and energy conditions can thereafter be looked at.
@@
 $\rho(r)=-\frac{b_{0}^{2}}{r^{4}},\hskip 42.67912pt\tau(r)=\frac{2Mr-b_{0}^{2}}{r^{3}(r-2M)},\hskip 42.67912ptp(r)=\frac{(M-r)\left(b_{0}^{2}(M-r)+Mr^{2}\right)}{r^{4}(r-2M)^{2}}$ (14)
+Figure 2: For Bronnikov–Ellis wormhole with redshift function: [left] $\rho+\tau$; [right] $\rho+p$. Blue line denotes $M=0.2b_{0}$ and red line represents $M=0.4b_{0}$. (here $b_{0}=1$)
 Note that the energy density is not affected by $M$, though the pressures change.
@@
 The Null Energy Condition is also violated with complementary behaviour for the $\rho+\tau$ and $\rho+p$ inequalities w.r.t.
-increase in $M$ value.
+increase in $M$ value (Fig.
+2).
 As $M$ increases the region of violation (near $r=b_{0}$) of the $\rho+\tau$ inequality shrinks whereas the region of violation (away from $r=b_{0}$) of the $\rho+p$ inequality becomes larger.
@@
 2), in order to show this feature in the NEC violations.
-Putting $M=0$, we get back the original Bronnikov–Ellis wormhole. We will discuss the effective potential and the $\omega_{QNM}$ for this spacetime in a later section (VI.4).
-Figure 2: For Bronnikov–Ellis wormhole with redshift function: [left] $\rho+\tau$; [right] $\rho+p$. Blue line denotes $M=0.2b_{0}$ and red line represents $M=0.4b_{0}$. (here $b_{0}=1$)
+Putting $M=0$, we get back the original Bronnikov–Ellis wormhole. We will discuss the effective potential and the $\omega_{QNM}$ for this spacetime in a later section (V.2).
 II.4 Effective potentials for scalar and axial perturbation
@@
 For this, we may write the scalar field as $\Phi+\phi$.
+Here $\phi$ is the perturbation on the phantom scalar field and it satisfies
 $\Box\Phi=\frac{1}{\sqrt{-g}}\partial_{\mu}(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\Phi)=0$ (15)
@@
 (15), one gets the Schrodinger-like radial equation
-$\frac{d^{2}\psi}{dl_{*}^{2}}+(\omega^{2}-V_{sc}(l))\psi=0\hskip 56.9055ptwhere\hskip 28.45274pt\frac{dl_{*}}{dl}=e^{-f}$ (16)
+$\frac{d^{2}\psi}{dl_{*}^{2}}+(\omega^{2}-V_{sc}(l))\psi=0\hskip 56.9055pt\text{where}\hskip 28.45274pt\frac{dl_{*}}{dl}=e^{-f}$ (16)
 Here, $V_{sc}$ is the effective potential (scalar perturbation) for the massive generalised Bronnikov–Ellis wormhole (Eq. (7)), given by
@@
 The shift and asymmetry are also noticeable for $n>2$ cases, where the height of the peaks on the positive side of $l_{*}$ exceeds that of the negative side (if $M>0$).
-III Methods used to find the QNMs
+III Results: quasinormal spectra (fundamental and overtones for the ultra-static case)
+The problem of computing quasi-normal modes for the Bronnikov–Ellis wormhole has been tackled in many previous works.
+The authors of Kim (2008); Blázquez-Salcedo et al.
+(2018a) and Dutta Roy et al.
+(2020); Dutta Roy (2022) have presented the fundamental modes for the original Bronnikov–Ellis ($n=2$ case) and the generalised Bronnikov–Ellis wormholes, respectively.
+However, as pointed out in Sago et al.
+(2021); Berti et al.
+(2007); Giesler et al.
+(2019); Nee et al.
+(2023), the fundamental modes may not represent the full picture, as the overtones play a crucial role in fitting the ringdown (RD) waveform.
+The inclusion of overtones (at least, the first few overtones) can help us achieve a better fit of the RD profile at earlier times when we expect a better signal-to-noise ratio (SNR).
+Overtones can also be instrumental in providing better estimates of the parameters of the underlying geometry.
+Though the overtones corresponding to the $n=2$ geometry were computed in Alfaro et al.
+(2024); Batic and Dutykh (2024), we compute the full QNM spectra for the generalised Bronnikov–Ellis spacetimes (even $n>2$ cases).
+We note many key differences between the $n=2$ and the $n>2$ cases.
+[…]
+While computing the $\omega_{QNM}$, we have used a frequency-domain method (spectral method (SM)) in conjunction with a time-domain method (Prony fit) (see appendix A).
+While the frequency-domain method is indispensable for extracting all the modes (up to a certain overtone number), the time-domain provides us with some valuable insight into the relative amplitudes of the overtones, thus pointing to the fact that sometimes even higher-order overtones (in our case: even modes) can have greater importance than the lower-order (in our case: odd modes) ones.
+The results from the two methods are in good agreement with each other.
+However, we could not compute the same number of overtones for every geometry.
+For lower values of $n$ and angular momentum parameter $m$, we could get very few overtones.
+In the series of figures below, we demonstrate the effective potentials and QNMs for scalar and axial gravitational perturbations.
+Figs.
+3-6 are for $n=2$, Figs.
+7-10 for $n=4$, Figs.
+11-14 are for $n=6$ and Figs.
+15-18 for $n=8$.
+In the following section, we discuss the results here and try to arrive at specific conclusions on the nature of the QNMs and their link with the different geometries.
+QNMs for n = 2 geometry:
+Figure 3: Effective potential (Scalar) for Bronnikov–Ellis wormhole with $n=2$.
+Figure 4: Effective potential (Axial) for Bronnikov–Ellis wormhole with $n=2$.
+Figure 5: QNMs for n = 2 (Scalar) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 8)
+Figure 6: QNMs for n = 2 (Axial) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 8)
+QNMs for n = 4 geometry:
+Figure 7: Effective potential (Scalar) for generalised Bronnikov–Ellis wormhole with $n=4$.
+Figure 8: Effective potential (Axial) for generalised Bronnikov–Ellis wormhole with $n=4$.
+Figure 9: QNMs for n = 4 (Scalar) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 9)
+Figure 10: QNMs for n = 4 (Axial) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 9)
+QNMs for n = 6 geometry:
+Figure 11: Effective potential (Scalar) for generalised Bronnikov–Ellis wormhole with $n=6$.
+Figure 12: Effective potential (Axial) for generalised Bronnikov–Ellis wormhole with $n=6$.
+Figure 13: QNMs for n = 6 (Scalar) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 10)
+Figure 14: QNMs for n = 6 (Axial) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 10)
+QNMs for n = 8 geometry:
+Figure 15: Effective potential (Scalar) for generalised Bronnikov–Ellis wormhole with $n=8$.
+Figure 16: Effective potential (Axial) for generalised Bronnikov–Ellis wormhole with $n=8$.
+Figure 17: QNMs for n = 8 (Scalar) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 11)
+Figure 18: QNMs for n = 8 (Axial) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 11)
+IV Inferences drawn from the QNM spectra: emphasis on overtones
+[…]
+[…]
+[…]
+The imaginary parts of the quasinormal spectra are also an important feature distinguishing the geometries.
+In the scalar perturbation case, for a fixed angular momentum parameter value, the greater the $n$ value, the smaller the magnitude of the imaginary parts of the QNMs.
+The separation between the $|Im(\omega)|$ also keeps decreasing (See Figs.
+6, 10, 14, 18).
+This implies that for wormholes with larger values of $n$, the fundamental modes and overtones have a lower decay rate and, therefore, are less stable.
+As we keep on increasing $m$, for the axial perturbation case, the absolute value of the imaginary parts of the fundamental modes first increase and then decrease (See Figs.
+6, 10, 14, 18).
+Variation Scalar Axial
+The rate of increase The rate of increase
+[…]
+(for $n>2$) number gets expedited. number gets expedited.
+[…]
+Increase m: $|Im(\omega)|$ decreases $|Im(\omega)|$ (fundamental mode) first
+increases and then decreases.
+The rate of increase of The rate of increase of
+[…]
+(for $n>2$) with overtone number. certain overtone number.
+$|Im(\omega)|$ increases. $|Im(\omega)|$ increases.
+Table 1: Key features of the QNM spectra for massless generalised Bronnikov-Ellis wormhole.
+[…]
+Fixing the value of $n$, if we increase the angular momentum parameter ($m$), the height of the effective potential increases, which is related to increasing the real part of the QNMs.
+The double peak feature also becomes less prominent for the scalar perturbation case (See Figs.
+8, 12, 16).
+This may explain the decrease of $|Im(\omega)|$ with $m$ in the quasi-normal spectra.
+For axial perturbations in $n>2$ geometries, the transition of the potential from triple barrier to step-like and then to a single barrier, with increasing m (See Figs.
+8, 12, 16), may cause the initial increase and subsequent decrease of $|Im(\omega)|$ for the fundamental modes with m.
+IV.1 Approximate Analytic Fit
+To further advance our study of the QNM spectra and their dependence on different parameters such as the geometry ($n$), angular momentum parameter ($m$), overtone order ($p$) and the throat ($b_{0}$), we can fit our numerical results (for scalar perturbations) to analytical models.
+Here, the throat radius is simply a scaling parameter, and the quasi-normal frequencies are inversely proportional to $b_{0}$.
+However, the other parameters do not follow such simple rules.
+To get the fits corresponding to these parameters, we have used the NonLinearModelFit function of Mathematica 14.2.
+For all the fits obtained in this section, we have used the results from the spectral method.
+[…]
+Figure 19: QNMs for n = 2 (Scalar Perturbation) case.
+Here, we observe that the real parts of the quasi-normal frequencies decrease with overtone mode $p$ but increase with angular momentum parameter $m$ (Fig. 19).
+Figure 20: [Left] QNMs for n = 4; [Right] QNMs for n = 8.
+[…]
+Figure 21: Fit of Im($\omega$) with n and Re($\omega$): [left] fundamental modes; [middle] 1st overtones; [right] 2nd overtones.
+[…]
+IV.2 Detectable Modes
+Since we have now obtained the quasinormal modes including some overtones, it is important to ask whether or not these modes are detectable.
+If they are within observable limits, this must be for certain choices of the parameters involved.
+To figure this out, we will need to determine the numerical values for the quasinormal frequencies ($\nu$) and decay times ($\tau$), which can be found from
+$\nu=\frac{c^{3}}{2\pi b_{0}GM_{\odot}}Re(\omega)\hskip 5.69046ptHz\hskip 56.9055pt\tau=\frac{2\pi b_{0}GM_{\odot}}{c^{3}}\frac{1}{Im(\omega)}\hskip 5.69046ptseconds$
+Advanced LIGO is sensitive to the frequency range of $10\hskip 2.84544ptHz$ to $10\hskip 2.84544ptkHz$.
+For the generalised Bronnikov-Ellis geometries, we have calculated the QNFs, and all of them can be within the aLIGO frequency range when the throat radius is roughly between $50M_{\odot}$ and $1000M_{\odot}$.
+For now, we choose $b_{0}=200$ (in units of solar mass) and present the frequencies and damping times for a few cases, as an illustration.
+In Table 2 below, we show the values for $n=2$ and $m=5$ (scalar perturbation).
+Mode Frequency (Hz) Damping Time (s) Amplitude
+0 889.5 0.01235 0.2396
+[…]
+2 845.8 0.00244 0.1219
+Table 2: Frequency and Damping time for $n=2$ and $m=5$. These values are calculated from the Prony fit (details in Appendix B).
+Figure 22: Log plot of time-domain profile for $\psi$ for $n=2$ and $m=5$.
+Figure 23: Time-domain waveform for $n=2$ and $m=5$.
+Frequencies and damping times for n = 8 and m = 5 case (scalar perturbation) are mentioned in Table 3
+Mode Frequency (Hz) Damping Time (s) Amplitude
+0 904.2 0.06467 0.1550
+[…]
+2 1067.7 0.00733 0.1368
+3 1200.3 0.00444 0.0006
+4 1359.8 0.00296 0.1100
+Table 3: Frequency and Damping time for $n=8$ and $m=5$. These values are calculated from the prony fit (details mentioned in Appendix B).
+Figure 24: Log plot time-domain profile for $\psi$ for $n=8$ and $m=5$.
+Figure 25: Time-domain waveform for $n=8$ and $m=5$.
+In the ringdown waveforms, we have represented the actual time-domain (obtained using the procedures mentioned in A.1) wave as the blue points, the fundamental mode as green lines and the resultant wave of the fundamental and overtones ($\omega$ and amplitudes computed using the Prony fit) as the magenta lines (Figs.
+23,25).
+For both cases, the magenta lines better fit the time domain than only the fundamental mode.
+However, the difference between the green and magenta lines is only noticeable for the first few cycles, as the overtones have higher imaginary parts and decay faster.
+While the inclusion of overtones makes little improvement in the fitting for the $n=2$ case, the effect is more significant for the $n=8$ case.
+This is because for the $n=8$ case, we could find more overtones, and these overtones have amplitudes comparable to the fundamental mode, whereas, in the $n=2$ case, the amplitudes of the overtones are significantly lower than the fundamental mode (here the fundamental mode fits well from the second cycle) .
+For $n=4,6$ cases, we computed 232-3 overtones and their corresponding amplitudes using the Prony fit, and similar to the $n=8$ case, we observe that inclusion of the overtones significantly improves the fit at early times.
+For all the geometries (i.e.
+different $n$), the more overtones we could get (up to $p=4$), the better we were able to model the time-domain signal, indicating that obtaining more overtones may result in a better fit of the time-domain profile.
+However, we need to be cautious not to overfit our results by incorporating an excessive number of overtones.
+Figure 26: Amplitude calculated for $n=2$ and $m=5$ with different initial time.
+Figure 27: Amplitude calculated for $n=8$ and $m=5$ with different initial time.
+One would expect the amplitudes of the modes to depend on our initial value of $\psi$ but not on which part of the time domain we are fitting.
+However, for the generalised Bronnikov–Ellis wormhole, only the amplitudes associated with the fundamental modes and 2nd overtones (to some extent) are time-independent (Figs.
+27,27).
+For almost all the other $\omega_{QNM}$, if we change the initial time of the fitting, the corresponding amplitudes remain stable up to some fitting domain, but beyond it, they begin to vary.
+Thus, we again fit the ringdown profiles with the modes with a stable amplitude
+Figure 28: Time-domain profile for all computed overtones and the modes with stable amplitude for $n=2$ and $m=5$.
+Figure 29: Time-domain profile for all computed overtones and the modes with stable amplitude for $n=8$ and $m=5$.
+In both plots (Figs.
+29, 29), the black line represents the contribution of the fundamental and 2nd overtone (the modes with stable amplitude), and the magenta line has the contribution of all the modes from tables 2 and 3.
+For the $n=2$ case, the black and the magenta lines overlap (this is expected as we filtered out the 1st overtone with a very small amplitude).
+In the $n=8$ case, the two lines deviate at a very early time.
+The inclusion of all the modes gives a slightly better fit as the modes that were filtered out are also stable in the initial times, and their inclusion does not change the late-time behaviour, as these modes are damped quickly.
+Thus, we conclude that the inclusion of overtones can indeed help us achieve a better fit of the time-domain profile at earlier times.
+V Redshift
+In the previous sections, we have computed the quasinormal modes for the massless generalised Bronnikov–Ellis wormhole.
+However, if the geometry under study is surrounded by a dark matter halo or any generic matter distribution, the QNMs calculated for the isolated compact object may differ from the observed values, as the environment may affect the generation of GW as well as their propagation.
+Such environments cause the spectrum to be red-shifted as documented in Barausse et al.
+(2014); Pezzella et al.
+(2025); Liu et al.
+(2025); Chen and Kotlařík (2023) for the black hole spacetime embedded in a halo.
+In the case of the Bronnikov–Ellis wormhole, similar redshifts in the fundamental mode can be found, if one considers the wormhole with non-zero ADM mass, as noted by the authors of Blázquez-Salcedo et al.
+(2018a).
+We will first find the QNMs (fundamental and overtones) for the $M\neq 0$ case in the $n=2$ Bronnikov–Ellis geometry.
+Thereafter, in our second example, we will find the QNMs for a modified $n=2$ geometry with a new redshift function having an extra parameter $M$ Alfaro et al.
+(2024).
+In these two scenarios, using the QNMs obtained, we then verify if we have a parameter-controlled redshift, similar to the environment-induced redshift mentioned above.
+The modelling of such red-shifts is crucial as the red-shifted modes may appear to be QNMs of the original isolated geometry with a larger throat radius (as $\omega$ varies inversely with $b_{0}$).
+V.1 Bronnikov–Ellis geometry with asymptotic mass:
+Let us now see if redshift effects arise for the QNMs of perturbations of Bronnikov–Ellis geometries.
+The ultrastatic Bronnikov–Ellis geometry is not of use since it has only one parameter $b_{0}$ (throat radius) which essentially appears as an overall scale factor in the QNMs.
+Thus, we either modify the shape function with extra parameters, in which case it is no longer Bronnikov–Ellis or, alternatively, we may bring in the redshift function which will have an extra parameter.
+We choose to work with this latter modification.
+The well-known, original line element for a Bronnikov–Ellis geometry with a redshift function is given as,
+$ds^{2}=-e^{f(l)}dt^{2}+e^{-f(l)}[dl^{2}+(l^{2}+b_{0}^{2})d\Omega_{2}^{2}]$ (20)
+where, $f(l)=\frac{C}{b_{0}}\left(\arctan\left(\frac{l}{b_{0}}\right)-\frac{\pi}{2}\right)$.
+Note that due to the presence of $f(l)$, the energy density as well as the pressures change, as discussed in Section II.2.
+We will work with this line element in order to demonstrate the redshift effect in the QNMs.
+The ADM mass ($M$) of this spacetime is related to the metric parameter $C$ as $C=2M$.
+One can plot the effective potentials for both scalar and axial perturbations numerically and observe that the symmetry in the potential barriers is lost with the introduction of non-zero asymptotic mass ($M$).
+The QNM spectra computed using the Prony fit (from $t=0$ to $t=40$ with $h=1/40$ using Mathematica 14.2 for $M\neq 0$, $b_{0}=1$), for this geometry, are presented in the tables 4,5 .
+However, in the figures we have presented some results from the spectral method.
+The effective potential (scalar perturbation) for this metric can be expressed as Eq.
+(17).
+Here, we have plotted the effective potentials for the $n=2$ case ($m=5$ in Fig.
+31 and $m=7$ in Fig.
+31) and obtained single-peaked potentials with the maxima of the potential shifting to the right as we increase the value of $C$ (=2M=2M).
+The QNMs corresponding to $m=5$ and $m=7$ are plotted in Figs.
+33 and 33 respectively.
+Figure 30: Effective potentials (Scalar perturbation) for different values of M. (n = 2 and m = 5).
+Figure 31: Effective potentials (Scalar perturbation) for different values of M. (n = 2 and m = 7).
+Figure 32: QNMs (Scalar) for different values of M. (n = 2 and m = 5). The results of SM and Prony fit are represented by circles and plus signs, respectively.
+Figure 33: QNMs (Scalar) for different values of M. (n = 2 and m = 7). The results of SM and Prony fit are represented by circles and plus signs, respectively.
+$m$ Mode $M=0b_{0}$ $M=0.025b_{0}$ $M=0.05b_{0}$ $M=0.075b_{0}$ $M=0.1b_{0}$
+3 0 3.534430.50686i3.53443-0.50686i 3.271660.46911i3.27166-0.46911i 3.035760.43526i3.03576-0.43526i 2.823830.40465i2.82383-0.40465i 2.632550.37783i2.63255-0.37783i
+5 0 5.522820.50282i5.52282-0.50282i 5.112430.46532i5.11243-0.46532i 4.743780.43185i4.74378-0.43185i 4.412590.40161i4.41259-0.40161i 4.114270.37699i4.11427-0.37699i
+1 5.417961.52258i5.41796-1.52258i 5.043811.38243i5.04381-1.38243i 4.659621.30864i4.65962-1.30864i 4.338111.21722i4.33811-1.21722i 4.086991.09940i4.08699-1.09940i
+2 5.251232.54327i5.25123-2.54327i 4.858552.36301i4.85855-2.36301i 4.442922.19880i4.44292-2.19880i 4.148092.04224i4.14809-2.04224i -
+7 0 7.519340.50106i7.51934-0.50106i 6.959850.46386i6.95985-0.46386i 6.458100.43037i6.45810-0.43037i 6.007110.40048i6.00711-0.40048i 5.601080.37349i5.60108-0.37349i
+1 7.455761.50817i7.45576-1.50817i 6.913981.38873i6.91398-1.38873i 6.405561.30364i6.40556-1.30364i 5.953961.20411i5.95396-1.20411i 5.550571.11919i5.55057-1.11919i
+2 7.319852.52412i7.31985-2.52412i 6.791142.33889i6.79114-2.33889i 6.284852.12915i6.28485-2.12915i 5.845092.02687i5.84509-2.02687i 5.468271.89265i5.46827-1.89265i
+9 0 9.515610.50062i9.51561-0.50062i 8.810080.46296i8.81008-0.46296i 8.174730.42973i8.17473-0.42973i 7.603780.39983i7.60378-0.39983i 7.090230.37301i7.09023-0.37301i
+1 9.433511.50092i9.43351-1.50092i 8.762701.41073i8.76270-1.41073i 8.131211.29187i8.13121-1.29187i 7.559411.20308i7.55941-1.20308i 7.045881.12225i7.04588-1.12225i
+2 9.357642.51354i9.35764-2.51354i 8.646602.31498i8.64660-2.31498i 8.029582.14977i8.02958-2.14977i 7.507362.00358i7.50736-2.00358i 6.984781.87115i6.98478-1.87115i
+Table 4: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Scalar perturbation) with $M=0,0.025b_{0},0.05b_{0},0.075b_{0},0.1b_{0}$.
+(The results are obtained using Prony fit.) [The fundamental modes were computed in Blázquez-Salcedo et al.
+(2018a).
+We extended the analysis to include overtones.]
+The effective potential for axial perturbation has the form Eq.
+(19) and are plotted below.
+With increasing M ($\frac{C}{2}$), the height of the barrier reduces and the potentials become more asymmetric (See Figs.
+35, 35).
+The figures 37 and 37 display the QNM spectra for axial perturbation in $m=5$ and $m=7$ cases.
+Figure 34: Effective potentials (Axial Perturbation) for different values of M. (n = 2 and m = 5).
+Figure 35: Effective potentials (Axial Perturbation) for different values of M. (n = 2 and m = 7).
+Figure 36: QNMs (Axial Perturbation) for different values of M. (n = 2 and m = 5). The results of SM and Prony fit are represented by circles and plus signs, respectively.
+Figure 37: QNMs (Axial Perturbation) for different values of M. (n = 2 and m = 7). The results of SM and Prony fit are represented by circles and plus signs, respectively.
+[…]
+$m$ Mode $M=0b_{0}$ $M=0.025b_{0}$ $M=0.05b_{0}$ $M=0.075b_{0}$ $M=0.1b_{0}$
+5 0 5.151760.44366i5.15176-0.44366i 4.772360.43237i4.77236-0.43237i 4.428820.40144i4.42882-0.40144i 4.120300.37367i4.12030-0.37367i 3.843890.34924i3.84389-0.34924i
+1 5.066861.39798i5.06686-1.39798i 4.731551.33422i4.73155-1.33422i 4.360341.21916i4.36034-1.21916i 4.060041.12686i4.06004-1.12686i 3.764550.99294i3.76455-0.99294i
+2 4.919582.39447i4.91958-2.39447i 4.555092.21569i4.55509-2.21569i - - -
+7 0 7.250730.48257i7.25073-0.48257i 6.711070.44682i6.71107-0.44682i 6.227500.41477i6.22750-0.41477i 5.793550.38607i5.79355-0.38607i 5.402970.36046i5.40297-0.36046i
+1 7.193871.45787i7.19387-1.45787i 6.684081.34773i6.68408-1.34773i 6.173471.25075i6.17347-1.25075i 5.742081.16371i5.74208-1.16371i 5.356771.08268i5.35677-1.08268i
+2 7.059932.43004i7.05993-2.43004i 6.529632.26610i6.52963-2.26610i 6.048652.09636i6.04865-2.09636i 5.636291.94025i5.63629-1.94025i 5.245201.82708i5.24520-1.82708i
+9 0 9.304080.48927i9.30408-0.48927i 8.614130.45250i8.61413-0.45250i 7.992830.41976i7.99283-0.41976i 7.435550.39076i7.43555-0.39076i 6.973560.37212i6.97356-0.37212i
+1 9.252551.47571i9.25255-1.47571i 8.563641.36779i8.56364-1.36779i 7.880561.26327i7.88056-1.26327i 7.408401.17056i7.40840-1.17056i 6.881581.09045i6.88158-1.09045i
+2 9.150682.45648i9.15068-2.45648i 8.475192.27130i8.47519-2.27130i 7.842632.09637i7.84263-2.09637i 7.271801.96059i7.27180-1.96059i 6.887561.86712i6.88756-1.86712i
+Table 5: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Axial perturbation) with $M=0,0.025b_{0},0.05b_{0},0.075b_{0},0.1b_{0}$.
+(The results are obtained using Prony fit.) [The fundamental modes were computed in Blázquez-Salcedo et al.
+(2018a).
+We extended the analysis to include overtones.]
+It is interesting to note that both the real and imaginary parts of the $\omega_{QNM}$ decrease with increasing M, with the fundamental mode and the overtones following the same pattern.
+We can estimate the QNMs for a certain M value from the spectra of the original Bronnikov–Ellis wormhole ($M=0$ case) using the purely empirical formula
+$\omega_{M}=\omega_{0}exp\left(-\frac{\pi M}{1+0.8M}\right)$
+where, $\omega_{0}$ is quasi-normal mode corresponding to $M=0$ case and $\omega_{M}$ denotes quasi-normal modes corresponding to $M\neq 0$ cases.
+This relation is true for both scalar and axial perturbations of the $n=2$ geometry and holds for all the values of the overtone number and the angular momentum parameter (i.e.
+the ones we have obtained numerically).
+We are not aware of the origin of this empirical relation, though, for small values of the exponent, an approximate linear relation seems to work, as evident from the plots.
+In Figs. 39 (scalar) and 39 (axial), we have plotted the QNMs (obtained using the Prony Fit method) and the above empirical fitting function for the $n=2$ and $m=7$ case.
+Figure 38: Fit of the scalar $\omega_{QNM}$ ($m=7$) using the empirical fitting function.
+Figure 39: Fit of the axial $\omega_{QNM}$ ($m=7$) using the empirical fitting function.
+For the $n=2$ case, one can observe that the increasing ADM mass ($M$) has a very similar effect on fundamental mode and overtones.
+Due to this feature of the QNM spectra, the ringdown signal of a Bronnikov-Ellis wormhole with non-zero $M$ can be mistaken as that of the massless Bronnikov-Ellis wormhole with a larger throat radius ($b_{0}$).
+V.2 Bronnikov–Ellis geometry with a chosen redshift function:
+Apart from the geometry we studied in the previous subsection, we may also consider the Bronnikov–Ellis wormhole (the $n=2$ case) with a different non-zero redshift function, introduced and discussed earlier in Section II.3.
+The line element is given as,
+$ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\frac{dr^{2}}{1-\frac{b_{0}^{2}}{r^{2}}}+r^{2}d\theta^{2}+r^{2}sin^{2}\theta d\varphi^{2}$
+Here, as stated before, the energy density does not change on varying $M$, but the pressures change. We will now see if a redshift effect of the QNMs is present when we tune the value of $M$.
+The Schrodinger-like equation for perturbations of this spacetime is
+$\frac{d^{2}\psi(r)}{dr_{*}^{2}}+(\omega^{2}-V_{eff}(r))\psi(r)=0$ (21)
+where the tortoise coordinate $r_{*}$ has the form
+$r_{*}=$ $\frac{1}{\sqrt{r\left(r^{2}-b_{0}^{2}\right)(r-2M)}(4M^{2}-b_{0}^{2})}\bigg(b_{0}\bigg(b_{0}^{2}k_{1}(r-2M)^{2}\left(E\left(k_{0}|a_{1}\right)-F\left(k_{0}|a_{1}\right)\right)+2b_{0}M\big(k_{1}(r-2M)^{2}\big(E\left(k_{0}|a_{1}\right)$
+$-\Pi\left(a_{0};k_{0}|a_{1}\right)\big)-2Mr\big)-4k_{1}M^{2}(r-2M)^{2}\left(F\left(k_{0}|a_{1}\right)-\Pi\left(a_{0};k_{0}|a_{1}\right)\right)+b_{0}^{3}r\bigg)+r^{3}(4M^{2}-b_{0}^{2})\bigg)$
+with $a_{0}=\frac{2b_{0}}{b_{0}+2M}$, $a_{1}=\frac{4M}{b_{0}+2M}$, $k_{0}=sin^{-1}\sqrt{\frac{\left(b_{0}+2M\right)\left(r-b_{0}\right)}{2b_{0}(r-2M)}}$, $k_{1}=\sqrt{\frac{r\left(b_{0}-2M\right)\left(b_{0}^{2}-4M^{2}\right)\left(r^{2}-b_{0}^{2}\right)}{b_{0}^{3}(r-2M)^{3}}}$, E()=EllipticEE(...)=EllipticE, F()=EllipticFF(...)=EllipticF and $\Pi(...)=EllipticPi$ function.
+Now, to compute the QNMs with the spectral method, it is convenient to consider two separate ansatz Alfaro et al. (2024) (one for odd overtones and one for even).
+$\psi(r)_{even}=e^{i\omega r}r^{iM\omega}\phi(r)_{even}\hskip 56.9055pt\psi(r)_{odd}=(r-b_{0})^{\frac{1}{2}}e^{i\omega r}r^{iM\omega-\frac{1}{2}}\phi(r)_{odd}$
+Further, we have used the transformation $x=1-\frac{2b_{0}}{r}$ to map the radial coordinate $r\in(b_{0},\infty)$ to $x\in[-1,1]$ Batic and Dutykh (2024).
+This procedure is applicable for both scalar and axial perturbations.
+The tables below (6, 7) contain results obtained using the spectral method.
+However, we have also shown the results from the Prony fit method in the plots.
+For scalar perturbation the expression for $V_{eff}$ is
+$V_{sc}(r)=\frac{b_{0}^{2}(r-3M)+r^{2}\left(\left(-2m^{2}-2m+1\right)M+m(m+1)r\right)}{r^{5}}$
+As shown in the figures below (Figs. 41, 41), the effective potential changes from being a single barrier to a double barrier for higher $M$. The potentials have been plotted numerically.
+Figure 40: Effective potentials (Scalar Perturbation) for different values of $M$. (n = 2 and m = 5).
+Figure 41: Effective potentials (Scalar Perturbation) for different values of $M$. (n = 2 and m = 7).
+For this geometry, as we increase the value of $M$, both the real and imaginary parts of the QNM spectra decrease (See Figs.
+43 ,43).
+Thus, an increase in $M$ causes a redshift.
+On careful observation, one notices that for lower values of $M$, the real parts of quasinormal modes decrease with the overtone number.
+However, this pattern reverses when we go to higher $M$ values.
+The height of the effective potential decreases as we keep on increasing $M$, and after a certain value, the effective potential becomes double-peaked.
+This may explain the above feature of the quasi-normal spectra.
+For a fixed overtone number, the $\omega_{QNM}$ of scalar perturbation corresponding to different $M$ values can be roughly estimated by linear fits.
+The fitting parameter will vary with the overtone number and angular momentum parameter (m).
+Figure 42: QNMs (Scalar) for different values of $M$. (n = 2 and m = 5). The results of SM and Prony fit are represented by circles and plus signs, respectively.
+Figure 43: QNMs (Scalar) for different values of $M$. (n = 2 and m = 7). The results of SM and Prony fit are represented by circles and plus signs, respectively.
+$m$ Mode $M=0b_{0}$ $M=0.1b_{0}$ $M=0.2b_{0}$ $M=0.3b_{0}$ $M=0.4b_{0}$
+1 0 1.572710.52970i1.57271-0.52970i 1.424430.44987i1.42443-0.44987i 1.261840.35838i1.26184-0.35838i 1.079800.25189i1.07980-0.25189i 0.861520.12862i0.86152-0.12862i
+1 1.259571.70681i1.25957-1.70681i 1.185311.43536i1.18531-1.43536i 1.129611.14815i1.12961-1.14815i 1.081470.84604i1.08147-0.84604i 1.011040.51803i1.01104-0.51803i
+2 - - 0.973852.14007i0.97385-2.14007i 1.098391.63586i1.09839-1.63586i 1.181061.10171i1.18106-1.10171i
+3 0 3.534290.50692i3.53429-0.50692i 3.170400.42628i3.17040-0.42628i 2.762720.32941i2.76272-0.32941i 2.295670.20688i2.29567-0.20688i 1.738320.06152i1.73832-0.06152i
+1 3.390291.51877i3.39029-1.51877i 3.064061.29199i3.06406-1.29199i 2.715761.00573i2.71576-1.00573i 2.347540.67242i2.34754-0.67242i 1.934050.30467i1.93405-0.30467i
+2 3.100632.62534i3.10063-2.62534i 2.851462.20091i2.85146-2.20091i 2.616151.73263i2.61615-1.73263i 2.387301.23570i2.38730-1.23570i 2.089890.70557i2.08989-0.70557i
+3 2.718633.55029i2.71863-3.55029i - 2.463052.54081i2.46305-2.54081i 2.392441.88891i2.39244-1.88891i 2.234651.18172i2.23465-1.18172i
+5 0 5.522310.50296i5.52231-0.50296i 4.945510.42182i4.94551-0.42182i 4.295170.32260i4.29517-0.32260i 3.541420.19058i3.54142-0.19058i 2.636860.03204i2.63686-0.03204i
+1 5.413721.50367i5.41372-1.50367i 4.878851.27072i4.87885-1.27072i 4.270800.97686i4.27080-0.97686i 3.604200.61373i3.60420-0.61373i 2.858990.20555i2.85899-0.20555i
+2 5.250112.54570i5.25011-2.54570i 4.745242.13600i4.74524-2.13600i 4.217291.65629i4.21729-1.65629i 3.670731.11220i3.67073-1.11220i 3.028660.52636i3.02866-0.52636i
+3 4.970143.61705i4.97014-3.61705i 4.544923.03031i4.54492-3.03031i 4.129082.37426i4.12908-2.37426i 3.713541.67104i3.71354-1.67104i 3.191440.91448i3.19144-0.91448i
+7 0 7.516490.50162i7.51649-0.50162i 6.727600.42027i6.72760-0.42027i 5.835750.31996i5.83575-0.31996i 4.795240.18192i4.79524-0.18192i 3.538880.01635i3.53888-0.01635i
+1 7.452251.50671i7.45225-1.50671i 6.679101.26365i6.67910-1.26365i 5.820190.96550i5.82019-0.96550i 4.859280.58145i4.85928-0.58145i 3.780570.14253i3.78057-0.14253i
+2 7.315422.52531i7.31542-2.52531i 6.581912.11572i6.58191-2.11572i 5.785901.62677i5.78590-1.62677i 4.935871.04502i4.93587-1.04502i 3.958450.40847i3.95845-0.40847i
+3 7.113303.56071i7.11330-3.56071i 6.435132.98262i6.43513-2.98262i 5.728572.31139i5.72857-2.31139i 4.998351.55865i4.99835-1.55865i 4.134490.74607i4.13449-0.74607i
+Table 6: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Scalar perturbation) with $M=0,0.1b_{0}$, $0.2b_{0}$, $0.3b_{0}$ and $0.4b_{0}$.
+[The QNM spectra for this geometry (single-peaked scalar potentials) were published in Alfaro et al.
+(2024).
+We extended this to double-peaked potentials (Higher $M$)].
+[Results obtained using the spectral method ($b_{0}=1$).]
+In the axial perturbation case, the effective potential in Eq. (21) has the form
+$V_{ax}(r)=\frac{b_{0}^{2}(7M-3r)+r^{2}(m(m+1)r-(2m(m+1)+1)M)}{r^{5}}$
+Similar to the scalar perturbation, the effective potential due to axial perturbation changes from being a single barrier to a double barrier for higher $M$ (See Figs. 45 ,45).
+The quasi-normal spectra for axial perturbation have a similar redshift effect as the QNMs for scalar perturbation (Figs.
+47 ,47).
+However, unlike the scalar perturbation case, here, a linear fit does not estimate the $\omega_{QNM}$ of the same overtone order.
+Figure 44: Effective potentials (Axial Perturbation) for different values of $M$. (n = 2 and m = 5).
+Figure 45: Effective potentials (Axial Perturbation) for different values of $M$. (n = 2 and m = 7).
+Figure 46: QNMs (Axial) for different values of $M$. (n = 2 and m = 5). The results of SM and Prony fit are represented by circles and plus signs, respectively.
+Figure 47: QNMs (Axial) for different values of $M$. (n = 2 and m = 7). The results of SM and Prony fit are represented by circles and plus signs, respectively.
+$m$ Mode $M=0b_{0}$ $M=0.1b_{0}$ $M=0.2b_{0}$ $M=0.3b_{0}$ $M=0.4b_{0}$
+3 0 2.952410.40998i2.95241-0.40998i 2.652020.34116i2.65202-0.34116i 2.317040.25890i2.31704-0.25890i 1.935830.15765i1.93583-0.15765i 1.478950.04526i1.47895-0.04526i
+1 2.866431.25767i2.86643-1.25767i 2.599431.05329i2.59943-1.05329i 2.320190.81701i2.32019-0.81701i 2.023200.54789i2.02320-0.54789i 1.681430.25349i1.68143-0.25349i
+2 2.660562.19085i2.66056-2.19085i 2.472321.84041i2.47232-1.84041i 2.283661.46458i2.28366-1.46458i 2.092621.06263i2.09262-1.06263i 1.837720.62125i1.83772-0.62125i
+5 0 5.154820.46725i5.15482-0.46725i 4.617430.39009i4.61743-0.39009i 4.012260.29556i4.01226-0.29556i 3.313170.17059i3.31317-0.17059i 2.474820.02713i2.47482-0.02713i
+1 5.072861.40857i5.07286-1.40857i 4.561441.17693i4.56144-1.17693i 4.000380.89951i4.00038-0.89951i 3.387860.56119i3.38786-0.56119i 2.699340.18747i2.69934-0.18747i
+2 4.907462.37118i4.90746-2.37118i 4.447731.98388i4.44773-1.98388i 3.966461.53674i3.96646-1.53674i 3.465291.03582i3.46529-1.03582i 2.867940.49395i2.86794-0.49395i
+3 4.656613.37224i4.65661-3.37224i 4.273002.82560i4.27300-2.82560i 3.899882.21918i3.89988-2.21918i 3.517281.57299i3.51728-1.57299i 3.030590.87000i3.03059-0.87000i
+7 0 7.248180.48308i7.24818-0.48308i 6.487870.40372i6.48787-0.40372i 5.628680.30567i5.62868-0.30567i 4.627880.17089i4.62788-0.17089i 3.420550.01444i3.42055-0.01444i
+1 7.184381.45261i7.18438-1.45261i 6.443081.21430i6.44308-1.21430i 5.618090.92386i5.61809-0.92386i 4.698040.55209i4.69804-0.55209i 3.664050.13378i3.66405-0.13378i
+2 7.056342.43256i7.05634-2.43256i 6.353142.03436i6.35314-2.03436i 5.592221.56056i5.59222-1.56056i 4.780411.00184i4.78041-1.00184i 3.841390.39161i3.84139-0.39161i
+3 6.865883.43084i6.86588-3.43084i 6.217332.87112i6.21733-2.87112i 5.544942.22360i5.54494-2.22360i 4.847721.50282i4.84772-1.50282i 4.016930.72270i4.01693-0.72270i
+9 0 9.301710.48964i9.30171-0.48964i 8.323540.40939i8.32354-0.40939i 7.216380.30982i7.21638-0.30982i 5.920450.16947i5.92045-0.16947i 4.347390.00712i4.34739-0.00712i
+1 9.250351.47100i9.25035-1.47100i 8.287121.23004i8.28712-1.23004i 7.207680.93382i7.20768-0.93382i 5.986670.54158i5.98667-0.54158i 4.608390.09362i4.60839-0.09362i
+2 9.147412.45870i9.14741-2.45870i 8.214142.05635i8.21414-2.05635i 7.187611.56991i7.18761-1.56991i 6.070490.97279i6.07049-0.97279i 4.792710.31009i4.79271-0.31009i
+3 8.992393.45726i8.99239-3.45726i 8.104242.89222i8.10424-2.89222i 7.152272.22372i7.15227-2.22372i 6.146251.44932i6.14625-1.44932i 4.975510.60743i4.97551-0.60743i
+Table 7: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Axial perturbation) with $M=0,0.1b_{0}$, $0.2b_{0}$, $0.3b_{0}$ and $0.4b_{0}$.
+[The QNM spectra for this geometry (single-peaked scalar potentials) were published in Alfaro et al.
+(2024).
+We extended this to axial perturbation and double-peaked potentials (Higher $M$)].
+[Results obtained using the spectral method ($b_{0}=1$).]
+[…]
+VI Remarks and future directions
+In our work here, we have revisited the problem of computing the QNMs of the generalised Bronnikov-Ellis wormholes with emphasis on the overtones.
+Further, we have also studied the ‘redshift’ of QNMs induced by varying the metric parameters in analogy with the environment-induced proposed recently.
+Below, we briefly summarise our main results.
+[…]
+$\bullet$ We have used the spectral method (frequency domain) and the Prony Fit (time domain) to compute the QNM spectra (scalar and axial gravitational) for the ultrastatic case.
+Beyond the fundamental modes, we have computed 2-3 overtones (4 in some cases), which, as we show, may help in distinguishing different geometries (i.e.
+with different $n$).
+We have also pointed out key features of the QNM spectra and how the shape of the effective potentials may explain the trends in QNM behaviour.
+$\bullet$ The range of the throat radius ($b_{0}$) for detectable QNMs (in aLIGO) in the ultrastatic case is found.
+We also establish, for a few cases, using the time-domain waveform, the crucial role that overtones may play in fitting the signal at early times (when SNR is high).
+[…]
+Various extensions of this work here are possible.
+We have not really succeeded in obtaining a large number of overtones which can give us a clearer picture of the full quasinormal mode spectrum.
+This aspect surely needs to be handled better.
+Further, though we have identified distinct features and have tried to demonstrate how overtones may affect a signal, a better and more complete understanding is desirable.
+In addition, learning more about polar perturbations and their overtones remains a difficult task which needs to be addressed.
+Finally, in the context of the redshift, though we see a correlation with the behaviour of the effective potentials, the results are case-specific and need to be understood in a more generic way.
+We hope to return to the abovementioned issues in future investigations.
+ACKNOWLEDGEMENTS
+ABM thanks the Indian Institute of Technology Kharagpur, India, for supporting him through a fellowship and allowing access to available computational facilities.
+Appendix A Methods used to find the QNMs
 Computation of quasinormal modes requires solving linearised perturbation equations of spacetime under specific boundary conditions (purely outgoing at spatial infinities of the tortoise coordinate).
@@
 Below we provide short summaries of the key methods we have used here to determine the QNMs of the generalised Bronnikov–Ellis wormholes.
-III.1 Time domain and Prony fit
+A.1 Time domain and Prony fit
 The time domain profile represents the temporal evolution of the field $\psi$ in a fixed background geometry.
@@
 (1994).
-In the previous section, we derived the master wave equation, which may also be rewritten without the separability ansatz for the field.
+In section II.4, we derived the master wave equation, which may also be rewritten without the separability ansatz for the field.
 Thus, the equation, using the new coordinates $u$ and $v$, such that $du=dt-dx,\hskip 5.69046ptdv=dt+dx$, reads Konoplya and Zhidenko (2011):
-$\frac{\partial^{2}\psi}{\partial u\partial v}+V(u,v)\psi=0.$ (20)
+$4\frac{\partial^{2}\psi}{\partial u\partial v}+V(u,v)\psi=0.$ (22)
 Rewriting the time evolution operator in the new coordinates and discretising the wave equation, we obtain,
@@
 Eqn.
-(21) enables us to compute $\psi$ iteratively on the $u-v$ grid based on the initial data specified on null surfaces $u=u_{0}$ and $v=v_{0}$.
+(23) enables us to compute $\psi$ iteratively on the uvu-v grid based on the initial data specified on null surfaces $u=u_{0}$ and $v=v_{0}$.
 The values of $\psi(N)$, thus calculated, give us the time-domain profile for the given potential.
-The time-domain profile of $\psi$ is then fitted using the Prony method $\psi(t)\approx\sum_{j=1}^{p}C_{j}e^{i\omega_{j}t}$.
+The time-domain profile of $\psi$ is then fitted using the Prony method $\psi(t)\approx\sum_{j=1}^{p}C_{j}e^{-i\omega_{j}t}$.
 Here $\omega_{j}$ are the quasinormal frequencies and $C_{j}$ are the corresponding amplitudes.
 We assume that the ringdown starts at $t=0$ and ends at $t=N_{1}h$, and approximate the profile data as a sum of $p\hskip 2.84544pt(\leq\frac{N_{1}+1}{2})$ terms in the following way:
-$x_{n}\approx\psi(nh)=\sum_{j=1}^{p}C_{j}z_{j}^{n},\hskip 56.9055ptz_{j}=e^{-i\omega_{j}nh}$ (22)
+$x_{n}\approx\psi(nh)=\sum_{j=1}^{p}C_{j}z_{j}^{n},\hskip 56.9055ptz_{j}=e^{-i\omega_{j}h}$ (24)
 In the Prony fit method, it is assumed that the $\psi(nh)$ satisfies the recurrence relation
@@
 where $a_{0}=1$. This leads to the matrix equation:
-$\begin{bmatrix}x_{p-1}&x_{p-2}&\ldots&x_{0}\\ x_{p}&x_{p-1}&\ldots&x_{1}\\ \vdots&\vdots&\ddots&\vdots\\ x_{N-1}&x_{N-2}&\ldots&x_{N-p}\end{bmatrix}\begin{bmatrix}a_{1}\\ a_{2}\\ \vdots\\ a_{p}\end{bmatrix}=-\begin{bmatrix}x_{p}\\ x_{p+1}\\ \vdots\\ x_{N}\end{bmatrix}\hskip 28.45274pt\implies X\cdot a=-x,$ (24)
+$\begin{bmatrix}x_{p-1}&x_{p-2}&\ldots&x_{0}\\ x_{p}&x_{p-1}&\ldots&x_{1}\\ \vdots&\vdots&\ddots&\vdots\\ x_{N_{1}-1}&x_{N_{1}-2}&\ldots&x_{N_{1}-p}\end{bmatrix}\begin{bmatrix}a_{1}\\ a_{2}\\ \vdots\\ a_{p}\end{bmatrix}=-\begin{bmatrix}x_{p}\\ x_{p+1}\\ \vdots\\ x_{N_{1}}\end{bmatrix}\hskip 28.45274pt\implies X\cdot a=-x,$ (26)
 Eqn.
@@
 Having computed the quasi-normal frequencies, we can also obtain their corresponding amplitudes using the Prony fit.
-For this, we need to define a $(p\times N)$ matrix $A_{ij}=z_{i-1}^{j}$ and a vector $v=[x_{1},x_{2},\ldots,x_{N}]$.
+For this, we need to define a $(p\times N_{1})$ matrix $A_{ij}=z_{i-1}^{j}$ and a vector $v=[x_{1},x_{2},\ldots,x_{N_{1}}]$.
 We get the amplitudes by calculating $C=[C_{1},C_{2},\ldots,C_{p}]=(A^{\dagger}A)^{-1}A^{\dagger}v$.
+These amplitudes are generally complex numbers; however, we have reported only the absolute values in the appendix B.
 In the method discussed above, we get $p$-number of solutions, but most of them are numerical artefacts and not actual QNMs of the system.
-To mitigate this issue, we can repeat the above steps with different values of $h$ and $N$ and accept only the results common to both sets of solutions (with appropriate tolerance) as the actual QNMs.
+To mitigate this issue, we can repeat the above steps with different values of $h$ and $N_{1}$ and accept only the results common to both sets of solutions (with appropriate tolerance) as the actual QNMs.
 The time-domain and Prony fit method works well for most effective potentials which do not have any discontinuities.
 But this method often fails to return the $\omega_{QNM}$ with low relative amplitude (we notice this feature for odd overtones in Sec.
-IV).
-III.2 The spectral method
+III).
+A.2 The spectral method
 Using the corresponding equation for a perturbation (scalar or axial gravitational), another method known as the spectral method, may also be employed to find the QNMs of the system.
 In our analysis, we will use the Chebyshev grid and thus we have to map our coordinate to a finite interval, i.e.
-$[-1,1]$.
-The tortoise coordinate $l\in(-\infty,+\infty)$ may be mapped to $x\in[-1,1]$ using the transformation $x=\frac{2}{\pi}arctan(l)$ Batic and Dutykh (2024).
+[1,1][-1,1].
+The tortoise coordinate $l\in(-\infty,+\infty)$ may be mapped to $x\in[-1,1]$ using the transformation $x=\frac{2}{\pi}\arctan(l)$ Batic and Dutykh (2024).
 For the generalised Bronnikov–Ellis wormhole with $M=0$, we write the Schrodinger-like equation in terms of the tortoise coordinate ($l$),
@@
 For that, we examine the behaviour of $\psi$ near the throat and at infinity and define a new variable $\phi(x)$ using a suitable ansatz: Batic and Dutykh (2024).
-$\psi(x)=exp(i\omega\frac{2l}{\pi}arctan(l))\phi(x)$
+$\psi(x)=exp\left(i\omega\frac{2l}{\pi}\arctan(l)\right)\phi(x)$ (29)
 Next, the function $\phi$ is discretized with the Chebyshev polynomial of first kind $C_{i}(x)$ as Jansen (2017)
 $\phi(x)\approx\sum_{i=0}^{N}\phi(x_{i})C_{i}(x)$ (30)
-with $x_{i}$ being the collocation points (Chebyshev roots) given by $x_{i}=\frac{(2i+1)\pi}{2(N+1)}$ (i = 0,1,2,…,N).
+with $x_{i}$ being the collocation points (Chebyshev roots) given by $x_{i}=\frac{(2i+1)\pi}{2(N+1)}$ (i=0,1,2,,N)(i={0,1,2,...,N}).
 Thus, we can express the master equation in the form
@@
 We compute these expressions and $\phi$ at the co-location points and get numerical matrices of the form
-[…]
+$(M_{a})_{ij}=c_{0,a}D^{(0)}_{ij}+c_{1,a}D^{(1)}_{ij}+c_{2,a}D^{(2)}_{ij}\hskip 28.45274pta={1,2,...}$ (32)
 where, $D^{(n)}_{ij}$ represents n-th derivative of $C_{i}(x_{j})$. If the highest power of $\omega$ in Eqn. (31) is $p$, then we will need to compute $p$ such matrices. Finally, (31) is expressed as
@@
 Solving the matrix equation (33) will return the quasi-normal frequencies of the system.
-III.3 Other methods: merits and demerits
-Apart from the methods discussed above, there are a plethora of analytic and numerical methods that can be employed to compute the quasinormal spectra.
-One can compare numerically obtained QNMs with exact results for the few effective potentials which are analytically solvableFerrari and Mashhoon (1984); Heidari and Hassanabadi (2023).
-In some cases, the exactly solvable potentials are used as approximations for the real potentials in black hole perturbations (for example, in the Schwarzschild, a matching may be done using the Poschl–Teller potential, as shown many years ago in Ferrari and Mashhoon (1984) ).
-The QNMs found following this route however do not match accurately when compared with results using other numerical methods.
-In particular, the $Re(\omega)$ for the overtones show a mismatch.
+We can consider the example of scalar perturbation in the ultra-static Bronnikov–Ellis geometry (Eqn. (5)), where, after using the ansatz (Eqn. (29)) we have
+$c_{0}=$ $-m(m+1)\left(\tan^{n}\left(\frac{\pi x}{2}\right)+1\right)^{-2/n}-\frac{(n-1)\tan^{n-2}\left(\frac{\pi x}{2}\right)}{\left(\tan^{n}\left(\frac{\pi x}{2}\right)+1\right)^{2}}-\left(x^{2}-1\right)\omega^{2}$
+$+\frac{2i\pi\omega\cos(\pi x)-\omega\sin(\pi x)(2\pi x\omega+(\omega+i\pi)\sin(\pi x))}{\pi^{2}}+\frac{2i\omega}{\pi}$
+$c_{1}=\frac{2i\cos^{2}\left(\frac{\pi x}{2}\right)(2\pi x\omega+(2\omega+i\pi)\sin(\pi x))}{\pi^{2}}\hskip 85.35826ptc_{2}=\frac{4\cos^{4}\left(\frac{\pi x}{2}\right)}{\pi^{2}}$
+Now we can take the coefficient of $\omega^{0}$ in $c_{0}$ to get $c_{0,0}$, the coefficient of $\omega^{1}$ in $c_{0}$ to get $c_{1,0}$ and so on.
+A.3 Other methods: merits and demerits
+[…]
 The direct integration method is another method that we have used to verify the fundamental modes for our geometries.
@@
 In the results presented below, we have primarily used the Prony fit (time domain) and the spectral method (SM) while finding the QNMs.
-IV Results: quasinormal spectra (fundamental and overtones for the ultra-static case)
-The problem of computing quasi-normal modes for the Bronnikov–Ellis wormhole has been tackled in many previous works.
-The authors of Kim (2008); Blázquez-Salcedo et al.
-(2018b) and Dutta Roy et al.
-(2020); Dutta Roy (2022) have presented the fundamental modes for the original Bronnikov–Ellis ($n=2$ case) and the generalised Bronnikov–Ellis wormholes respectively.
-However, as pointed out in Sago et al.
-(2021); Berti et al.
-(2007); Giesler et al.
-(2019); Nee et al.
-(2023), the fundamental modes may not represent the full picture, as the overtones play a crucial role in fitting the ringdown (RD) waveform.
-The inclusion of overtones (at least, the first few overtones) can help us achieve a better fit of the RD profile at earlier times when we expect a better signal-to-noise ratio (SNR).
-Overtones can also be instrumental in providing better estimates of the parameters of the underlying geometry.
-Though the overtones corresponding to the $n=2$ geometry were computed in Alfaro et al.
-(2024); Batic and Dutykh (2024), we compute the full QNM spectra for the generalised Bronnikov–Ellis spacetimes (even $n>2$ cases).
-We note many key differences between the $n=2$ and the $n>2$ cases.
-The authors of Baibhav et al.
-(2023) have argued that the overtones having lower real parts compared to the fundamental mode (true for $n=2$), fit away the non-linearities of the RD signal.
-The same argument will not be true for $n>2$ geometries (where $Re(\omega)$ for overtones are higher in value than the fundamental mode).
-Any possible future detection of such RD signals may further cement the role of overtones.
-As stated above, while computing the $\omega_{QNM}$, we have used a frequency-domain method (spectral method (SM)) in conjunction with a time-domain method (Prony fit).
-While the frequency-domain method is indispensable for extracting all the modes (up to a certain overtone number), the time-domain provides us with some valuable insight into the relative amplitudes of the overtones, thus pointing to the fact that sometimes even higher-order overtones (in our case: even modes) can have greater importance than the lower-order (in our case: odd modes) ones.
-The results from the two methods are in good agreement with each other.
-However, we could not compute the same number of overtones for every geometry.
-For lower values of $n$ and angular momentum parameter $m$, we could get very few overtones.
-In the series of figures below, we demonstrate the effective potentials and QNMs for scalar and axial gravitational perturbations.
-Figs.
-2-5 are for $n=2$, Figs.
-$6-9$ for $n=4$, Figs.
-10-13 are for $n=6$ and Figs.
-14-17 for $n=8$.
-In the following section, we discuss the results here and try to arrive at specific conclusions on the nature of the QNMs and their link with the different geometries.
-QNMs for n = 2 geometry:
-Figure 3: Effective potential (Scalar) for Bronnikov–Ellis wormhole with $n=2$.
-Figure 4: Effective potential (Axial) for Bronnikov–Ellis wormhole with $n=2$.
-Figure 5: QNMs for n = 2 (Scalar) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 8)
-Figure 6: QNMs for n = 2 (Axial) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 8)
-QNMs for n = 4 geometry:
-Figure 7: Effective potential (Scalar) for generalised Bronnikov–Ellis wormhole with $n=4$.
-Figure 8: Effective potential (Axial) for generalised Bronnikov–Ellis wormhole with $n=4$.
-Figure 9: QNMs for n = 4 (Scalar) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 9)
-Figure 10: QNMs for n = 4 (Axial) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 9)
-QNMs for n = 6 geometry:
-Figure 11: Effective potential (Scalar) for generalised Bronnikov–Ellis wormhole with $n=6$.
-Figure 12: Effective potential (Axial) for generalised Bronnikov–Ellis wormhole with $n=6$.
-Figure 13: QNMs for n = 6 (Scalar) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 10)
-Figure 14: QNMs for n = 6 (Axial) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 10)
-QNMs for n = 8 geometry:
-Figure 15: Effective potential (Scalar) for generalised Bronnikov–Ellis wormhole with $n=8$.
-Figure 16: Effective potential (Axial) for generalised Bronnikov–Ellis wormhole with $n=8$.
-Figure 17: QNMs for n = 8 (Scalar) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 11)
-Figure 18: QNMs for n = 8 (Axial) case. The results of SM and Prony fit are represented by circles and plus signs, respectively. (Table: 11)
-V Inferences drawn from the QNM spectra: emphasis on overtones
-In Section IV, we presented the fundamental modes and overtones of the generalised, ultrastatic, Bronnikov–Ellis wormhole.
-Now, one may ask whether these QNMs can help distinguish different geometries (i.e.
-different $n$ values) of this wormhole family.
-To answer this query, we focus on the plots ($Re(\omega)$ vs $-Im(\omega)$) given above in the previous section and try to identify some general trends in the QNM spectra.
-In our analysis, we denote the mode numbers with $p$, where the mode with the smallest value of $|Im(\omega)|$ (the longest-lived mode) is the fundamental mode (see e.g.
-Giesler et al.
-(2019)) with $p=0$ and modes with higher $|Im(\omega)|$ are the overtones.
-For both scalar and axial perturbations, in the $n=2$ case, the real parts of the $\omega_{QNM}$ decrease with increase in mode number ($p$), while for the geometries with higher $n$ values, the higher overtones correspond to an increase in $Re(\omega)$.
-From the tables in Appendix A, we notice that for a fixed large value of the angular momentum parameter ($m$), the fundamental modes of higher $n$ geometries are very close in value.
-However, these wormhole spacetimes can also be distinguished from each other if we take the overtones into account.
-As we move to larger $n(>2)$ values, we observe that the rate of increase of $Re(\omega)$ with the mode number gets expedited.
-This is true for both scalar and axial perturbations (see Figs.
-6, 6, 10, 10…, 18).
-The imaginary parts of the quasinormal spectra are also an important feature distinguishing the geometries.
-In the scalar perturbation case, for a fixed angular momentum parameter value, the greater the $n$ value, the smaller the magnitude of the imaginary parts of the QNMs.
-The separation between the $|Im(\omega)|$ also keeps decreasing (See Figs.
-6, 10, 14, 18).
-This implies that for wormholes with larger values of $n$, the fundamental modes and overtones have a lower decay rate and, therefore, are less stable.
-As we keep on increasing $m$, for the axial perturbation case, the absolute value of the imaginary parts of the fundamental modes first increase and then decrease (See Figs.
-6, 10, 14, 18).
-Having identified the broader trends from the plots, we can try to explain them from the effective potential.
-The decrease of $Re(\omega)$ with overtone mode number is only observed in the $n=2$ case, which has a comparatively narrower and sharper single-peaked potential (See Figs.
-4, 4).
-This trend is reversed for all even $n>2$ wormholes, which have a comparatively broader and flatter effective potential with two peaks for scalar perturbation (See Figs.
-8, 12, 16) and one (for higher values of m)/three (for smaller values of m) peak(s) for axial perturbations (See Figs.
-8, 12, 16).
-On careful observation, one can notice that as we go to higher values of $n$, the central region of the potential becomes flatter.
-This can be a possible explanation behind the enhanced rate of increase of the real parts for higher $n$ geometries.
-Fixing the value of $n$, if we increase the angular momentum parameter ($m$), the height of the effective potential increases, which is related to increasing the real part of the QNMs.
-The double peak feature also becomes less prominent for the scalar perturbation case (See Figs.
-8, 12, 16).
-This may explain the decrease of $|Im(\omega)|$ with $m$ in the quasi-normal spectra.
-For axial perturbations in $n>2$ geometries, the transition of the potential from triple barrier to step-like and then to a single barrier, with increasing m (See Figs.
-8, 12, 16), may cause the initial increase and subsequent decrease of $|Im(\omega)|$ for the fundamental modess with m.
-Variation Scalar Axial
-The rate of increase The rate of increase
-Increase n: of $Re(\omega)$ with the mode of $Re(\omega)$ with the mode
-(for $n>2$) number gets expedited. number gets expedited.
-$Re(\omega)$ increases and $Re(\omega)$ increases and
-Increase m: $|Im(\omega)|$ decreases $|Im(\omega)|$ (fundamental mode) first
-increases and then decreases.
-The rate of increase of The rate of increase of
-Increase p: $Re(\omega)$ does not change $Re(\omega)$ increases after a
-(for $n>2$) with overtone number. certain overtone number.
-$|Im(\omega)|$ increases. $|Im(\omega)|$ increases.
-Table 1: Key features of the QNM spectra for massless generalised Bronnikov-Ellis wormhole.
-V.1 Approximate Analytic Fit
-To further advance our study of the QNM spectra and their dependence on different parameters such as the geometry ($n$), angular momentum parameter ($m$), overtone order ($p$) and the throat ($b_{0}$), we can fit our numerical results (for scalar perturbations) to analytical models.
-Here, the throat radius is simply a scaling parameter, and the quasi-normal frequencies are inversely proportional to $b_{0}$.
-However, the other parameters do not follow such simple rules.
-To get the fits corresponding to these parameters, we have used the NonLinearModelFit function of Mathematica 14.2.
-For all the fits obtained in this section, we have used the results from the spectral method.
-As mentioned above, the change in the QNMs with changing $n$ values may give us some valuable insights.
-To this end, we have analysed the dependence of $Re(\omega)$ on overtone mode number ($p$) and the angular momentum parameter ($m$), keeping the geometry parameter $n$ fixed.
-Figure 19: QNMs for n = 2 (Scalar Perturbation) case.
-Here, we observe that the real parts of the quasi-normal frequencies decrease with overtone mode $p$ but increase with angular momentum $m$.
-Figure 20: [Left] QNMs for n = 4; [Right] QNMs for n = 8.
-For geometries with higher $n$ $(>2)$, the real parts increase with $m$ as well as with overtone mode number $(p)$.
-We notice, with increasing $p$ $Re(\omega)$ increases faster for $n=8$ case as compared to say, the $n=4$ case.
-We have also fitted the imaginary part with the geometry parameter and the real part, for different overtone numbers ($p$), for scalar perturbations.
-Figure 21: Fit of Im($\omega$) with n and Re($\omega$): [left] fundamental modes; [middle] 1st overtones; [right] 2nd overtones.
-Thus far, we have (i) fitted $Re(\omega_{QNM})$ with angular momentum and overtone mode number and (ii) the $-Im(\omega_{QNM})$ with geometry parameter $n$ and $Re(\omega_{QNM})$.
-Hence, for given values of $n$, $m$ and $p$, we can numerically estimate the value of $\omega_{QNM}$ (scalar perturbation).
-The fitting functions discussed so far are approximate.
-They do not return exact results but provide information about overall trends, which are useful while interpreting our results.
-V.2 Detectable Modes
-Since we have now obtained the quasinormal modes including some overtones, it is important to ask whether or not these modes are detectable.
-If they are within observable limits, this must be for certain choices of the parameters involved.
-To figure this out, we will need to determine the numerical values for the quasinormal frequencies ($\nu$) and decay times ($\tau$), which can be found from
-[…]
-Advanced LIGO is sensitive to the frequency range of $10\hskip 5.69046ptHz$ to $10\hskip 5.69046ptkHz$.
-For the generalised Bronnikov-Ellis geometries, we have calculated the QNFs, and all of them can be within the aLIGO frequency range when the throat radius is roughly between $50M_{\odot}$ and $1000M_{\odot}$.
-For now, we choose $b_{0}=200M_{\odot}$ and present the frequencies and damping times for a few cases, as an illustration.
-In Table II below we show the values for $n=2$ and $m=5$ (scalar perturbation).
-Mode Frequency (Hz) Damping Time (s) Amplitude
-0 889.5 0.01235 0.2396
-1 872.6 0.00408 2.01×$\times 10^{-5}$
-2 845.8 0.00244 0.1219
-Table 2: Frequency and Damping time for $n=2$ and $m=5$. These values are calculated from the Prony fit (details in Appendix A).
-Figure 22: Log plot of time-domain profile for $\psi$ for $n=2$ and $m=5$.
-Figure 23: Time-domain waveform for $n=2$ and $m=5$.
-For $n=8$ and $m=5$ case (scalar):
-Mode Frequency (Hz) Damping Time (s) Amplitude
-0 904.2 0.06467 0.1550
-1 967.6 0.01617 1.789×$\times 10^{-5}$
-2 1067.7 0.00733 0.1368
-3 1200.3 0.00444 0.0006
-4 1359.8 0.00296 0.1100
-Table 3: Frequency and Damping time for $n=8$ and $m=5$. These values are calculated from the prony fit (details mentioned in Appendix A).
-Figure 24: Log plot time-domain profile for $\psi$ for $n=8$ and $m=5$.
-Figure 25: Time-domain waveform for $n=8$ and $m=5$.
-In the ringdown waveforms, we have represented the actual time-domain (obtained using the procedures mentioned in III.1) wave as the blue points, the fundamental mode as green lines and the resultant wave of the fundamental and overtones ($\omega$ and amplitudes computed using the Prony fit) as the magenta lines (Figs.
-23,25).
-For both cases, the magenta lines better fit the time domain than only the fundamental mode.
-However, the difference between the green and magenta lines is only noticeable for the first few cycles, as the overtones have higher imaginary parts and decay faster.
-While the inclusion of overtones makes minimal improvement in the fitting for the $n=2$ case, the effect is more significant for the $n=8$ case.
-This is because for the $n=8$ case, we could find more overtones, and these overtones have amplitudes comparable to the fundamental mode, whereas, in the $n=2$ case, the amplitudes of the overtones are significantly lower than the fundamental mode.
-For $n=4,6$ cases, we computed $2-3$ overtones and their corresponding amplitudes using the Prony fit, and similar to the $n=8$ case, we observe that inclusion of the overtones significantly improves the fit at early times.
-For all the geometries (i.e.
-different $n$), the more overtones we could get (up to $p=4$), the better we were able to model the time-domain signal, indicating that obtaining more overtones may result in a better fit of the time-domain profile.
-However, we need to be cautious not to overfit our results by incorporating an excessive number of overtones.
-One would expect the amplitudes of the modes to depend on our initial value of $\psi$ but not on which part of the time domain we are fitting.
-However, for the generalised Bronnikov–Ellis wormhole, only the amplitudes associated with the fundamental modes and 2nd overtones (to some extent) are time-independent.
-For almost all the other $\omega_{QNM}$, if we change the initial time of the fitting, the corresponding amplitudes remain stable up to some fitting domain, but beyond it, they begin to vary.
-Figure 26: Amplitude calculated for $n=2$ and $m=5$ with different initial time.
-Figure 27: Amplitude calculated for $n=8$ and $m=5$ with different initial time.
-Thus, we again fit the ringdown profiles with the modes with a stable amplitude
-Figure 28: Time-domain profile for all computed overtones and the modes with stable amplitude for $n=2$ and $m=5$.
-Figure 29: Time-domain profile for all computed overtones and the modes with stable amplitude for $n=8$ and $m=5$.
-In both plots (Figs.
-29, 29), the black line represents the contribution of the fundamental and 2nd overtone (the modes with stable amplitude), and the magenta line has the contribution of all the modes from tables 2 and 3.
-For the $n=2$ case, the black and the magenta lines overlap (this is expected as we filtered out the 1st overtone with a very small amplitude).
-In the $n=8$ case, the two lines deviate at a very early time.
-The inclusion of all the modes gives a slightly better fit as the modes that were filtered out are also stable in the initial times, and their inclusion does not change the late-time behaviour, as these modes are damped quickly.
-Thus, we conclude that the inclusion of overtones can indeed help us achieve a better fit of the time-domain profile at earlier times.
-VI Redshift
-In the previous sections, we have computed the quasinormal modes for the massless generalised Bronnikov–Ellis wormhole.
-However, if the geometry under study is surrounded by a dark matter halo or any generic matter distribution, the QNMs calculated for the isolated compact object may differ from the observed values, as the environment may affect the generation of GW as well as their propagation.
-Such environments cause the spectrum to be red-shifted as documented in Barausse et al.
-(2014); Pezzella et al.
-(2025); Liu et al.
-(2025) for the black hole spacetime embedded in a halo.
-In the case of the Bronnikov–Ellis wormhole, similar redshifts in the fundamental mode can be found, if one considers the wormhole with non-zero ADM mass, as noted by the authors of Blázquez-Salcedo et al.
-(2018b).
-We will first find the QNMs (fundamental and overtones) for the $M\neq 0$ case in the $n=2$ Bronnikov–Ellis geometry.
-Thereafter, in our second example, we will find the QNMs for a modified $n=2$ geometry with a new redshift function having an extra parameter $M$ Alfaro et al.
-(2024).
-In these two scenarios, using the QNMs obtained, we then verify if we have a parameter-controlled redshift, similar to the environment-induced redshift mentioned above.
-The modelling of such red-shifts is crucial as the red-shifted modes may appear to be QNMs of the original isolated geometry with a larger throat radius (as $\omega$ varies inversely with $b_{0}$).
-Let us first fix our general idea using some well-studied potentials as toy examples.
-We will try to find out how the $\omega_{QNM}$ vary if we change the width or other parameters present in two well-known potentials which have been investigated and used extensively in the past.
-VI.1 Rectangular Barrier:
-The rectangular barrier has the form:
-$V(x)=\begin{cases}V_{0}&\text{for }0\leq x\leq a,\\ 0&\text{otherwise}.\end{cases}$ (30)
-The exact transmission amplitude is $a_{t}=4\omega ke^{-i\omega a}((\omega+k)^{2}e^{-ika}-(\omega-k)^{2}e^{ika})^{-1}$, where $k=\sqrt{\omega^{2}-V_{0}}$.
-In order to obtain the quasi-normal spectra, we need purely outgoing waves, i.e.
-$a_{t}\rightarrow\infty$, Boonserm and Visser (2011) which implies $(\omega+k)^{2}e^{-ika}-(\omega-k)^{2}e^{ika}=0$, and we can solve this equation numerically to get the QNMs of the rectangular barrier.
-Here, we are interested in how the QNM spectra change if we change the width (a) of the potential while keeping the height of the barrier fixed (say $V_{0}=1$).
-Figure 30: Rectangular Barriers with the same height ($V_{0}=1$) and slightly different widths.
-Figure 31: QNMs for Rect. Barrier of different widths. Results of the semi-analytical method and Prony fit are represented by a circle and a plus, respectively.
-We have computed the QNM spectra (Fig.
-31) for a rectangular barrier with three different barrier widths (Fig.
-31).
-We observe that, as we increase the width of the barrier, the $Re(\omega)$ decreases.
-Thus, we encounter a redshift for wider barriers.
-The QNM separation is more pronounced for higher overtones (see Figure 31, rightwards).
-VI.2 Poschl-Teller Potential:
-The potential is defined as: $V=V_{0}sech^{2}(\alpha x)$ and the corresponding master equation is:
-$\frac{d^{2}\psi}{dx^{2}}+\left(\omega^{2}-\frac{V_{0}}{\cosh^{2}(\alpha x)}\right)\psi=0.$
-Using $\xi=\tanh(\alpha x)$ and rewriting the equation, we get
-$\frac{d}{d\xi}\left((1-\xi^{2})\frac{d\psi}{d\xi}\right)+\left[s(s+1)+\frac{\omega^{2}}{\alpha^{2}(1-\xi^{2})}\right]\psi=0\hskip 28.45274ptwhere\hskip 8.5359pts(s+1)=-\frac{V_{0}}{\alpha^{2}}$ (31)
-The asymptotic solution at $x\to-\infty$ is given by Landau and Lifshitz (2013):
-$\psi\sim e^{-i\omega x}\frac{\Gamma(\frac{i\omega}{\alpha})\Gamma(1-\frac{i\omega}{\alpha})}{\Gamma(-s)\Gamma(1+s)}+e^{i\omega x}\frac{\Gamma(-\frac{i\omega}{\alpha})\Gamma(1-\frac{i\omega}{\alpha})}{\Gamma(-\frac{i\omega}{\alpha}-s)\Gamma(-\frac{i\omega}{\alpha}+s+1)}$
-At the asymptotic limits, the solution must be purely outgoing.
-Thus, we have to find the values of $\omega$ such that the coefficient of $e^{i\omega x}$ becomes zero.
-We can solve this analytically, and the QNMs are given by Ferrari and Mashhoon (1984); Beyer (1999); Cardona and Molina (2017)
-$\omega=\alpha\left[-i\left(n+\frac{1}{2}\right)\pm\sqrt{\frac{V_{0}}{\alpha^{2}}-\frac{1}{4}}\right]$ (32)
-where, $n=0$ corresponds to the fundamental mode and $n=1,2,3,...$ represent overtones.
-Figure 32: Poschl-Teller potential of different values of $\alpha$.
-Figure 33: QNMs for Poschl-Teller of different values of $\alpha$.
-From the plots above, one can note that contrary to the rectangular barrier case, for the Poschl-Teller potential, as we increase the value of $\alpha$ (decrease width), keeping the height fixed (Fig.
-33), the real part of the $\omega_{QNM}$ decreases (Fig.
-33).
-So, increasing the value of $\alpha$ in the Poschl-Teller potential causes a redshift.
-In both the rectangular barrier and the Poschl-Teller potential case, reducing the height of the potential barrier (keeping the width fixed) causes a reduction in the $Re(\omega)$, thus causing a redshift.
-VI.3 Bronnikov–Ellis geometry with asymptotic mass:
-Let us now see if a similar redshift effect arises for the QNMs of perturbations of Bronnikov–Ellis geometries.
-The ultrastatic Bronnikov–Ellis geometry is not of use since it has only one parameter $b_{0}$ (throat radius) which essentially appears as an overall scale factor in the QNMs.
-Thus, we either modify the shape function with extra parameters, in which case it is no longer Bronnikov–Ellis or, alternatively, we may bring in the redshift function which will have an extra parameter.
-We choose to work with this latter modification.
-The well-known, original line element for a Bronnikov–Ellis geometry with a redshift function is given as,
-$ds^{2}=-e^{f(l)}dt^{2}+e^{-f(l)}[dl^{2}+(l^{2}+b_{0}^{2})d\Omega_{2}^{2}]$ (33)
-where, $f(l)=\frac{C}{b_{0}}\left(arctan\left(\frac{l}{b_{0}}\right)-\frac{\pi}{2}\right)$.
-Note that due to the presence of $f(l)$, the energy density as well as the pressures change, as discussed in Section II.2.
-We will work with this line element in order to demonstrate the redshift effect in the QNMs.
-The ADM mass ($M$) of this spacetime is related to the metric parameter $C$ as $C=2M$.
-One can plot the effective potentials for both scalar and axial perturbations numerically and observe that the symmetry in the potential barriers is lost with the introduction of non-zero asymptotic mass ($M$).
-The QNM spectra computed using the Prony fit (from $t=0$ to $t=40$ with $h=1/40$ using Mathematica 14.2 for $M\neq 0$, $b_{0}=1$), for this geometry, are presented in the tables 4,5 .
-However, in the figures we have presented some results from the spectral method.
-The effective potential (scalar perturbation) for this metric can be expressed as Eq.
-(17).
-Here, we have plotted the effective potentials for the $n=2$ case ($m=5$ in Fig.
-35 and $m=7$ in Fig.
-35) and obtained single-peaked potentials with the maxima of the potential shifting to the right as we increase the value of $C$ ($=2M$).
-The QNMs corresponding to $m=5$ and $m=7$ are plotted in Figs.
-37 and 37 respectively.
-Figure 34: Effective potentials (Scalar perturbation) for different values of M. (n = 2 and m = 5).
-Figure 35: Effective potentials (Scalar perturbation) for different values of M. (n = 2 and m = 7).
-Figure 36: QNMs (Scalar) for different values of M. (n = 2 and m = 5). The results of SM and Prony fit are represented by circles and plus signs, respectively.
-Figure 37: QNMs (Scalar) for different values of M. (n = 2 and m = 7). The results of SM and Prony fit are represented by circles and plus signs, respectively.
-$m$ Mode $M=0b_{0}$ $M=0.025b_{0}$ $M=0.05b_{0}$ $M=0.075b_{0}$ $M=0.1b_{0}$
-3 0 $3.53443-0.50686i$ $3.27166-0.46911i$ $3.03576-0.43526i$ $2.82383-0.40465i$ $2.63255-0.37783i$
-5 0 $5.52282-0.50282i$ $5.11243-0.46532i$ $4.74378-0.43185i$ $4.41259-0.40161i$ $4.11427-0.37699i$
-1 $5.41796-1.52258i$ $5.04381-1.38243i$ $4.65962-1.30864i$ $4.33811-1.21722i$ $4.08699-1.09940i$
-2 $5.25123-2.54327i$ $4.85855-2.36301i$ $4.44292-2.19880i$ $4.14809-2.04224i$-
-7 0 $7.51934-0.50106i$ $6.95985-0.46386i$ $6.45810-0.43037i$ $6.00711-0.40048i$ $5.60108-0.37349i$
-1 $7.45576-1.50817i$ $6.91398-1.38873i$ $6.40556-1.30364i$ $5.95396-1.20411i$ $5.55057-1.11919i$
-2 $7.31985-2.52412i$ $6.79114-2.33889i$ $6.28485-2.12915i$ $5.84509-2.02687i$ $5.46827-1.89265i$
-9 0 $9.51561-0.50062i$ $8.81008-0.46296i$ $8.17473-0.42973i$ $7.60378-0.39983i$ $7.09023-0.37301i$
-1 $9.43351-1.50092i$ $8.76270-1.41073i$ $8.13121-1.29187i$ $7.55941-1.20308i$ $7.04588-1.12225i$
-2 $9.35764-2.51354i$ $8.64660-2.31498i$ $8.02958-2.14977i$ $7.50736-2.00358i$ $6.98478-1.87115i$
-Table 4: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Scalar perturbation) with $M=0,0.025b_{0},0.05b_{0},0.075b_{0},0.1b_{0}$.
-(The results are obtained using Prony fit.) [The fundamental modes were computed in Blázquez-Salcedo et al.
-(2018b).
-We extended the analysis to include overtones.]
-The effective potential for axial perturbation has the form Eq.
-(19) and are plotted below.
-With increasing M ($=\frac{C}{2}$), the height of the barrier reduces and the potentials become more asymmetric (See Figs.
-39, 39).
-The figures 41 and 41 display the QNM spectra for axial perturbation in $m=5$ and $m=7$ cases.
-Figure 38: Effective potentials (Axial Perturbation) for different values of M. (n = 2 and m = 5).
-Figure 39: Effective potentials (Axial Perturbation) for different values of M. (n = 2 and m = 7).
-Figure 40: QNMs (Axial Perturbation) for different values of M. (n = 2 and m = 5). The results of SM and Prony fit are represented by circles and plus signs, respectively.
-Figure 41: QNMs (Axial Perturbation) for different values of M. (n = 2 and m = 7). The results of SM and Prony fit are represented by circles and plus signs, respectively.
-$m$ Mode $M=0b_{0}$ $M=0.025b_{0}$ $M=0.05b_{0}$ $M=0.075b_{0}$ $M=0.1b_{0}$
-5 0 $5.15176-0.44366i$ $4.77236-0.43237i$ $4.42882-0.40144i$ $4.12030-0.37367i$ $3.84389-0.34924i$
-1 $5.06686-1.39798i$ $4.73155-1.33422i$ $4.36034-1.21916i$ $4.06004-1.12686i$ $3.76455-0.99294i$
-2 $4.91958-2.39447i$ $4.55509-2.21569i$- -- --
-7 0 $7.25073-0.48257i$ $6.71107-0.44682i$ $6.22750-0.41477i$ $5.79355-0.38607i$ $5.40297-0.36046i$
-1 $7.19387-1.45787i$ $6.68408-1.34773i$ $6.17347-1.25075i$ $5.74208-1.16371i$ $5.35677-1.08268i$
-2 $7.05993-2.43004i$ $6.52963-2.26610i$ $6.04865-2.09636i$ $5.63629-1.94025i$ $5.24520-1.82708i$
-9 0 $9.30408-0.48927i$ $8.61413-0.45250i$ $7.99283-0.41976i$ $7.43555-0.39076i$ $6.97356-0.37212i$
-1 $9.25255-1.47571i$ $8.56364-1.36779i$ $7.88056-1.26327i$ $7.40840-1.17056i$ $6.88158-1.09045i$
-2 $9.15068-2.45648i$ $8.47519-2.27130i$ $7.84263-2.09637i$ $7.27180-1.96059i$ $6.88756-1.86712i$
-Table 5: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Axial perturbation) with $M=0,0.025b_{0},0.05b_{0},0.075b_{0},0.1b_{0}$.
-(The results are obtained using Prony fit.) [The fundamental modes were computed in Blázquez-Salcedo et al.
-(2018b).
-We extended the analysis to include overtones.]
-For this geometry, the increase of asymptotic mass (M) can be associated with redshift, i.e.
-for higher values of M, we have lower values of $Re(\omega)$.
-This phenomenon is similar to that observed by the authors of Pezzella et al.
-(2025).
-We have presented the quasi-normal modes for small values of M, as we could not get overtones for higher values of M.
-The fundamental modes are in agreement with the results of Blázquez-Salcedo et al.
-(2018b).
-It is interesting to note that both the real and imaginary parts of the $\omega_{QNM}$ decrease with increasing M, with the fundamental mode and the overtones following the same pattern.
-We can estimate the QNMs for a certain M value from the spectra of the original Bronnikov–Ellis wormhole ($M=0$ case) using the purely empirical formula
-$\omega_{M}=\omega_{0}exp\left(-\frac{\pi M}{1+0.8M}\right)$
-where, $\omega_{0}$ is quasi-normal mode corresponding to $M=0$ case and $\omega_{M}$ denotes quasi-normal modes corresponding to $M\neq 0$ cases.
-This relation is true for both scalar and axial perturbations of the $n=2$ geometry and holds for all the values of the overtone number and the angular momentum parameter (i.e.
-the ones we have obtained numerically).
-We are not aware of the origin of this empirical relation, though, for small values of the exponent, an approximate linear relation seems to work, as evident from the plots.
-In Figs. 43 (scalar) and 43 (axial), we have plotted the QNMs (obtained using the Prony Fit method) and the above empirical fitting function for the $n=2$ and $m=7$ case.
-Figure 42: Fit of the scalar $\omega_{QNM}$ ($m=7$) using the empirical fitting function.
-Figure 43: Fit of the axial $\omega_{QNM}$ ($m=7$) using the empirical fitting function.
-For the $n=2$ case, one can observe that the increasing ADM mass ($M$) has a very similar effect on fundamental mode and overtones.
-Due to this feature of the QNM spectra, the ringdown signal of a Bronnikov-Ellis wormhole with non-zero $M$ can be mistaken as that of the massless Bronnikov-Ellis wormhole with a larger throat radius ($b_{0}$).
-VI.4 Bronnikov–Ellis geometry with a chosen redshift function:
-Apart from the geometry we studied in the previous subsection, we may also consider the Bronnikov–Ellis wormhole (the $n=2$ case) with a different non-zero redshift function, introduced and discussed earlier in Section II.3.
-The line element is given as,
-$ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\frac{dr^{2}}{1-\frac{b_{0}^{2}}{r^{2}}}+r^{2}d\theta^{2}+r^{2}sin^{2}\theta d\varphi^{2}$
-Here, as stated before, the energy density does not change on varying $M$, but the pressures change. We will now see if a redshift effect of the QNMs is present when we tune the value of $M$.
-The Schrodinger-like equation for perturbations of this spacetime is
-$\frac{d^{2}\psi(r)}{dr_{*}^{2}}+(\omega^{2}-V_{eff}(r))\psi(r)=0$ (34)
-where the tortoise coordinate $r_{*}$ has the form
-$r_{*}=$ $\frac{1}{\sqrt{r\left(r^{2}-b_{0}^{2}\right)(r-2M)}(4M^{2}-b_{0}^{2})}\bigg{(}b_{0}\bigg{(}b_{0}^{2}k_{1}(r-2M)^{2}\left(E\left(k_{0}|a_{1}\right)-F\left(k_{0}|a_{1}\right)\right)+2b_{0}M\big{(}k_{1}(r-2M)^{2}\big{(}E\left(k_{0}|a_{1}\right)$
-$-\Pi\left(a_{0};k_{0}|a_{1}\right)\big{)}-2Mr\big{)}-4k_{1}M^{2}(r-2M)^{2}\left(F\left(k_{0}|a_{1}\right)-\Pi\left(a_{0};k_{0}|a_{1}\right)\right)+b_{0}^{3}r\bigg{)}+r^{3}(4M^{2}-b_{0}^{2})\bigg{)}$
-with $a_{0}=\frac{2b_{0}}{b_{0}+2M}$, $a_{1}=\frac{4M}{b_{0}+2M}$, $k_{0}=sin^{-1}\sqrt{\frac{\left(b_{0}+2M\right)\left(r-b_{0}\right)}{2b_{0}(r-2M)}}$, $k_{1}=\sqrt{\frac{r\left(b_{0}-2M\right)\left(b_{0}^{2}-4M^{2}\right)\left(r^{2}-b_{0}^{2}\right)}{b_{0}^{3}(r-2M)^{3}}}$, $E(...)=EllipticE$, $F(...)=EllipticF$ and $\Pi(...)=EllipticPi$ function.
-Now, to compute the QNMs with the spectral method, it is convenient to consider two separate ansatz Alfaro et al. (2024) (one for odd overtones and one for even).
-$\psi(r)_{even}=e^{i\omega r}r^{iM\omega}\phi(r)_{even}\hskip 56.9055pt\psi(r)_{odd}=(r-b_{0})^{\frac{1}{2}}e^{i\omega r}r^{iM\omega-\frac{1}{2}}\phi(r)_{odd}$
-Further, we have used the transformation $x=1-\frac{2b_{0}}{r}$ to map the radial coordinate $r\in(b_{0},\infty)$ to $x\in[-1,1]$ Batic and Dutykh (2024).
-This procedure is applicable for both scalar and axial perturbations.
-The tables below (6, 7) contain results obtained using the spectral method.
-However, we have also shown the results from the Prony fit method in the plots.
-For scalar perturbation the expression for $V_{eff}$ is
-$V_{sc}(r)=\frac{b_{0}^{2}(r-3M)+r^{2}\left(\left(-2m^{2}-2m+1\right)M+m(m+1)r\right)}{r^{5}}$
-As shown in the figures below (Figs. 45, 45), the effective potential changes from being a single barrier to a double barrier for higher $M$. The potentials have been plotted numerically.
-Figure 44: Effective potentials (Scalar Perturbation) for different values of $M$. (n = 2 and m = 5).
-Figure 45: Effective potentials (Scalar Perturbation) for different values of $M$. (n = 2 and m = 7).
-For this geometry, as we increase the value of $M$, both the real and imaginary parts of the QNM spectra decrease (See Figs.
-47 ,47).
-Thus, an increase in $M$ causes a redshift.
-On careful observation, one notices that for lower values of $M$, the real parts of quasinormal modes decrease with the overtone number.
-However, this pattern reverses when we go to higher $M$ values.
-The height of the effective potential decreases as we keep on increasing $M$, and after a certain value, the effective potential becomes double-peaked.
-This may explain the above feature of the quasi-normal spectra.
-For a fixed overtone number, the $\omega_{QNM}$ of scalar perturbation corresponding to different $M$ values can be roughly estimated by linear fits.
-The fitting parameter will vary with the overtone number and angular momentum parameter (m).
-Figure 46: QNMs (Scalar) for different values of $M$. (n = 2 and m = 5). The results of SM and Prony fit are represented by circles and plus signs, respectively.
-Figure 47: QNMs (Scalar) for different values of $M$. (n = 2 and m = 7). The results of SM and Prony fit are represented by circles and plus signs, respectively.
-$m$ Mode $M=0b_{0}$ $M=0.1b_{0}$ $M=0.2b_{0}$ $M=0.3b_{0}$ $M=0.4b_{0}$
-1 0 $1.57271-0.52970i$ $1.42443-0.44987i$ $1.26184-0.35838i$ $1.07980-0.25189i$ $0.86152-0.12862i$
-1 $1.25957-1.70681i$ $1.18531-1.43536i$ $1.12961-1.14815i$ $1.08147-0.84604i$ $1.01104-0.51803i$
-2 -- -- $0.97385-2.14007i$ $1.09839-1.63586i$ $1.18106-1.10171i$
-3 0 $3.53429-0.50692i$ $3.17040-0.42628i$ $2.76272-0.32941i$ $2.29567-0.20688i$ $1.73832-0.06152i$
-1 $3.39029-1.51877i$ $3.06406-1.29199i$ $2.71576-1.00573i$ $2.34754-0.67242i$ $1.93405-0.30467i$
-2 $3.10063-2.62534i$ $2.85146-2.20091i$ $2.61615-1.73263i$ $2.38730-1.23570i$ $2.08989-0.70557i$
-3 $2.71863-3.55029i$- $2.46305-2.54081i$ $2.39244-1.88891i$ $2.23465-1.18172i$
-5 0 $5.52231-0.50296i$ $4.94551-0.42182i$ $4.29517-0.32260i$ $3.54142-0.19058i$ $2.63686-0.03204i$
-1 $5.41372-1.50367i$ $4.87885-1.27072i$ $4.27080-0.97686i$ $3.60420-0.61373i$ $2.85899-0.20555i$
-2 $5.25011-2.54570i$ $4.74524-2.13600i$ $4.21729-1.65629i$ $3.67073-1.11220i$ $3.02866-0.52636i$
-3 $4.97014-3.61705i$ $4.54492-3.03031i$ $4.12908-2.37426i$ $3.71354-1.67104i$ $3.19144-0.91448i$
-7 0 $7.51649-0.50162i$ $3.19144-0.91448i$ $5.83575-0.31996i$ $4.79524-0.18192i$ $3.53888-0.01635i$
-1 $7.45225-1.50671i$ $6.67910-1.26365i$ $5.82019-0.96550i$ $4.85928-0.58145i$ $3.78057-0.14253i$
-2 $7.31542-2.52531i$ $6.58191-2.11572i$ $5.78590-1.62677i$ $4.93587-1.04502i$ $3.95845-0.40847i$
-3 $7.11330-3.56071i$ $6.43513-2.98262i$ $5.72857-2.31139i$ $4.99835-1.55865i$ $4.13449-0.74607i$
-Table 6: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Scalar perturbation) with $M=0,0.1b_{0}$, $0.2b_{0}$, $0.3b_{0}$ and $0.4b_{0}$.
-[The QNM spectra for this geometry (single-peaked scalar potentials) were published in Alfaro et al.
-(2024).
-We extended this to double-peaked potentials (Higher $M$)].
-[Results obtained using the spectral method ($b_{0}=1$).]
-In the axial perturbation case, the effective potential in Eq. (34) has the form
-$V_{ax}(r)=\frac{b_{0}^{2}(7M-3r)+r^{2}(m(m+1)r-(2m(m+1)+1)M)}{r^{5}}$
-Similar to the scalar perturbation, the effective potential due to axial perturbation changes from being a single barrier to a double barrier for higher $M$ (See Figs. 49 ,49).
-Figure 48: Effective potentials (Axial Perturbation) for different values of $M$. (n = 2 and m = 5).
-Figure 49: Effective potentials (Axial Perturbation) for different values of $M$. (n = 2 and m = 7).
-The quasi-normal spectra for axial perturbation have a similar redshift effect as the QNMs for scalar perturbation (Figs.
-51 ,51).
-However, unlike the scalar perturbation case, here, a linear fit does not estimate the $\omega_{QNM}$ of the same overtone order.
-For this modified Bronnikov-Ellis wormhole, the spectra change with increasing $M$.
-For the $M=0$ case, the real part decreases with increasing overtone number, but this trend gradually reverses with increasing $M$.
-Though both $Re(\omega)$ and $|Im(\omega)|$ decrease with $M$ (similar effect as increasing $b_{0}$ of the original Bronnikov-Ellis spacetime), the detection of the first few overtones can distinguish this geometry (with non-zero redshift function) from the original Bronnikov-Ellis wormhole.
-Figure 50: QNMs (Axial) for different values of $M$. (n = 2 and m = 5). The results of SM and Prony fit are represented by circles and plus signs, respectively.
-Figure 51: QNMs (Axial) for different values of $M$. (n = 2 and m = 7). The results of SM and Prony fit are represented by circles and plus signs, respectively.
-$m$ Mode $M=0b_{0}$ $M=0.1b_{0}$ $M=0.2b_{0}$ $M=0.3b_{0}$ $M=0.4b_{0}$
-3 0 $2.95241-0.40998i$ $2.65202-0.34116i$ $2.31704-0.25890i$ $1.93583-0.15765i$ $1.47895-0.04526i$
-1 $2.86643-1.25767i$ $2.59943-1.05329i$ $2.32019-0.81701i$ $2.02320-0.54789i$ $1.68143-0.25349i$
-2 $2.66056-2.19085i$ $2.47232-1.84041i$ $2.28366-1.46458i$ $2.09262-1.06263i$ $1.83772-0.62125i$
-5 0 $5.15482-0.46725i$ $4.61743-0.39009i$ $4.01226-0.29556i$ $3.31317-0.17059i$ $2.47482-0.02713i$
-1 $5.07286-1.40857i$ $4.56144-1.17693i$ $4.00038-0.89951i$ $3.38786-0.56119i$ $2.69934-0.18747i$
-2 $4.90746-2.37118i$ $4.44773-1.98388i$ $3.96646-1.53674i$ $3.46529-1.03582i$ $2.86794-0.49395i$
-3 $4.65661-3.37224i$ $4.27300-2.82560i$ $3.89988-2.21918i$ $3.51728-1.57299i$ $3.03059-0.87000i$
-7 0 $7.24818-0.48308i$ $6.48787-0.40372i$ $5.62868-0.30567i$ $4.62788-0.17089i$ $3.42055-0.01444i$
-1 $7.18438-1.45261i$ $6.44308-1.21430i$ $5.61809-0.92386i$ $4.69804-0.55209i$ $3.66405-0.13378i$
-2 $7.05634-2.43256i$ $6.35314-2.03436i$ $5.59222-1.56056i$ $4.78041-1.00184i$ $3.84139-0.39161i$
-3 $6.86588-3.43084i$ $6.21733-2.87112i$ $5.54494-2.22360i$ $4.84772-1.50282i$ $4.01693-0.72270i$
-9 0 $9.30171-0.48964i$ $8.32354-0.40939i$ $7.21638-0.30982i$ $5.92045-0.16947i$ $4.34739-0.00712i$
-1 $9.25035-1.47100i$ $8.28712-1.23004i$ $7.20768-0.93382i$ $5.98667-0.54158i$ $4.60839-0.09362i$
-2 $9.14741-2.45870i$ $8.21414-2.05635i$ $7.18761-1.56991i$ $6.07049-0.97279i$ $4.79271-0.31009i$
-3 $8.99239-3.45726i$ $8.10424-2.89222i$ $7.15227-2.22372i$ $6.14625-1.44932i$ $4.97551-0.60743i$
-Table 7: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Axial perturbation) with $M=0,0.1b_{0}$, $0.2b_{0}$, $0.3b_{0}$ and $0.4b_{0}$.
-[The QNM spectra for this geometry (single-peaked scalar potentials) were published in Alfaro et al.
-(2024).
-We extended this to aial perturbation and double-peaked potentials (Higher $M$)].
-[Results obtained using the spectral method ($b_{0}=1$).]
-VII Remarks and future directions
-In our work here, we have revisited the problem of computing the QNMs of the generalised Bronnikov-Ellis wormholes with emphasis on the overtones.
-Further, we have also studied the ‘redshift’ of QNMs induced by varying the metric parameters in analogy with the environment-induced proposed recently.
-Below, we briefly summarise our main results.
-$\bullet$ We extend the ultrastatic, generalised (two parameter–$b_{0}$, $n>2$) Bronnikov-Ellis wormhole formulated in Kar et al.
-(1995) by constructing a modified metric with a non-zero asymptotic mass $M$ (furthering the analysis in Blázquez-Salcedo et al.
-(2018b) for $n=2$).
-The geometry, energy-momentum tensor and the energy conditions for the new three-parameter ($b_{0}$, $n$, $M$) space-times are discussed briefly, followed by the computation of the corresponding effective potentials for scalar and axial gravitational perturbations.
-$\bullet$ We have used the spectral method (frequency domain) and the Prony Fit (time domain) to compute the QNM spectra (scalar and axial gravitational) for the ultrastatic case.
-Beyond the fundamental modes, we have computed 2-3 overtones (4 in some cases), which, as we show, may help in distinguishing different geometries (i.e.
-with different $n$).
-We have also pointed out key features of the QNM spectra and how the shape of the effective potentials may explain the trends in QNM behaviour.
-$\bullet$ The range of the throat radius ($b_{0}$) for detectable QNMs (in aLIGO) in the ultrastatic case is found.
-We also establish, for a few cases, using the time-domain waveform, the crucial role that overtones may play in fitting the signal at early times (when SNR is high).
-$\bullet$ Finally, we focused on the $n=2$ wormhole family with asymptotic mass $M$ and observed how the increase in $M$ causes a systematic redshift (lowering of values of $Re(\omega)$).
-We further observed this redshift effect (w.r.t.
-changing $M$) for the original Bronnikov–Ellis wormhole but with a chosen redshift function ($e^{2\Psi(r)}=1-\frac{2M}{r}$) Alfaro et al.
-(2024).
-The distinguishability of these geometries from their massless counterparts (with larger $b_{0}$) has also been mentioned.
-Various extensions of this work here are possible.
-We have not really succeeded in obtaining a large number of overtones which can give us a clearer picture of the full quasinormal mode spectrum.
-This aspect surely needs to be handled better.
-Further, though we have identified distinct features and have tried to demonstrate how overtones may affect a signal, a better and more complete understanding is desirable.
-In addition, learning more about polar perturbations and their overtones remains a difficult task which needs to be addressed.
-Finally, in the context of the redshift, though we see a correlation with the behaviour of the effective potentials, the results are case-specific and need to be understood in a more generic way.
-We hope to return to the abovementioned issues in future investigations.
-ACKNOWLEDGEMENTS
-ABM thanks the Indian Institute of Technology Kharagpur, India, for supporting him through a fellowship and allowing access to available computational facilities.
-Appendix A QNMs for ultrastatic, generalized Bronnikov–Ellis wormhole
+Appendix B QNMs for ultrastatic, generalized Bronnikov–Ellis wormhole
 We present here, in table form, the values of QNMs (accurate up to 3 decimal places) for the family of wormholes (with $b_{0}=1$).
@@
 In this ultra-static spacetime, if $(\omega_{R}-i\omega_{I})$ is a solution, then $(-\omega_{R}-i\omega_{I})$ is also a solution, both having the same amplitude.
-In the tables of the following sections, we show only the solution with a positive real part and double its amplitude (accounting for the amplitude of both solutions).
+In the tables below, we show only the solution with a positive real part and double its amplitude (accounting for the amplitude of both solutions).
 Scalar Perturbation Axial Perturbation
 $m$ Mode Spectral Method Prony Fit Amplitude Spectral Method Prony Fit Amplitude
-0 0 $0.68137-0.61775i$ $0.68402-0.61594i^{*}$- -- -- --
-1 $0.49683-2.17988i$- -- -- -- --
-1 0 $1.57271-0.52970i$ $1.57274-0.52969i^{*}$- -- -- --
-1 $1.25957-1.70681i$- -- -- -- --
-2 0 $2.54666-0.51267i$ $2.54925-0.51285i$ $0.35745915$ $1.73769-0.30514i$ $1.73771-0.30513i^{*}$ $0.2689666$
-1 $2.35300-1.57522i$- -- $1.71239-1.02010i$- --
-2 $1.94641-2.75472i$- -- $1.52495-2.03050i$- --
-3 0 $3.53429-0.50692i$ $3.53443-0.50686i$ $0.30122677$ $2.95241-0.40998i$ $2.95253-0.41006i$ $0.29811729$
-1 $3.39038-1.53568i$- -- $2.86643-1.25767i$- --
-2 $3.09988-2.62324i$ $3.09682-2.62481i$ $0.16837004$ $2.66056-2.19085i$ $2.67518-2.17517i$ $0.171151$
-4 0 $4.52705-0.50433i$ $4.52734-0.50424i$ $0.26517902$ $4.07625-0.44911i$ $4.07646-0.44904i$ $0.2640608$
-1 $4.41536-1.52340i$- -- $3.99432-1.36285i$- --
-2 $4.19023-2.57208i$ $4.19075-2.57242i$ $0.1397532$ $3.80100-2.30237i$ $3.79816-2.30549i$ $0.14061934$
-5 0 $5.52231-0.50296i$ $5.52282-0.50282i$ $0.23964086$ $5.15482-0.46725i$ $5.15524-0.46714i$ $0.239214$
-1 $5.42996-1.51789i$ $5.41796-1.52258i$ $2.009\times 10^{-5}$ $5.07306-1.40721i$- --
-2 $5.24705-2.54683i$ $5.25123-2.54327i$ $0.12190640$ $4.90770-2.37002i$ $4.90815-2.37157i$ $0.12362226$
-6 0 $6.51897-0.50214i$ $6.51980-0.50195i$ $0.22033499$ $6.20882-0.47711i$ $6.20954-0.47695i$ $0.2201443$
-1 $6.44246-1.50964i$ $6.44313-1.53221i$ $1.886\times 10^{-5}$ $6.13729-1.43537i$- --
-2 $6.28646-2.53346i$ $6.28819-2.53269i$ $0.11151401$ $5.99151-2.40927i$ $5.99267-2.40806i$ $0.1115693$
-7 0 $7.51649-0.50162i$ $7.51934-0.50106i$ $0.20454891$ $7.24818-0.48307i$ $7.24931-0.48285i$ $0.20497677$
-1 $7.44933-1.50895i$ $7.45576-1.50817i$ $2.309\times 10^{-5}$ $7.24818-0.48307i$- --
-2 $7.31541-2.52544i$ $7.31985-2.52412i$ $0.10156691$ $7.05635-2.43256i$ $7.05785-2.43166i$ $0.10294975$
-Table 8: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=2$) under scalar perturbations for $m=0$ $7$ and axial perturbations for $m=2$ $7$.
+0 0 0.681370.61775i0.68137-0.61775i $0.68402-0.61594i^{*}$ - - - -
+1 0.496832.17988i0.49683-2.17988i - - - - -
+1 0 1.572710.52970i1.57271-0.52970i $1.57274-0.52969i^{*}$ - - - -
+1 1.259571.70681i1.25957-1.70681i - - - - -
+2 0 2.546660.51267i2.54666-0.51267i 2.549250.51285i2.54925-0.51285i $0.35745915$ 1.737690.30514i1.73769-0.30514i $1.73771-0.30513i^{*}$ $0.2689666$
+1 2.353001.57522i2.35300-1.57522i - - 1.712391.02010i1.71239-1.02010i - -
+2 1.946412.75472i1.94641-2.75472i - - 1.524952.03050i1.52495-2.03050i - -
+3 0 3.534290.50692i3.53429-0.50692i 3.534430.50686i3.53443-0.50686i $0.30122677$ 2.952410.40998i2.95241-0.40998i 2.952530.41006i2.95253-0.41006i $0.29811729$
+1 3.390381.53568i3.39038-1.53568i - - 2.866431.25767i2.86643-1.25767i - -
+2 3.099882.62324i3.09988-2.62324i 3.096822.62481i3.09682-2.62481i $0.16837004$ 2.660562.19085i2.66056-2.19085i 2.675182.17517i2.67518-2.17517i $0.171151$
+4 0 4.527050.50433i4.52705-0.50433i 4.527340.50424i4.52734-0.50424i $0.26517902$ 4.076250.44911i4.07625-0.44911i 4.076460.44904i4.07646-0.44904i $0.2640608$
+1 4.415361.52340i4.41536-1.52340i - - 3.994321.36285i3.99432-1.36285i - -
+2 4.190232.57208i4.19023-2.57208i 4.190752.57242i4.19075-2.57242i $0.1397532$ 3.801002.30237i3.80100-2.30237i 3.798162.30549i3.79816-2.30549i $0.14061934$
+5 0 5.522310.50296i5.52231-0.50296i 5.522820.50282i5.52282-0.50282i $0.23964086$ 5.154820.46725i5.15482-0.46725i 5.155240.46714i5.15524-0.46714i $0.239214$
+1 5.429961.51789i5.42996-1.51789i 5.417961.52258i5.41796-1.52258i $2.009\times 10^{-5}$ 5.073061.40721i5.07306-1.40721i - -
+2 5.247052.54683i5.24705-2.54683i 5.251232.54327i5.25123-2.54327i $0.12190640$ 4.907702.37002i4.90770-2.37002i 4.908152.37157i4.90815-2.37157i $0.12362226$
+6 0 6.518970.50214i6.51897-0.50214i 6.519800.50195i6.51980-0.50195i $0.22033499$ 6.208820.47711i6.20882-0.47711i 6.209540.47695i6.20954-0.47695i $0.2201443$
+1 6.442461.50964i6.44246-1.50964i 6.443131.53221i6.44313-1.53221i $1.886\times 10^{-5}$ 6.137291.43537i6.13729-1.43537i - -
+2 6.286462.53346i6.28646-2.53346i 6.288192.53269i6.28819-2.53269i $0.11151401$ 5.991512.40927i5.99151-2.40927i 5.992672.40806i5.99267-2.40806i $0.1115693$
+7 0 7.516490.50162i7.51649-0.50162i 7.519340.50106i7.51934-0.50106i $0.20454891$ 7.248180.48307i7.24818-0.48307i 7.249310.48285i7.24931-0.48285i $0.20497677$
+1 7.449331.50895i7.44933-1.50895i 7.455761.50817i7.45576-1.50817i $2.309\times 10^{-5}$ 7.184381.45261i7.18438-1.45261i - -
+2 7.315412.52544i7.31541-2.52544i 7.319852.52412i7.31985-2.52412i $0.10156691$ 7.056352.43256i7.05635-2.43256i 7.057852.43166i7.05785-2.43166i $0.10294975$
+Table 8: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=2$) under scalar perturbations for $m=0$ - 77 and axial perturbations for $m=2$ - 77.
 [The QNM spectra for the $n=2$ case were published in Kim (2008); Batic and Dutykh (2024); Blázquez-Salcedo et al.
-(2018b); Alfaro et al.
+(2018a); Alfaro et al.
 (2024)]
@@
 $m$ Mode Spectral Method Prony Fit Amplitude Spectral Method Prony Fit Amplitude
-0 0 $0.83639-0.39420i$ $0.83639-0.39418i^{*}$- -- -- --
-1 $1.78264-1.42784i$- -- -- -- --
-1 0 $1.68973-0.28405i$ $1.68974-0.28404i^{*}$- -- -- --
-1 $2.28101-1.18068i$- -- -- -- --
-2 0 $2.64867-0.24369i$ $2.64856-0.24379i$ $0.30535863$ $1.86212-0.27064i$ $1.86214-0.27064i^{*}$ $0.24995736$
-1 $3.09550-0.97920i$- -- $1.97161-0.71811i$- --
-2 $3.73267-2.07179i$ $3.72699-2.07264i$ $0.22283351$ $2.55089-1.37091i$- --
-3 0 $3.60155-3.36888i$ $3.62865-0.22143i$ $0.24259909$ $3.14398-0.31728i$ $3.14405-0.31724i$ $0.24589997$
-1 $4.01130-0.86237i$- -- $3.18751-0.89811i$- --
-2 $4.52123-1.81580i$ $4.52041-1.81625i$ $0.18309362$ $3.48149-1.49460i$ $3.48082-1.49494i$ $0.1574885$
-4 0 $4.6164-0.20649i$ $4.61663-0.20641i$ $0.20468468$ $4.28394-0.28792i$ $4.28414-0.28787i$ $0.21497073$
-1 $4.96092-0.78763i$ $4.95844-0.78954i$ $0.00019476$ $4.40055-0.92088i$- --
-2 $5.40832-1.63672i$ $5.41382-1.63803i$ $0.15290777$ $4.59293-1.60512i$ $4.59135-1.60323i$ $0.1200131$
-3 $5.94108-2.65940i$- -- $4.98385-2.34634i$- --
-4 -- -- -- $5.60020-3.24404i$ $5.57299-3.21739i$ $0.12091509$
-5 0 $5.60815-0.19523i$ $5.60859-0.19513i$ $0.1788866$ $5.35662-0.25860i$ $5.35702-0.25851i$ $0.18764768$
-1 $5.92623-0.73455i$ $5.92481-0.73471i$ $0.00011546$ $5.53163-0.86711i$ $5.54394-0.87820i$ $1.795\times 10^{-5}$
-2 $6.33667-1.50966i$ $6.33719-1.51452i$ $0.13574164$ $5.74869-1.60115i$ $5.74805-1.60159i$ $0.11547397$
-3 $6.81719-2.45107i$- -- $5.91553-2.26906i$- --
-4 -- -- -- $6.46920-3.21880i$ $6.41666-3.21904i$ $0.0932583$
-6 0 $6.60210-0.18623i$ $6.60282-0.18610i$ $0.15997014$ $6.39998-0.23613i$ $6.40069-0.23600i$ $0.1668965$
-1 $6.90042-0.69409i$ $6.90291-0.69404i$ $1.259\times 10^{-5}$ $6.60250-0.80713i$ $6.60367-0.81068i$ $2.414\times 10^{-5}$
-2 $7.28536-1.41469i$ $7.28540-1.41346i$ $0.11879929$ $6.85764-1.53123i$ $6.85660-1.52978i$ $0.10977835$
-3 $7.72908-2.28317i$- -- $7.21059-2.33977i$- --
-4 $8.24197-3.28172i$ $8.25637-3.28542i$ $0.09481262$ $7.48045-3.19207i$ $7.50483-3.1369i$ $0.07558219$
-7 0 $7.59739-0.17875i$ $7.59853-0.17859i$ $0.1454134$ $7.42853-0.21893i$ $7.42962-0.21877i$ $0.15090914$
-1 $7.88016-0.66170i$ $7.88124-0.66145i$ $4.624\times 10^{-5}$ $7.64284-0.75603i$ $7.64344-0.75595i$ $6.13\times 10^{-6}$
-2 $8.24579-1.34029i$ $8.24539-1.33931i$ $0.10773933$ $7.91941-1.45107i$ $7.91918-1.45030i$ $0.10293865$
-3 $8.65746-2.16049i$- -- $8.22111-2.20332i$- --
-4 $9.13470-3.08560i$ $9.10375-3.06571i$ $0.0848906$ $8.54449-3.11196i$ $8.54457-3.06192i$ $0.06786849$
-Table 9: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=4$) under scalar perturbations for $m=0$ $7$ and axial perturbations for $m=2$ $7$.
+0 0 0.836390.39420i0.83639-0.39420i $0.83639-0.39418i^{*}$ - - - -
+1 1.782641.42784i1.78264-1.42784i - - - - -
+1 0 1.689730.28405i1.68973-0.28405i $1.68974-0.28404i^{*}$ - - - -
+1 2.281011.18068i2.28101-1.18068i - - - - -
+2 0 2.648670.24369i2.64867-0.24369i 2.648560.24379i2.64856-0.24379i $0.30535863$ 1.862120.27064i1.86212-0.27064i $1.86214-0.27064i^{*}$ $0.24995736$
+1 3.095500.97920i3.09550-0.97920i - - 1.971610.71811i1.97161-0.71811i - -
+2 3.732672.07179i3.73267-2.07179i 3.726992.07264i3.72699-2.07264i $0.22283351$ 2.550891.37091i2.55089-1.37091i - -
+3 0 3.601553.36888i3.60155-3.36888i 3.628650.22143i3.62865-0.22143i $0.24259909$ 3.143980.31728i3.14398-0.31728i 3.144050.31724i3.14405-0.31724i $0.24589997$
+1 4.011300.86237i4.01130-0.86237i - - 3.187510.89811i3.18751-0.89811i - -
+2 4.521231.81580i4.52123-1.81580i 4.520411.81625i4.52041-1.81625i $0.18309362$ 3.481491.49460i3.48149-1.49460i 3.480821.49494i3.48082-1.49494i $0.1574885$
+4 0 4.61640.20649i4.6164-0.20649i 4.616630.20641i4.61663-0.20641i $0.20468468$ 4.283940.28792i4.28394-0.28792i 4.284140.28787i4.28414-0.28787i $0.21497073$
+1 4.960920.78763i4.96092-0.78763i 4.958440.78954i4.95844-0.78954i $0.00019476$ 4.400550.92088i4.40055-0.92088i - -
+2 5.408321.63672i5.40832-1.63672i 5.413821.63803i5.41382-1.63803i $0.15290777$ 4.592931.60512i4.59293-1.60512i 4.591351.60323i4.59135-1.60323i $0.1200131$
+3 5.941082.65940i5.94108-2.65940i - - 4.983852.34634i4.98385-2.34634i - -
+4 - - - 5.600203.24404i5.60020-3.24404i 5.572993.21739i5.57299-3.21739i $0.12091509$
+5 0 5.608150.19523i5.60815-0.19523i 5.608590.19513i5.60859-0.19513i $0.1788866$ 5.356620.25860i5.35662-0.25860i 5.357020.25851i5.35702-0.25851i $0.18764768$
+1 5.926230.73455i5.92623-0.73455i 5.924810.73471i5.92481-0.73471i $0.00011546$ 5.531630.86711i5.53163-0.86711i 5.543940.87820i5.54394-0.87820i $1.795\times 10^{-5}$
+2 6.336671.50966i6.33667-1.50966i 6.337191.51452i6.33719-1.51452i $0.13574164$ 5.748691.60115i5.74869-1.60115i 5.748051.60159i5.74805-1.60159i $0.11547397$
+3 6.817192.45107i6.81719-2.45107i - - 5.915532.26906i5.91553-2.26906i - -
+4 - - - 6.469203.21880i6.46920-3.21880i 6.416663.21904i6.41666-3.21904i $0.0932583$
+6 0 6.602100.18623i6.60210-0.18623i 6.602820.18610i6.60282-0.18610i $0.15997014$ 6.399980.23613i6.39998-0.23613i 6.400690.23600i6.40069-0.23600i $0.1668965$
+1 6.900420.69409i6.90042-0.69409i 6.902910.69404i6.90291-0.69404i $1.259\times 10^{-5}$ 6.602500.80713i6.60250-0.80713i 6.603670.81068i6.60367-0.81068i $2.414\times 10^{-5}$
+2 7.285361.41469i7.28536-1.41469i 7.285401.41346i7.28540-1.41346i $0.11879929$ 6.857641.53123i6.85764-1.53123i 6.856601.52978i6.85660-1.52978i $0.10977835$
+3 7.729082.28317i7.72908-2.28317i - - 7.210592.33977i7.21059-2.33977i - -
+4 8.241973.28172i8.24197-3.28172i 8.256373.28542i8.25637-3.28542i $0.09481262$ 7.480453.19207i7.48045-3.19207i 7.504833.1369i7.50483-3.1369i $0.07558219$
+7 0 7.597390.17875i7.59739-0.17875i 7.598530.17859i7.59853-0.17859i $0.1454134$ 7.428530.21893i7.42853-0.21893i 7.429620.21877i7.42962-0.21877i $0.15090914$
+1 7.880160.66170i7.88016-0.66170i 7.881240.66145i7.88124-0.66145i $4.624\times 10^{-5}$ 7.642840.75603i7.64284-0.75603i 7.643440.75595i7.64344-0.75595i $6.13\times 10^{-6}$
+2 8.245791.34029i8.24579-1.34029i 8.245391.33931i8.24539-1.33931i $0.10773933$ 7.919411.45107i7.91941-1.45107i 7.919181.45030i7.91918-1.45030i $0.10293865$
+3 8.657462.16049i8.65746-2.16049i - - 8.221112.20332i8.22111-2.20332i - -
+4 9.134703.08560i9.13470-3.08560i 9.103753.06571i9.10375-3.06571i $0.0848906$ 8.544493.11196i8.54449-3.11196i 8.544573.06192i8.54457-3.06192i $0.06786849$
+Table 9: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=4$) under scalar perturbations for $m=0$ - 77 and axial perturbations for $m=2$ - 77.
 [The fundamental modes for the $n>2$ cases were published in Dutta Roy et al.
@@
 $m$ Mode Spectral Method Prony Fit Amplitude Spectral Method Prony Fit Amplitude
-0 0 $0.86815-0.34320i$ $0.86815-0.34319i^{*}$- -- -- --
-1 $2.04851-1.11401i$- -- -- -- --
-2 $3.38964-1.91968i$ $3.39861-1.92833i^{*}$- -- -- --
-1 0 $1.71264-0.22384i$ $1.71265-0.22383i^{*}$- -- -- --
-1 $2.49623-0.91263i$- -- -- -- --
-2 $3.64769-1.77155i$ $3.64832-1.77127i^{*}$- -- -- --
-2 0 $2.66518-0.17757i$ $2.66523-0.17779i$ $0.29315965$ $1.89443-0.24761i$ $1.89445-0.24761i^{*}$ $0.2266797$
-1 $3.24903-0.73077i$- -- $2.04700-0.61385i$- --
-2 $4.15868-1.54933i$ $4.15833-1.54843i$ $0.20476605$ $2.78063-1.07127i$ $2.77464-1.07131i^{*}$ $0.30966731$
-3 $5.28869-2.41854i$- -- -- -- --
-4 $6.52829-3.26724i$ $6.53412-3.27928i$ $0.40308849$- -- --
-3 0 $3.64016-0.15237i$ $3.64024-0.15233i$ $0.22714669$ $3.18363-0.28391i$ $3.18372-0.28388i$ $0.23197406$
-1 $4.12259-0.61681i$ $4.14739-0.63000i$ $0.00015229$ $3.25390-0.76629i$- --
-2 $4.86332-1.34256i$ $4.86269-1.34213i$ $0.18258948$ $3.65647-1.17197i$ $3.65640-1.17160i$ $0.16536179$
-3 $5.81271-2.19639i$- -- $4.52375-1.72969i$- --
-4 $6.94662-3.06045i$ $6.95026-3.06342i$ $0.12904273$ $5.60501-2.47686i$ $5.60406-2.47709i$ $0.15103107$
-4 0 $4.62419-0.13598i$ $4.62440-0.13592i$ $0.18790256$ $4.32457-0.23666i$ $4.32477-0.23661i$ $0.20448213$
-1 $5.04395-0.54241i$ $5.04408-0.54139i$ $1.859\times 10^{-5}$ $4.47738-0.78946i$- --
-2 $5.67924-1.18324i$ $5.67895-1.18265i$ $0.15644501$ $4.71713-1.30051i$ $4.71834-1.29988i$ $0.10481977$
-3 $6.51135-1.97082i$- -- $5.30511-1.77828i$- --
-4 $7.49665-2.83332i$ $7.49472-2.83864i$ $0.11727697$ $6.23010-2.39260i$ $6.22995-2.38931i$ $0.13343803$
-5 0 $5.61290-0.12418i$ $5.61332-0.12410i$ $0.16150451$ $5.39035-0.19645i$ $5.39075-0.19637i$ $0.17505116$
-1 $5.98915-0.48994i$ $5.98930-0.48917i$ $7.8\times 10^{-6}$ $5.62591-0.70862i$ $5.62610-0.71015i$ $4.315\times 10^{-5}$
-2 $6.55411-1.06444i$ $6.55411-1.06349i$ $0.13585498$ $5.87683-1.32855i$ $5.87679-1.32565i$ $0.09890309$
-3 $7.28065-1.79073i$- -- $6.26018-1.85524i$- --
-4 $8.15775-2.61273i$ $8.15569-2.60335i$ $0.10450563$ $6.96511-2.38395i$ $6.9809-2.36109i$ $0.10370454$
-6 0 $6.60441-0.11513i$ $6.60513-0.11501i$ $0.14239417$ $6.42671-0.16857i$ $6.42738-0.16847i$ $0.15284811$
-1 $6.94825-0.45059i$ $6.94923-0.45040i$ $0.00018714$ $6.69168-0.62009i$ $6.69217-0.62012i$ $0.00011249$
-2 $7.46242-0.97399i$ $7.46683-0.96882i$ $0.1139611$ $7.01628-1.24539i$ $7.01625-1.24492i$ $0.10544664$
-3 $8.11603-1.64203i$- -- $7.35083-1.88535i$- --
-4 $8.90307-2.41503i$ $8.79939-2.40900i$ $0.06811125$ $7.84927-2.43946i$ $7.84363-2.43251i$ $0.08016603$
-7 0 $7.59771-0.10787i$ $7.59880-0.10775i$ $0.12789715$ $7.44938-0.14872i$ $7.45045-0.14859i$ $0.13601879$
-1 $7.91626-0.41968i$ $7.91704-0.41916i$ $8.235\times 10^{-5}$ $7.71959-0.55112i$ $7.72040-0.55064i$ $3.859\times 10^{-5}$
-2 $8.39152-0.90296i$ $8.39220-0.90173i$ $0.10758787$ $8.08261-1.12967i$ $8.08311-1.12881i$ $0.10269501$
-3 $8.99190-1.52108i$ $8.93263-1.51227i$ $0.00374277$ $8.46623-1.80756i$- --
-4 $9.70803-2.24495i$ $9.69532-2.23123i$ $0.09946345$ $8.87902-2.45825i$ $8.88319-2.49385i$ $0.06576324$
-Table 10: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=6$) under scalar perturbations for $m=0$ $7$ and axial perturbations for $m=2$ $7$.
+0 0 0.868150.34320i0.86815-0.34320i $0.86815-0.34319i^{*}$ - - - -
+1 2.048511.11401i2.04851-1.11401i - - - - -
+2 3.389641.91968i3.38964-1.91968i $3.39861-1.92833i^{*}$ - - - -
+1 0 1.712640.22384i1.71264-0.22384i $1.71265-0.22383i^{*}$ - - - -
+1 2.496230.91263i2.49623-0.91263i - - - - -
+2 3.647691.77155i3.64769-1.77155i $3.64832-1.77127i^{*}$ - - - -
+2 0 2.665180.17757i2.66518-0.17757i 2.665230.17779i2.66523-0.17779i $0.29315965$ 1.894430.24761i1.89443-0.24761i $1.89445-0.24761i^{*}$ $0.2266797$
+1 3.249030.73077i3.24903-0.73077i - - 2.047000.61385i2.04700-0.61385i - -
+2 4.158681.54933i4.15868-1.54933i 4.158331.54843i4.15833-1.54843i $0.20476605$ 2.780631.07127i2.78063-1.07127i $2.77464-1.07131i^{*}$ $0.30966731$
+3 5.288692.41854i5.28869-2.41854i - - - - -
+4 6.528293.26724i6.52829-3.26724i 6.534123.27928i6.53412-3.27928i $0.40308849$ - - -
+3 0 3.640160.15237i3.64016-0.15237i 3.640240.15233i3.64024-0.15233i $0.22714669$ 3.183630.28391i3.18363-0.28391i 3.183720.28388i3.18372-0.28388i $0.23197406$
+1 4.122590.61681i4.12259-0.61681i 4.147390.63000i4.14739-0.63000i $0.00015229$ 3.253900.76629i3.25390-0.76629i - -
+2 4.863321.34256i4.86332-1.34256i 4.862691.34213i4.86269-1.34213i $0.18258948$ 3.656471.17197i3.65647-1.17197i 3.656401.17160i3.65640-1.17160i $0.16536179$
+3 5.812712.19639i5.81271-2.19639i - - 4.523751.72969i4.52375-1.72969i - -
+4 6.946623.06045i6.94662-3.06045i 6.950263.06342i6.95026-3.06342i $0.12904273$ 5.605012.47686i5.60501-2.47686i 5.604062.47709i5.60406-2.47709i $0.15103107$
+4 0 4.624190.13598i4.62419-0.13598i 4.624400.13592i4.62440-0.13592i $0.18790256$ 4.324570.23666i4.32457-0.23666i 4.324770.23661i4.32477-0.23661i $0.20448213$
+1 5.043950.54241i5.04395-0.54241i 5.044080.54139i5.04408-0.54139i $1.859\times 10^{-5}$ 4.477380.78946i4.47738-0.78946i - -
+2 5.679241.18324i5.67924-1.18324i 5.678951.18265i5.67895-1.18265i $0.15644501$ 4.717131.30051i4.71713-1.30051i 4.718341.29988i4.71834-1.29988i $0.10481977$
+3 6.511351.97082i6.51135-1.97082i - - 5.305111.77828i5.30511-1.77828i - -
+4 7.496652.83332i7.49665-2.83332i 7.494722.83864i7.49472-2.83864i $0.11727697$ 6.230102.39260i6.23010-2.39260i 6.229952.38931i6.22995-2.38931i $0.13343803$
+5 0 5.612900.12418i5.61290-0.12418i 5.613320.12410i5.61332-0.12410i $0.16150451$ 5.390350.19645i5.39035-0.19645i 5.390750.19637i5.39075-0.19637i $0.17505116$
+1 5.989150.48994i5.98915-0.48994i 5.989300.48917i5.98930-0.48917i $7.8\times 10^{-6}$ 5.625910.70862i5.62591-0.70862i 5.626100.71015i5.62610-0.71015i $4.315\times 10^{-5}$
+2 6.554111.06444i6.55411-1.06444i 6.554111.06349i6.55411-1.06349i $0.13585498$ 5.876831.32855i5.87683-1.32855i 5.876791.32565i5.87679-1.32565i $0.09890309$
+3 7.280651.79073i7.28065-1.79073i - - 6.260181.85524i6.26018-1.85524i - -
+4 8.157752.61273i8.15775-2.61273i 8.155692.60335i8.15569-2.60335i $0.10450563$ 6.965112.38395i6.96511-2.38395i 6.98092.36109i6.9809-2.36109i $0.10370454$
+6 0 6.604410.11513i6.60441-0.11513i 6.605130.11501i6.60513-0.11501i $0.14239417$ 6.426710.16857i6.42671-0.16857i 6.427380.16847i6.42738-0.16847i $0.15284811$
+1 6.948250.45059i6.94825-0.45059i 6.949230.45040i6.94923-0.45040i $0.00018714$ 6.691680.62009i6.69168-0.62009i 6.692170.62012i6.69217-0.62012i $0.00011249$
+2 7.462420.97399i7.46242-0.97399i 7.466830.96882i7.46683-0.96882i $0.1139611$ 7.016281.24539i7.01628-1.24539i 7.016251.24492i7.01625-1.24492i $0.10544664$
+3 8.116031.64203i8.11603-1.64203i - - 7.350831.88535i7.35083-1.88535i - -
+4 8.903072.41503i8.90307-2.41503i 8.799392.40900i8.79939-2.40900i $0.06811125$ 7.849272.43946i7.84927-2.43946i 7.843632.43251i7.84363-2.43251i $0.08016603$
+7 0 7.597710.10787i7.59771-0.10787i 7.598800.10775i7.59880-0.10775i $0.12789715$ 7.449380.14872i7.44938-0.14872i 7.450450.14859i7.45045-0.14859i $0.13601879$
+1 7.916260.41968i7.91626-0.41968i 7.917040.41916i7.91704-0.41916i $8.235\times 10^{-5}$ 7.719590.55112i7.71959-0.55112i 7.720400.55064i7.72040-0.55064i $3.859\times 10^{-5}$
+2 8.391520.90296i8.39152-0.90296i 8.392200.90173i8.39220-0.90173i $0.10758787$ 8.082611.12967i8.08261-1.12967i 8.083111.12881i8.08311-1.12881i $0.10269501$
+3 8.991901.52108i8.99190-1.52108i 8.932631.51227i8.93263-1.51227i $0.00374277$ 8.466231.80756i8.46623-1.80756i - -
+4 9.708032.24495i9.70803-2.24495i 9.695322.23123i9.69532-2.23123i $0.09946345$ 8.879022.45825i8.87902-2.45825i 8.883192.49385i8.88319-2.49385i $0.06576324$
+Table 10: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=6$) under scalar perturbations for $m=0$ - 77 and axial perturbations for $m=2$ - 77.
 [The fundamental modes for the $n>2$ cases were published in Dutta Roy et al.
@@
 $m$ Mode Spectral Method Prony Fit Amplitude Spectral Method Prony Fit Amplitude
-0 0 $0.87997-0.32519i$ $0.87998-0.32518i^{*}$- -- -- --
-1 $2.14466-0.97491i$- -- -- -- --
-2 $3.57108-1.59477i$ $3.57054-1.59493i^{*}$- -- -- --
-1 0 $1.72193-0.20105i$ $1.72194-0.20104i^{*}$- -- -- --
-1 $2.57737-0.79070i$- -- -- -- --
-2 $3.82365-1.47070i$ $3.82587-1.46941i^{*}$- -- -- --
-2 0 $2.67187-0.15181i$ $2.67251-0.15238i$ $0.29083766$ $1.91003-0.23457i$ $1.91005-0.23456i^{*}$ $0.21776703$
-1 $3.30697-0.61908i$- -- $2.08571-0.56896i$- --
-2 $4.31495-1.28267i$ $4.31427-1.28190i$ $0.23210154$ $2.87446-0.93747i$ $2.87405-0.93741i^{*}$ $0.29184553$
-3 $5.55107-1.96109i$- -- $4.04950-1.50405i$- --
-4 $6.88381-2.61925i$ $6.88773-2.61918i$ $0.25385346$ $5.37019-2.11453i$ $5.36933-2.11477i^{*}$ $0.1670686$
-3 0 $3.64456-0.12512i$ $3.64471-0.12510i$ $0.22202515$ $3.19594-0.26645i$ $3.19603-0.26648i$ $0.22253281$
-1 $4.16276-0.50814i$ $4.17354-0.50844i$ $0.00037032$ $3.28608-0.69780i$- --
-2 $4.98966-1.10188i$ $4.99007-1.10179i$ $0.18332356$ $3.73814-1.01798i$ $3.73866-1.01791i$ $0.17541431$
-3 $6.06575-1.77885i$- -- $4.69073-1.44514i$- --
-4 $7.29017-2.45688i$ $7.29374-2.46700i$ $0.13255104$ $5.86562-2.02004i$ $5.8667-2.02183i$ $0.15024292$
-4 0 $4.62667-0.10807i$ $4.62688-0.10801i$ $0.18186996$ $4.33741-0.21472i$ $4.33761-0.21467i$ $0.19968836$
-1 $5.07164-0.43530i$ $5.07146-0.43473i$ $5.0481\times 10^{-6}$ $4.49808-0.72204i$- --
-2 $5.77715-0.95771i$ $5.77702-0.95720i$ $0.15714858$ $4.76792-1.13746i$ $4.76742-1.13632i$ $0.10174903$
-3 $6.71074-1.59843i$- -- $5.44241-1.46518i$- --
-4 $7.81676-2.27620i$ $7.81867-2.27681i$ $0.11845874$ $6.47562-1.94993i$ $6.47398-1.95101i$ $0.13691266$
-5 0 $5.61380-0.09608i$ $5.61421-0.09601i$ $0.15498730$ $5.40063-0.17075i$ $5.40103-0.17067i$ $0.16992681$
-1 $6.00767-0.38426i$ $6.00786-0.38394i$ $1.7893\times 10^{-5}$ $5.65481-0.64034i$- --
-2 $6.62899-0.84805i$ $6.62886-0.84759i$ $0.13681384$ $5.90429-1.18229i$ $5.90559-1.17820i$ $0.08731672$
-3 $7.45181-1.43958i$ $7.45249-1.39775i$ $0.00057971$ $6.34149-1.56146i$- --
-4 $8.44520-2.09680i$ $8.44251-2.10081i$ $0.11003456$ $7.17592-1.93339i$ $7.18692-1.93325i$ $0.11707173$
-6 0 $6.60399-0.08711i$ $6.60469-0.08702i$ $0.13560445$ $6.43410-0.14111i$ $6.43478-0.14102i$ $0.14713892$
-1 $6.95973-0.34643i$ $6.96018-0.34614i$ $1.3489\times 10^{-5}$ $6.72107-0.54020i$ $6.72167-0.53992i$ $8.009\times 10^{-5}$
-2 $7.51913-0.76397i$ $7.51916-0.76313i$ $0.12066340$ $7.05523-1.11725i$ $7.05465-1.11561i$ $0.09874688$
-3 $8.25731-1.30704i$ $8.23586-1.33897i$ $0.00457058$ $7.39302-1.62858i$- --
-4 $9.15431-1.93145i$ $9.15519-1.93202i$ $0.10141119$ $7.98533-1.99226i$ $7.98248-1.98655i$ $0.08462117$
-7 0 $7.59621-0.08009i$ $7.59732-0.07999i$ $0.12093846$ $7.45434-0.12059i$ $7.45540-0.12047i$ $0.12992236$
-1 $7.92220-0.31715i$ $7.92296-0.31677i$ $3.2848\times 10^{-5}$ $7.74342-0.46352i$ $7.74431-0.46321i$ $1.434\times 10^{-5}$
-2 $8.43378-0.69798i$ $8.43393-0.69726i$ $0.10853332$ $8.13502-0.98974i$ $8.13560-0.98861i$ $0.10233494$
-3 $9.10647-1.19796i$ $9.10056-1.19398i$ $0.00094986$ $8.51244-1.59506i$- --
-4 $9.92416-1.78541i$ $9.92248-1.77980i$ $0.09058615$ $8.94988-2.06737i$ $8.95234-2.06370i$ $0.06334415$
-Table 11: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=8$) under scalar perturbations for $m=0$ $7$ and axial perturbations for $m=2$ $7$.
+0 0 0.879970.32519i0.87997-0.32519i $0.87998-0.32518i^{*}$ - - - -
+1 2.144660.97491i2.14466-0.97491i - - - - -
+2 3.571081.59477i3.57108-1.59477i $3.57054-1.59493i^{*}$ - - - -
+1 0 1.721930.20105i1.72193-0.20105i $1.72194-0.20104i^{*}$ - - - -
+1 2.577370.79070i2.57737-0.79070i - - - - -
+2 3.823651.47070i3.82365-1.47070i $3.82587-1.46941i^{*}$ - - - -
+2 0 2.671870.15181i2.67187-0.15181i 2.672510.15238i2.67251-0.15238i $0.29083766$ 1.910030.23457i1.91003-0.23457i $1.91005-0.23456i^{*}$ $0.21776703$
+1 3.306970.61908i3.30697-0.61908i - - 2.085710.56896i2.08571-0.56896i - -
+2 4.314951.28267i4.31495-1.28267i 4.314271.28190i4.31427-1.28190i $0.23210154$ 2.874460.93747i2.87446-0.93747i $2.87405-0.93741i^{*}$ $0.29184553$
+3 5.551071.96109i5.55107-1.96109i - - 4.049501.50405i4.04950-1.50405i - -
+4 6.883812.61925i6.88381-2.61925i 6.887732.61918i6.88773-2.61918i $0.25385346$ 5.370192.11453i5.37019-2.11453i $5.36933-2.11477i^{*}$ $0.1670686$
+3 0 3.644560.12512i3.64456-0.12512i 3.644710.12510i3.64471-0.12510i $0.22202515$ 3.195940.26645i3.19594-0.26645i 3.196030.26648i3.19603-0.26648i $0.22253281$
+1 4.162760.50814i4.16276-0.50814i 4.173540.50844i4.17354-0.50844i $0.00037032$ 3.286080.69780i3.28608-0.69780i - -
+2 4.989661.10188i4.98966-1.10188i 4.990071.10179i4.99007-1.10179i $0.18332356$ 3.738141.01798i3.73814-1.01798i 3.738661.01791i3.73866-1.01791i $0.17541431$
+3 6.065751.77885i6.06575-1.77885i - - 4.690731.44514i4.69073-1.44514i - -
+4 7.290172.45688i7.29017-2.45688i 7.293742.46700i7.29374-2.46700i $0.13255104$ 5.865622.02004i5.86562-2.02004i 5.86672.02183i5.8667-2.02183i $0.15024292$
+4 0 4.626670.10807i4.62667-0.10807i 4.626880.10801i4.62688-0.10801i $0.18186996$ 4.337410.21472i4.33741-0.21472i 4.337610.21467i4.33761-0.21467i $0.19968836$
+1 5.071640.43530i5.07164-0.43530i 5.071460.43473i5.07146-0.43473i $5.0481\times 10^{-6}$ 4.498080.72204i4.49808-0.72204i - -
+2 5.777150.95771i5.77715-0.95771i 5.777020.95720i5.77702-0.95720i $0.15714858$ 4.767921.13746i4.76792-1.13746i 4.767421.13632i4.76742-1.13632i $0.10174903$
+3 6.710741.59843i6.71074-1.59843i - - 5.442411.46518i5.44241-1.46518i - -
+4 7.816762.27620i7.81676-2.27620i 7.818672.27681i7.81867-2.27681i $0.11845874$ 6.475621.94993i6.47562-1.94993i 6.473981.95101i6.47398-1.95101i $0.13691266$
+5 0 5.613800.09608i5.61380-0.09608i 5.614210.09601i5.61421-0.09601i $0.15498730$ 5.400630.17075i5.40063-0.17075i 5.401030.17067i5.40103-0.17067i $0.16992681$
+1 6.007670.38426i6.00767-0.38426i 6.007860.38394i6.00786-0.38394i $1.7893\times 10^{-5}$ 5.654810.64034i5.65481-0.64034i - -
+2 6.628990.84805i6.62899-0.84805i 6.628860.84759i6.62886-0.84759i $0.13681384$ 5.904291.18229i5.90429-1.18229i 5.905591.17820i5.90559-1.17820i $0.08731672$
+3 7.451811.43958i7.45181-1.43958i 7.452491.39775i7.45249-1.39775i $0.00057971$ 6.341491.56146i6.34149-1.56146i - -
+4 8.445202.09680i8.44520-2.09680i 8.442512.10081i8.44251-2.10081i $0.11003456$ 7.175921.93339i7.17592-1.93339i 7.186921.93325i7.18692-1.93325i $0.11707173$
+6 0 6.603990.08711i6.60399-0.08711i 6.604690.08702i6.60469-0.08702i $0.13560445$ 6.434100.14111i6.43410-0.14111i 6.434780.14102i6.43478-0.14102i $0.14713892$
+1 6.959730.34643i6.95973-0.34643i 6.960180.34614i6.96018-0.34614i $1.3489\times 10^{-5}$ 6.721070.54020i6.72107-0.54020i 6.721670.53992i6.72167-0.53992i $8.009\times 10^{-5}$
+2 7.519130.76397i7.51913-0.76397i 7.519160.76313i7.51916-0.76313i $0.12066340$ 7.055231.11725i7.05523-1.11725i 7.054651.11561i7.05465-1.11561i $0.09874688$
+3 8.257311.30704i8.25731-1.30704i 8.235861.33897i8.23586-1.33897i $0.00457058$ 7.393021.62858i7.39302-1.62858i - -
+4 9.154311.93145i9.15431-1.93145i 9.155191.93202i9.15519-1.93202i $0.10141119$ 7.985331.99226i7.98533-1.99226i 7.982481.98655i7.98248-1.98655i $0.08462117$
+7 0 7.596210.08009i7.59621-0.08009i 7.597320.07999i7.59732-0.07999i $0.12093846$ 7.454340.12059i7.45434-0.12059i 7.455400.12047i7.45540-0.12047i $0.12992236$
+1 7.922200.31715i7.92220-0.31715i 7.922960.31677i7.92296-0.31677i $3.2848\times 10^{-5}$ 7.743420.46352i7.74342-0.46352i 7.744310.46321i7.74431-0.46321i $1.434\times 10^{-5}$
+2 8.433780.69798i8.43378-0.69798i 8.433930.69726i8.43393-0.69726i $0.10853332$ 8.135020.98974i8.13502-0.98974i 8.135600.98861i8.13560-0.98861i $0.10233494$
+3 9.106471.19796i9.10647-1.19796i 9.100561.19398i9.10056-1.19398i $0.00094986$ 8.512441.59506i8.51244-1.59506i - -
+4 9.924161.78541i9.92416-1.78541i 9.922481.77980i9.92248-1.77980i $0.09058615$ 8.949882.06737i8.94988-2.06737i 8.952342.06370i8.95234-2.06370i $0.06334415$
+Table 11: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=8$) under scalar perturbations for $m=0$ - 77 and axial perturbations for $m=2$ - 77.
 [The fundamental modes for the $n>2$ cases were published in Dutta Roy et al.